### v1

# Yongyang Cai  Khyati Malik   Hyeseon Shin

Abstract
We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world. We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments outlined in the 2015 Paris Agreement. Our model shows that the emission permit price reaches \$749 per ton of carbon by 2050. We demonstrate that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
Keywords: Emission trading system, Paris Agreement, dynamic Nash equilibrium, integrated assessment model, carbon tax, social cost of carbon.
JEL Classification: C73, F18, Q54, Q58.
1 Introduction
Anthropogenic warming of our planet is projected to have disastrous consequences, from rising sea levels to increases in the frequency and intensity of extreme weather events (Arias et al., 2021). The first ever international effort to mitigate global emissions was undertaken in 1997 with the signing of the Kyoto protocol. More than two decades since the Kyoto protocol, global efforts to tackle climate change have failed, with the world experiencing accelerating rates of greenhouse gas emissions. The policy outlined in the 2015 Paris Agreement aimed to codify an objective of limiting climate change to a maximum of 2 degrees Celsius above pre-industrial levels. To accomplish this objective, the participating nations committed to Intended Nationally Determined Contributions (INDCs) specifying their emission mitigation goals. In November 2021, representatives of almost 200 countries met at Glasgow to renew their efforts in curbing climate change and approved the Glasgow Climate Pact with the overall aim of limiting the rise in global temperatures to 1.5 degrees Celsius through numerous actions.
Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETS). In a carbon tax regime, a tax is charged on firms for each unit of carbon emissions. The optimal global carbon tax rate is often determined at the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus, 2017). On the other hand, in an emission trading or a cap-and-trade system, the maximum amount of emissions is fixed in an economy and agents can sell and purchase emission permits at the market price. Under the assumptions of complete information and no frictions, properly designed carbon tax regimes and emission trading systems are argued to have equivalent outcomes (Nordhaus and Yang, 1996; Goulder and Schein, 2013). The ETS, however, is often considered a more efficient policy tool in practice, since it encourages agents that can reduce emissions at lower costs than others to undertake more emission reduction efforts (Nordhaus, 2017). In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, and as of the summer of 2023, the permit price stands at about 90 euros per ton of carbon. [^1] Source of EU permit price: https://tradingeconomics.com/commodity/carbon. [^2] In addition to the EU ETS, other national or sub-national ETS have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States. For a review of emission cap-and-trade systems in practice, see e.g., Schmalensee and Stavins (2017). Although a global-level ETS is not currently in place, it has emerged as a prominent potential mechanism to combat climate change, particularly since Article 17 of the Kyoto Protocol proposed emissions trading among countries.
In this study, we develop a dynamic multi-region model of climate and the economy under a hypothetical global ETS, and use it to project global prices of emission permits and the regional Social Cost of Carbon (SCC). In this integrated assessment framework, we formed 12 aggregated regions by grouping 198 countries around the world. Each region is allocated emission caps in accordance with the emission mitigation targets from the Paris Agreement and regional net zero commitments. Our main analysis involves solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions. Under the global ETS, regions with higher marginal abatement costs (MAC) are encouraged to purchase permits from the market to meet their emission cap constraints, whereas regions with lower MAC are incentivized to conduct more abatement and make profits by selling permits. The emission permit price is endogenously determined by annual supply and demand in the global permit market.
Our study makes four contributions. The first contribution of this study is a method to solve for the market equilibrium prices of emission permits under a global ETS. To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with endogenous prices in a dynamic regional integrated assessment model. While prior research on the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus, 2010) highlighted the importance of a multi-region ETS, it did not provide a solution for the market-clearing equilibrium permit price in a noncooperative setting. We present an algorithm for finding a Nash equilibrium solution in this context. Under the baseline emission cap scenario based on the Paris Agreement and the regional net zero commitments, our simulation results show that the oversupply of global permits in the initial decade results in zero permit prices, but by 2050, the emission permit price can reach up to \$749 per ton of carbon. Furthermore, we show that regions such as the United States, India and the EU become major emission permit buyers, and other regions such as Russia and China become permit suppliers.
The second contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS. While carbon taxes and ETS have often been considered as potentially interchangeable policies, we show that within the global ETS framework, the social planner of each region can introduce a carbon tax as a complementary policy tool to achieve its optimal level of net emissions. The necessity of a carbon tax arises from the discrepancy in optimal net emissions between the social planner’s choice and the market outcome under the ETS. When optimal net emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional MAC and the global permit price under the noncooperative Nash equilibrium with the ETS. We also numerically verify that the regional optimal carbon tax is equal to the regional SCC when net emissions are strictly positive.
The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Paris Agreement, enhancing the accuracy and relevance of our findings. We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz, 2017), also known as the “Middle-of-the-Road” scenario. Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al., 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on cumulative carbon emissions. We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al., 2011). Due to its simplicity and effectiveness, the TCRE scheme has found application in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, 2017; Dietz and Venmans, 2019; Mattauch et al., 2020; Barnett et al., 2020. Moreover, Dietz et al. (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus, 2017), compared to other more complicated climate systems.
Our final contribution is a comprehensive comparative analysis of the roles played by cooperation, emission caps, and the global ETS. Our simulations reveal that global cooperation leads to significantly lower emissions and lower temperature rises. We also find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the first decade. This implies that it is important to maintain a stringent emission cap at the global level to make the ETS work efficiently. Moreover, we show that a stricter emission cap leads to a higher permit price and smaller regional SCCs for every region. For example, if all the regions achieve net zero emissions in 2050, the emission permit price can reach up to \$1,538 per ton of carbon by 2049. However, even with this strict emission restriction, the global temperature is still projected to rise by 1.64 degrees Celsius by the end of this century. These findings suggest that achieving the global target of limiting the temperature rise to 1.5 degrees Celsius requires even stronger emission commitments than those outlined in the Paris Agreement. We also find that under cooperation between the regions, the regional SCC is higher for a developed region compared to a developing one. That is, there are higher carbon taxes in developed regions under cooperation compared to developing regions because the marginal utility of consumption is higher in developing regions. Global aggregate welfare will be higher if developed regions can contribute more to mitigating emissions, thus reducing climate damages in developing regions. In the noncooperative setting, the regional SCCs of the developed and the developing regions are comparable.
In the literature, the ETS has garnered attention as a promising mechanism for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze the ETS as a potential tool for global climate policy. One closely relevant study is Carbone et al. (2009), which builds a computable general equilibrium model of the world, incorporating countries’ endogenous participation in an ETS and allocation of emission permits, and solves for the equilibrium permit price. In contrast to our work, their study is based on a static setting, so it does not address the critical dynamic aspects of climate change, and lacks a connection with emission abatement decisions. Fischer and Springborn (2011) is another relevant work that constructs a dynamic stochastic general equilibrium model to compare the outcomes of the emission cap and tax policies, but their model does not consider emission trading between regions.
More broadly, a series of recent studies have examined ETS from various perspectives. Some studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS or China ETS (Fuss et al., 2018; Hitzemann and Uhrig-Homburg, 2018; Borenstein et al., 2019; Goulder et al., 2022; Perino et al., 2022). Others analyze the efficiency gains from integrating regional ETS (Habla and Winkler, 2018; Doda et al., 2019; Holtsmark and Weitzman, 2020; Holtsmark and Midttømme, 2021), or compare the emission permit allocation methods between auctioning and free allocation (Goulder et al., 2010). While a global ETS is not yet in place, it is expected to offer multiple benefits compared to a global carbon tax. In an international ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al., 2016). An international ETS requires little governmental interference, and will help regions to maintain their emission target commitments and thereby contribute towards controlling the global temperature rise. On the other hand, a uniform global carbon tax regime is practically challenging. This paper therefore contributes to the literature by proposing a global ETS with endogenous permit prices and associated regional carbon taxes under various emission caps.
Our model framework is closely related to the recent developments in integrated assessment models (IAMs), which aim to capture the interaction between economic growth and climate systems under cooperation and noncooperation. A seminal work is the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010), which extends the global DICE model (Nordhaus, 2014) by building a multi-region dynamic framework. Nordhaus (2010) computes carbon prices under a cooperative model where global welfare is maximized with weights applied on regional utilities. Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region context under cooperation and noncooperation. For instance, Hambel et al. (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions. van der Ploeg and de Zeeuw (2016) and Jaakkola and Van der Ploeg (2019) focus on stochastic components of the model, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation. Cai et al. (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation. Nonetheless, none of these works examine the dynamics of an international ETS with an endogenous market price of emission permits under a noncooperative setting. Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw (2016) and Cai et al. (2023) and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
2 A Dynamic Regional Model of Climate and the Economy with an ETS
This section introduces a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement and net zero emissions targets.
The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model. We follow the RICE model and assume that the world consists of 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs). [^3] See Appendix 1 A.2 for the full list of countries and regional aggregation. Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions. Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial time period is the year 2020. All regions are forward-looking with complete information.
2.1 Emissions
The total industrial carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
$E_{i,t}^{\mathrm{Total}}=\sigma_{i,t}Q_{i,t},$ (1)
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al. (2019). The amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. That is,
$E_{i,t}^{A}=\mu_{i,t}\sigma_{i,t}Q_{i,t}.$ (2)
Thus the net emissions, $E_{i,t}$, becomes
$E_{i,t}=E_{i,t}^{\mathrm{Total}}-E_{i,t}^{A}.$ (3)
We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions. Regions with net emissions exceeding the emission caps can purchase emission permits from other regions. Regions with emissions below their allowance can sell their emission permits to others. The emission cap constraint is represented by
$E_{i,t}-E_{i,t}^{P}\leq\overline{E}_{i,t},$ (4)
where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions traded with other regions. Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$. In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$. The Paris Agreement establishes upper limits for net emissions to ensure that the increase in global temperature remains well below 2 degrees Celsius above preindustrial levels by the end of this century.
2.2 Climate System
The global average temperature rises as carbon emissions accumulate in the atmosphere. We follow the TCRE climate system and assume that the global average temperature increase above the pre-industrial level is approximately proportional to cumulative global emissions $\mathcal{E}_{t}$, i.e.,
$T_{t}=\zeta\mathcal{E}_{t},$ (5)
where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2 degrees Celsius above the pre-industrial level. Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al., 2011). Figure A.1 in Appendix A.3 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways. Cumulative global emissions evolve according to
$\mathcal{E}_{t+1}=\mathcal{E}_{t}+\sum_{i\in\mathcal{I}}E_{i,t}.$ (6)
2.3 Economic System
A representative consumer in region $i$ has a power utility per period as follows:
$u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (7)
where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus (2017) and set $\gamma=1.45$.
The production technology follows a Cobb-Douglas function of capital and labor inputs. The gross output (in trillion USD), or pre-damage output, $Q_{i,t}$, is given by
$Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (8)
where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus (2017)) is the elasticity of gross output with respect to capital. The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the no-climate-impact GDP projections in Burke et al. (2018). Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz, 2017).
We denote $Y_{i,t}$ as the output net of damages induced by climate changes:
$Y_{i,t}=\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}Q_{i,t},$ (9)
where $T_{t}$ is the global average temperature increase in degrees Celsius above the preindustrial level. We calibrate the climate damage parameters based on the projections of GDP loss in Kahn et al. (2021). We follow DICE and assume emission abatement cost is
$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$
where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$. We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al. (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
The total consumption of each region under the ETS is obtained by
$c_{i,t}L_{i,t}=Y_{i,t}-I_{i,t}-\Phi_{i,t}-m_{t}E_{i,t}^{P},$ (10)
where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits. The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption. The capital stock evolves according to the standard neoclassical growth model:
$K_{i,t+1}=(1-\delta)K_{i,t}+I_{i,t},$ (11)
where $\delta=0.1$ is the rate of depreciation of capital stock. We assume that the emission trading market clears in each period. That is,
$\sum_{i\in\mathcal{I}}E_{i,t}^{P}=0,$ (12)
for each time period $t$.
3 Solving for the Equilibrium
Based on the model components introduced in Section 2, we now define equilibrium under cooperative and noncooperative models in the ETS. These two extreme cases place bounds on what we would observe in a real world scenario.
3.1 The Cooperative and the Noncooperative Model
In the cooperative model, all regions work together to maximize their combined welfare. More specifically, the objective function is to maximize the overall sum of discounted social utility across all regions as follows:
${\color[rgb]{0,0,0}\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\sum_{i\in\mathcal{I}}\beta^{t}u(c_{i,t})L_{i,t}\end{split},}$ (13)
where $\beta$ is the discount factor. We follow DICE-2016R (Nordhaus, 2017) and set $\beta=0.985$. Note that under cooperation there is no market equilibrium permit price, so we assume permit prices follow an exogenous path. Therefore, we define the cooperative model equilibrium given an exogenous path of permit prices.
DEFINITION:
Given the initial capital and cumulative global emissions, $\{K_{i,0},\mathcal{E}_{0}:\ i\in\mathcal{I}\}$, and the exogenous paths of emission caps and permit prices, $\{\overline{E}_{i,t},m_{t}:\ i\in\mathcal{I},t\geq 0\}$, the dynamic equilibrium for the cooperative model is a sequence of quantities $\{c_{i,t},E_{i,t}^{P},\mu_{i,t},K_{i,t},\mathcal{E}_{t},T_{t}:\ i\in\mathcal{I},t\geq 0\}$ that solve the maximization problem (13) subject to equations (7)-(12).
In the noncooperative model, individual regions maximize their own lifetime social welfare, which represents another extreme but relatively more realistic perspective. The maximization problem for each region $i$ is defined as
$\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t}\end{split}.$ (14)
DEFINITION:
Given the initial capital and cumulative global emissions, $\{K_{i,0},\mathcal{E}_{0}:\ i\in\mathcal{I}\}$, and the exogenous paths of emission caps $\{\overline{E}_{i,t}:\ i\in\mathcal{I},t\geq 0\}$, the dynamic Nash equilibrium for the noncooperative model is a sequence of quantities $\{c_{i,t},E_{i,t}^{P},\mu_{i,t},K_{i,t},\mathcal{E}_{t},T_{t}:\ i\in\mathcal{I},t\geq 0\}$ and prices $\{m_{t}:\ t\geq 0\}$ that simultaneously solve the maximization problem (14) for all regions subject to equations (7)-(12).
The optimal solution for this dynamic multi-region model involves three choice problems. First, as in the standard Ramsey-type growth model, each region faces an intertemporal choice problem in which there is a trade-off between current consumption and future consumption. Each region may sacrifice present consumption to make investments, which can contribute to higher consumption in the future. Second, the intertemporal choice problem is further compounded by climate damages. Current production increases the global temperature, which subsequently lowers future productivity. Since emissions abatement has positive externalities, a region’s returns from abatement efforts may not be large enough to offset the cost of abatement. Therefore, the optimal solution of each region is highly dependent on the choices made by other regions. Lastly, the ETS allows each region to choose between purchasing emission permits from the market and undertaking further abatement. The ETS promotes efficient abatement globally by encouraging regions with better abatement technology or capacity (thus, with lower abatement cost) to conduct more abatement, and regions with less efficient abatement technology to purchase permits from other regions.
3.2 The Algorithm for the Noncooperative Model
Obtaining the optimal solution of the dynamic model involving multiple regions under noncooperation is challenging. In particular, finding an equilibrium solution for the emission permit prices that satisfies the optimality conditions for each region as well as the market clearing condition poses significant computational challenges. While several studies have incorporated emission prices in their models, no study has successfully solved for the emission permit price in a noncooperative context.
Here we outline the algorithm we develop to obtain the optimal solution for our model. Since $\gamma=1.45$ implies utilities of per capita consumption have negative values and that long-run per capita consumption will be large, the long-run per-capita utility has a small magnitude. Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years. Thus, we can transform the infinite horizon models to finite horizon models, where region $i$’s social welfare is rewritten as
$\sum_{t=0}^{299}\beta^{t}u(c_{i,t})L_{i,t}+\beta^{300}u\left(\frac{0.75Y_{i,300}}{L_{i,300}}\right)\frac{L_{i,300}}{1-\beta}.$
Here the second term is an approximation of the sum of discounted utility from time $t=300$ to infinity, assuming that consumption at any $t\geq 300$ is identical to 75 percent of the output at $t=300$ and that the exogenous population after 300 years stays at its value at $t=300$. The algorithm to solve the noncooperative model is as follows:
Step
1. Initialization. Give an initial guess of permit prices $\{m_{t}^{0}:\ t\geq 0\}$ and solve the cooperative model with emission cap and trade. The optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{i,t}^{0},E_{i,t}^{P,0}:\ t\geq 0\}$, for each region $i$. Iterate through steps 2, 3 and 4 for $j=0,1,...$, until convergence.
Step
2. Maximization Step. Solve the maximization problem (14) without the market clearing condition (12) for each region. Given the permit prices $\{m_{t}^{j}:\ t\geq 0\}$, we solve the optimization problem (14) without the market clearing condition (12) for region 1, where we fix the other regions’ net emissions at $\{E_{i,t}^{j}:\ i=2,...,12,\ t\geq 0\}$. We then obtain region 1’s optimal solution, where the optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{1,t}^{*},E_{1,t}^{P,j+1}:\ t\geq 0\}$ Similarly, we obtain $\{E_{i,t}^{*},E_{i,t}^{P,j+1}:\ t\geq 0\}$ for regions $i=2,...,12$.
Step
3. Update the emission permit prices and net emissions. After solving for the optimization problem of all regions, update the permit prices and net emissions as
$m_{t}^{j+1}$ $=$ $m_{t}^{j}\exp\left(\omega\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}\right),$
$E_{i,t}^{j+1}$ $=$ $\omega E_{i,t}^{*}+(1-\omega)E_{i,t}^{j},\ \ \forall i\in\mathcal{I},$
where $\omega=0.1$ is a weight parameter and $\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}$ is the net quantity of traded emission permits.
Step
4. Check the convergence criterion. Check if $m_{t}^{j+1}\simeq m_{t}^{j}$, $E_{i,t}^{j+1}\simeq E_{i,t}^{j}$, and $E_{i,t}^{P,j+1}\simeq E_{i,t}^{P,j}$ for every region $i$ and $t\geq 0$. If so, stop the iteration. Note that $m_{t}^{j+1}=m_{t}^{j}$ implies that the market clearing condition (12) holds at the solution.
This algorithm embodies the concept of market equilibrium. When there is a positive net quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices. Conversely, when there is a negative net quantity of traded emission permits, indicating excess supply, we lower the permit prices. This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears. Furthermore, this algorithm guarantees that each region obtains its optimal solution and reaches an equilibrium state. In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution. Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state.
4 SCC, MAC, and the Optimal Carbon Tax
In this section, we first define the regional social cost of carbon (SCC) and marginal abatement cost (MAC) and explain how the optimal carbon tax can be introduced together with the global ETS, in relation to the SCC and MAC.
4.1 SCC and MAC
The social cost of carbon (SCC), a central concept in the climate change literature, is widely used to quantify the monetary value of climate damages induced by an additional unit of carbon emissions. While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region. Similar to van der Ploeg and de Zeeuw (2016) and Cai et al. (2023), we define the cooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
$\mathrm{SCC}_{i,t}^{\mathrm{Coop}}=\frac{-1,000(\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t})}{\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}},$ (15)
for region $i$, where $V_{t}^{\mathrm{Coop}}$ is the value function of the cooperative model at time $t$, depending on the state variables $\{K_{i,t},\mathcal{E}_{t}:\ i\in\mathcal{I}\}$; that is,
$V_{t}^{\mathrm{Coop}}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{t}\sum_{i=1}^{12}u(c_{i,s})L_{i,s}.$
Similarly, the noncooperative SCC of a region is defined as
$\mathrm{SCC}_{i,t}^{\mathrm{OLNE}}=\frac{-1,000(\partial V_{i,t}^{\mathrm{OLNE}}/\partial\mathcal{E}_{t})}{\partial V_{i,t}^{\mathrm{OLNE}}/\partial K_{i,t}},$ (16)
where
$V_{i,t}^{\mathrm{OLNE}}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{t}u(c_{i,s})L_{i,s},$
for each region $i$ under the OLNE.
Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon. To compute the regional SCC in the cooperative or the noncooperative model, it is equivalent to replace the numerator in equation (15) or (16) with the shadow price of the transition equation of cumulative global emissions (6) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (11) for each region.
It should be emphasized that our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al. (2018). Note that, for the global SCC, the marginal rate of substitution between global emissions and global capital (or consumption) is equivalent to the present value of future global climate damages from an additional unit of global emissions released in the current period. In this regard, it has been well understood in the literature that the global SCC can be used as the optimal global carbon tax. However, in Nordhaus (2017) and Ricke et al. (2018), their regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period. Such definition implies that their regional SCC is the regional contribution to the global SCC, i.e., the sum of their regional SCCs over all regions is the global SCC. This does not make sense in the context of carbon tax.
We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the OLNE. With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional net emissions is not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional net emissions is binding at zero. This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al. (2017) and Cai and Lontzek (2019). Similarly, the regional cooperative SCC in our definition is equal to the optimal regional carbon tax under cooperation, when the regional net emissions is not binding at zero and the exogenous emission permit price is zero (or there is no ETS). However, when the exogenous emission permit price is nonzero, the regional cooperative SCC is not equal to the optimal regional carbon tax under cooperation, because the ETS can be used as a means for transferring capital between regions to improve the global welfare under cooperation. For instance, a developed region can purchase a large amount of permits from a developing region even when its net emissions has been below its emission cap.
The marginal abatement cost (MAC) is another important concept that captures the additional cost incurred due to an increase in emission abatement. From our model equation (2), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD. Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows
$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$
When there is no ETS, it is well known that the optimal carbon tax is equal to the MAC. This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019). However, in the presence of the ETS, the equality does not hold as firms can purchase or sell emission permits instead of abatement.
4.2 Optimal Carbon Tax under the ETS
In this subsection, we discuss the relationship between the MAC and carbon taxation under the global ETS. While carbon taxes and ETS have often been considered as substitutable policies, we show that each region can internally choose an optimal carbon tax under the global ETS as a complementary policy tool to obtain the optimal level of net emissions. Before we discuss the optimal carbon tax, the mechanism behind the emission permit buyer and seller regions needs to be understood. Here we drop the region and time indices for simplicity.
First of all, the relationship between the optimal net emission ($E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller. A region becomes a permit buyer if its optimal net emission is greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement. On the other hand, a region becomes a permit seller if its optimal net emission is lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$). Note that this is the optimal solution from the perspective of each region’s social planner.
One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emission level. We assume that the social planner uses carbon taxation as the policy instrument to fix the inconsistency between the optimal net emissions of the regional social planner and the firms. Let us consider a representative firm in each region that produces emissions and faces the MAC function and the emission cap of the region. We assume that the representative firm is myopic and maximizes its profit in each period, thus it does not consider the dynamic optimization problem of the social planner. Given the total emissions, $E^{\mathrm{Total}}$, the representative firm’s profit maximization problem is equivalent to choosing its optimal level of net emissions to minimize its total cost associated with emissions (i.e. the abatement cost and revenue or expenditures from emission permit purchases or sales). [^4] It is worth mentioning that a region’s total emissions is proportional to its GDP, which, in turn, is determined by capital and the exogenous path of population size. Considering that the current level of capital is determined in the previous period, total emissions can be considered as exogenously given from the perspective of the single-period myopic firm.
Figure 1 depicts the relationship between emissions, permit prices, the carbon tax, and the MAC for buyer and seller regions. In Figure 1, the MAC curve is a decreasing function of net emissions ($E$),  [^5] Substituting $E^{A}=E^{\mathrm{Total}}-E$ from equation (3) into the MAC, it follows that $\mathrm{MAC}_{i,t}=1,000\left(\frac{b_{1,i,t}b_{2,i}}{\sigma_{i,t}}\right)\left(\frac{E_{i,t}^{\mathrm{Total}}-E_{i,t}}{\sigma_{i,t}Q_{i,t}}\right)^{b_{2,i}-1}$. which reaches zero when net emissions equals total emissions ($E^{\mathrm{Total}}$). That is, the higher the emission abatement, the smaller the net emissions, resulting in a higher MAC. The market equilibrium price of emission permits is given at $m$. Then in the absence of a carbon tax, there is a corresponding cut-off level of net emissions, $E^{m},$ above which conducting the emission abatement is more cost-effective, and below which purchasing emission permits from the permit markets is cheaper.
Figure 1: The optimal carbon tax in emission permit buyer and seller regions.
In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$. Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits. That is, in the absence of a carbon tax, the social planner’s optimal net emission level, $E^{*}$, may not align with the optimal choice of firms within each region, $E^{m}$, regardless of whether the region is a permit buyer or seller. Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at $m$ to increase its profits. This is also verified in our simulation results later on.
We now show that a regional carbon tax can serve as a supplementary policy for the representative firm to reach the social planner’s optimal net emissions at $E^{*}$. Assume that the social planner levies a tax on the firm of $\tau$ per unit of net emissions. Given the total emissions, $E^{\mathrm{Total}}$, the cost minimization problem of the representative firm that faces this carbon tax under the global ETS is as follows:
[…]
The first term of the objective function is the total tax on net emissions, the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E^{\mathrm{Total}}-E)$. The above cost minimization problem holds for both emission permit buyer and sellers. For simplicity, we assume that emissions cannot be negative in our model.
From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
$\tau=\mathrm{MAC}(E^{*})-m,$ (18)
when $E^{*}>0$. Importantly, it implies that for the representative firm to attain the optimal net emission level $E^{*},$ the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$. In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC of a region and the market price of emission permits. It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*},$ rather than $E^{m}$. When $E^{*}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
$\tau\geq\mathrm{MAC}(E^{*})-m.$ (19)
That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits. In particular, the regional $\mathrm{MAC}(E^{*})$ at $E^{*}=0$ could be smaller than the global permit price $m$, in which case the optimal carbon tax should be zero. Since the optimal carbon tax can always be chosen to be equal to the gap between the MAC and the permit price when the gap is positive, we have the following general formula for any $E^{*}\geq 0$:
$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\}.$ (20)
The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model. Therefore, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
5 Data and Calibration
To calibrate our model with data, we have two objectives. The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their commitments to the Paris Agreement and the net zero emission targets. The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al., 2018; Ueckerdt et al., 2019; Kahn et al., 2021). We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
Our model’s 12 regions are formed by aggregating 198 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010). For the initial year 2020, we obtain country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (GtC) from the World Bank. For future projections on population growth, we use the SSP2 scenario (Samir and Lutz, 2017).
5.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade or ETS currently in place, we use the emission commitments of the Paris Agreement and the regions’ net zero commitments to construct the baseline emission cap constraint, $\overline{E}_{i,t}$. We collect reports of the Intended Nationally Determined Contributions (INDCs) of 165 countries that were submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 (INDC, 2023). Although there is no international enforcement mechanism, most countries committed to achieving a self-defined mitigation target for 2030 (or 2025). We also obtain the target years to reach net zero emissions of different countries from Climate Action Tracker (2023). Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
As the first step, we obtain the level of emission targets by 2030 for each country. In their INDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level in 2015. Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in net emissions. For the countries that did not make a specific emissions reduction pledge, we assume that their carbon intensity reduction and emission reduction percentages are the same as those of the most populous country in that region.
Next, we generate annual emission cap pathways based on the regions’ historical emission levels (World Bank, 2020), their emission targets for 2030 (or 2025), and net zero emission target years. We use five-year emissions data (2014 - 2018) from the World Bank and the INDC emission targets in 2030 (or 2025) to fit a quadratic function and use this fitted function to project the emission pathways for the periods between 2018 and 2030. Emission projections for the years between 2030 (or 2025) and the net zero emission target year are obtained by linearly interpolating the emissions. After obtaining the country-level emission pathways, we aggregate them to find the regional emission cap pathways. [^6] We also find that our aggregated regional emission cap pathways are close to those used in Nordhaus (2010). To assess the implications of varying stringency in the emission caps, we also create additional emission cap pathways by choosing alternative net zero scenarios, wherein we assume that all regions achieve net zero emissions by 2050 (most stringent), 2070, or 2090 (most lax). See Figures A.2 and A.3 in Appendix A.4 for the regional emission cap pathways in the baseline cap scenario and the different emission cap pathways at the global level.
5.2 Total Factor Productivity
We calibrate total factor productivity (TFP), $A_{i,t}$, for future periods based on the GDP projections in Burke et al. (2018), who provide the projected GDP of 165 countries until 2099 assuming no climate-related impacts. We aggregate these projections according to our 12 regions and employ the SSP2 population scenario to obtain regional GDP per capita estimates, $y_{i,t}^{\mathrm{BDD}}$, which are used to calibrate the regional TFP under no climate impacts, $A_{i,t}$. For each region $i$, the dynamic path of the TFP is modeled by the relationship $A_{i,t+1}=A_{i,t}\exp\left(g_{i,t}\right)$, where $g_{i,t}$ is the growth rate of $A_{i,t}$ at time $t$. When $t<80$ (i.e., within this century), we assume
$g_{i,t}=g_{i,0}\exp\left(-d_{i}t\right).$ (21)
For $t\geq 80$ (i.e., beyond this century), since the cumulative effect is huge for a long horizon, it is often inappropriate to simply extrapolate TFP growth rate using the formula (21). Therefore, we follow RICE to generate $g_{t,i}$ for $t\geq 80$. See Appendix A.5 for the details.
In our structural estimation, we obtain $(g_{i,0},d_{i})$ by solving the following minimization problem:
$\min_{g_{i,0},d_{i}}\ \sum_{t=0}^{79}w_{t}\left(\frac{y_{i,t}^{\mathrm{NoCC}}}{y_{i,0}^{\mathrm{NoCC}}}-\frac{y_{i,t}^{\mathrm{BDD}}}{y_{i,0}^{\mathrm{BDD}}}\right)^{2},$ (22)
where $w_{t}$ are weights, and $y_{i,t}^{\mathrm{NoCC}}$ is GDP per capita obtained by solving the following optimal growth model with a choice of $(g_{i,0},d_{i})$ and its associated TFP $A_{it}$:
$\max_{c_{i,t}}$ $\sum_{t=0}^{500}\beta^{t}u(c_{i,t})L_{i,t},$ (23)
$\mathrm{s.t.}$ $K_{i,t+1}=(1-\delta)K_{i,t}+\left(y_{i,t}^{\mathrm{NoCC}}-c_{i,t}\right)L_{i,t},$
where $y_{i,t}^{\mathrm{NoCC}}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{-\alpha}$ is the GDP per capita under no climate impact and $u(c_{i,t})$ is defined as in equation (7). The initial TFP $A_{i,0}$ is chosen such that $A_{i,0}K_{i,0}^{\alpha}L_{i,0}^{-\alpha}$ is equal to the observed GDP per capita in 2020. Figure A.4 in Appendix A.6 shows that with our calibrated $A_{i,t}$ , the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD}}$ from Burke et al. (2018) for all regions.
5.3 Carbon Intensity
To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al. (2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation. As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun211’ in Ueckerdt et al. (2019). Specifically, this FFrun211 scenario corresponds to climate action from 2020 with full technology portfolio but with zero carbon tax. Based on the equation (1), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun211}}^{U}/Q_{i,t,\mathrm{FFrun211}}$, where $E_{i,t,\mathrm{FFrun211}}^{U}$ and $Q_{i,t,\mathrm{FFrun211}}$ are the projected regional emissions and GDP under the FFrun211 scenario for region $i$ at time $t$.  [^7] Ueckerdt et al. (2019) provide data for 11 regions. Upon comparison, we find that the countries constituting the ‘Rest of the World (ROW)’ region are the ones that are in the ‘Other High Income (OHI)’ and ‘Eurasia’ region in our study. Therefore, the carbon intensity obtained from the ROW in Ueckerdt et al. (2019) corresponds to that of OHI and Eurasia regions in our work.
5.4 Abatement Cost
Our estimation of the abatement cost parameters $b_{1,i}$, $b_{2,i}$, and $b_{3,i}$ relies on the simulation results under ten different levels of carbon taxes in Ueckerdt et al. (2019). For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$, Ueckerdt et al. (2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$. Since Ueckerdt et al. (2019) do not consider an ETS, according to the discussion in Section 4.2, we can assume that their carbon taxes $\tau_{t,j}^{U}$ are equal to the marginal abatement costs when net emissions are strictly positive. That is,
$\tau_{t,j}^{U}=\frac{1,000b_{2,i}\mu_{i,t,j}^{b_{2,i}-1}b_{1,i,t}}{\sigma_{i,t}}=1,000b_{2,i}\mu_{i,t,j}^{b_{2,i}-1}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t)),$ (24)
for $\mu_{i,t,j}\in(0,1)$. Therefore, we can use their simulation results under different carbon tax levels to estimate $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$. At first, we estimate the associated emission control rate $\mu_{i,t,j}^{U}$ using the equations (1)-(3) as follows:
$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun211}}^{U}},$ (25)
where $E_{i,t,\mathrm{FFrun211}}^{U}$ is the regional emissions from the FFrun211 scenario with zero carbon tax (see Ueckerdt et al. (2019) for details). Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (24) can hold in an approximate manner for every scenario $j$ and time $t$.
5.5 Climate Damage
We calibrate the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$ by considering projections on GDP loss across different climate scenarios in Kahn et al. (2021), which shows the percentage loss in GDP per capita by 2030, 2050, and 2100 under the RCP 2.6 and RCP 8.5 scenarios for China, EU, India, Russia, and the US. We use their method and data to project the percentage loss in GDP per capita ($\Delta_{i,t}^{\mathrm{RCP26}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}$) every year from 2020 to 2114 under the RCP 2.6 and RCP 8.5 scenarios for each of our 12 regions, employing the baseline setup in Kahn et al. (2021). Specifically, $\Delta_{i,t}^{\mathrm{RCP26}}=1-y_{i,t}^{\mathrm{RCP26}}/y_{i,t}^{\mathrm{base}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}=1-y_{i,t}^{\mathrm{RCP85}}/y_{i,t}^{\mathrm{base}}$, where $y_{i,t}^{\mathrm{RCP26}}$, $y_{i,t}^{\mathrm{RCP85}}$, and $y_{i,t}^{\mathrm{base}}$ are GDP per capita under RCP 2.6, RCP 8.5, and the baseline scenario, respectively. Thus from equation (9) we obtain $(\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem for each region $i$:
$\min_{\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{94}\left(\frac{1+\pi_{1,i}T_{t}^{\mathrm{RCP85}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP85}}\right)^{2}}{1+\pi_{1,i}T_{t}^{\mathrm{RCP26}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP26}}\right)^{2}}-\frac{1-\Delta_{t,i}^{\mathrm{RCP26}}}{1-\Delta_{t,i}^{\mathrm{RCP85}}}\right)^{2}.$ (26)
Here $T_{t}^{\mathrm{RCP26}}$ and $T_{t}^{\mathrm{RCP85}}$ are the global average temperature anomalies at time $t$ (deviation from the pre-industrial temperature) under the RCP 2.6 and RCP 8.5 scenarios. Figure A.5 in Appendix A.7 shows that with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al. (2021), for every region.
6 Results
In this section, we discuss our simulation results. We first present the results of the noncooperative model under the baseline emission cap scenario for 2020 to 2100. Next, we compare the economic outcomes of different models: noncooperation versus cooperation, with or without the emission caps, and with or without the global ETS. Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all countries.
6.1 Noncooperation with ETS under the Baseline Emission Caps
Figure 2 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2032, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 2 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
Global emissions are constrained by the global emission caps starting from 2032. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volumes of both emission abatement and traded permits, along with their associated costs. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.2 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$749 per ton of carbon in 2050 and \$2,225 in 2069 (Figure 2 bottom-left). The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions. As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050 for the US, EU, Japan, and OHI regions, and 2060 for Russia and China. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
Figure 2: Simulation results at the global scale for the noncooperative model with emission trading under the baseline emission cap scenario.
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^8] In the bottom-left panel of Figure 2, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.74 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right). Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 3 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage. After 2032, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Middle East, Africa, and OHI regions, with the US and India being the largest permit buyers. The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider. Japan is expected to be involved in a relatively small volume of permit trading.
Figure 3: Simulation results of regional emission permit purchase for the noncooperative model with emission trading, under the baseline emission cap scenario.
Based on our simulation results, we now discuss the optimal regional carbon tax under the noncooperative model with the ETS and the baseline emission cap scenario. Figure 4 shows the regional MAC along with the market equilibrium price of emission permits. We find that the MAC of every region increases drastically and remains strictly greater than the permit price until 2050, and the net emissions are strictly positive for all regions during this time.  [^9] For regional optimal net emissions, see Figure 3 in Appendix A.8.1. As discussed in Section 4.2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ is equivalent to $E_{i,t}^{*}<E_{i,t}^{m}$, implying that when there is no carbon tax, the optimal net emissions for the social planner do not align with the optimal net emissions for the representative firm in each region. Therefore, the optimal carbon tax can be determined by $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s net emissions are equal to the social planner’s optimal net emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050.
Figure 4: Simulation results of regional MAC for the noncooperative model with emission trading under the baseline emission cap scenario.
In Figure 4, the MAC of each region gradually decreases below the permit price after its net zero emissions (before trading of permits) is achieved. Russia is the first region to achieve net zero emissions (before trading of permits) in 2051 (as shown in Figure A.6 in Appendix A.8.1). But Russia still sells permits afterwards, as shown in Figure 3, because its after-trade net zero emission target year is 2060 under the baseline emission cap scenario. After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound. Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China and Eurasia) and the 2060s (the US, the EU, Japan, MidEast, Latin America, and the OHI). India, Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. [^10] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Recall that, when the regional net zero emissions (before trading of permits) are achieved, the optimal carbon tax should be larger than the gap between MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding, then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
Figure 5: Simulation results of regional SCC for the noncooperative model with emission trading under the baseline emission cap scenario.
We further confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4. For example, in 2050, the MAC for the US is \$908 per ton of carbon, the permit price is \$749, and their difference is exactly equal to the regional SCC of the US, \$159. Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC, starting at \$105 per ton of carbon in 2020 and steadily increasing up to \$210 per ton of carbon by the end of this century. The US is followed by the EU (from \$42 to \$138 per ton of carbon) and China (from \$39 to \$115 per ton of carbon) in this century. The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon). Our results show that there is considerable heterogeneity in the SCC across regions and suggest that a significant amount of regional carbon tax implementation is required, particularly for the US, the EU, and China, under the global ETS with the baseline emission cap scenario.
6.2 Comparative Analysis of Different Cases
We now modify our baseline model with various cases to understand the implications of cooperation among regions and the ETS. We compare outcomes between the noncooperative and the cooperative models, and for each model, we examine the impact of the ETS under the baseline emission caps. Figure 6 displays the global emission and temperature time profiles for all the cases. Note that for the cooperative model with the ETS, we use the equilibrium solution of permit prices obtained under the noncooperative model with the ETS, so we can compare the difference in the solutions between cooperation and noncooperation, holding permit prices fixed across scenarios.
Figure 6: Simulation results of different cases under cooperation or noncooperation.
We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation. Specifically, cooperation results in approximately 0.19 degrees Celsius lower temperature with the ETS and 0.15 degrees Celsius lower temperature without the ETS. Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps.
When comparing the cases with and without the ETS, we find that the ETS may play different roles under noncooperation and cooperation. In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS. This is because with the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level. This happens until 2050. In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time. With the ETS, slightly higher levels of global emissions and temperature rise are observed before 2040 and 2050, respectively, and the opposite trends are observed after that. This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
We also compare the regional MAC and SCC across the different cases. Figure 7 presents the regional MAC and SCC under both the cooperative and the noncooperative models with or without the ETS, using the US as an illustrative example. [^11] For comparisons of other regions, refer to Appendices A.8.4 and A.8.3. Under cooperation, the MAC of the US (Figure 7 top-left) is very large in the initial years. This is because the US, as a developed region, needs to mitigate a significant amount of emissions so that developing regions can mitigate less (as shown in Figure A.9 in Appendix A.8.3), thus improving global social welfare. After the 2030s, however, the US net emissions become zero in the cooperative case [^12] For regional net emissions under cooperation and noncooperation, see Appendix A.8.2. and cannot be further mitigated, so its MAC starts to decline over time. The effect of the ETS depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU at around 2040); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm).
Figure 7: Simulation results of different cases for the US in the cooperative or noncooperative model.
The top-right panel of Figure 7 shows that the cooperative SCC of the US is sustained at a high level throughout all periods. This is consistent with the high MAC of the US. When the US regional net emissions are not binding at zero before 2032 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2032. Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions. This also implies that a developed region will contribute more to mitigating emissions under cooperation. This pattern is consistent across regions: Figure A.11 in Appendix A.8.4 shows that the US, EU, Japan and OHI have a high cooperative SCC, while the developing regions like India and Africa have a small cooperative SCC.
Under noncooperation, the MAC of the US is substantially lower in the case with the ETS compared to the case without the ETS (Figure 7, bottom-left panel). Such discrepancy persists until 2058, the period during which the US is a net permit buyer (shown in Figure 3), reflecting the fact that implementation of the ETS offers the US a more cost-effective means of keeping its net emissions below its cap. Consequently, the US conducts less emission abatement. For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2032 (see Figure A.10 in Appendix A.8.3).
The bottom-right panel of Figure 7 shows that the noncooperative SCC of the US is much smaller than its cooperative SCC, which is also consistent with the smaller MAC of the US under noncooperation. Moreover, every region has a much smaller SCC under noncooperation than under cooperation, as shown in Figures A.11 and A.12 in Appendix A.8.4. This result implies that, in the dynamic Nash equilibrium under noncooperation, no region is willing to impose a high carbon tax thereby allowing other regions to free ride on lower emissions and climate damages. Therefore, in the noncooperative model, emissions and temperature rises are much larger than in the cooperative model, as shown in Figure 6.
Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS. For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2032 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2032 (see the two top panels of Figure 7), which is equal to the optimal regional carbon tax under cooperation. For the noncooperative case without the ETS, the US regional net emissions are binding at their caps (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC (see the two bottom panels of Figure 7), which is equal to the optimal regional carbon tax under noncooperation. Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Russian under cooperation until 2057 in Figures A.7, A.9, A.11); and Russian under noncooperation until 2042 in Figures A.8, A.10, and A.12). This relationship holds for every region as shown in Figures A.7-A.12.
The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under both cooperation and noncooperation. This pattern holds for every region as shown in Figures A.11 and A.12.
6.3 Alternative Net Zero Scenarios
Along with the baseline emission cap scenario, we further analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090. The top-left and top-right panels in Figure 8 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,538 per ton of carbon in 2049, and the temperature rise is restricted to 1.64 degrees Celsius by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$344 and \$158 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.86 degrees Celsius and 2.08 degrees Celsius above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
Figure 8: Simulation results of alternative emission cap scenarios under noncooperation with the ETS.
The bottom-left and bottom-right panels in Figure 8 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.8.5, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2086 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$216 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$193 in 2100 in the net zero 2090 scenario. This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
7 Conclusion
In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system. In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the Paris Agreement and their net zero commitments. We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting. The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth. We further show that carbon taxation can be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal net emissions. For strictly positive net emissions, we have shown both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
This work has several policy implications. First, our results indicate that the current global target of restricting the global temperature rise to 1.5 degrees Celsius above pre-industrial levels by 2100 is unattainable under noncooperation, even with the current emission commitments outlined in the Paris Agreement. Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature. Second, our baseline simulation shows that the current emission commitments of the Paris Agreement lead to excess emission permit supply in the initial decade, resulting in permit prices of zero. This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits. Third, we demonstrate that the global ETS is not a pure substitute for the carbon tax, but instead that both can be implemented together to improve the efficiency of global climate policies.
Appendix for Online Publication
A.1 List of Parameters
Table A.1: Key parameters.
Parameter Description
Economic parameters
$\beta=0.985$
Discount factor
$\gamma=1.45$
Elasticity of marginal utility
$\alpha=0.3$
Output elasticity of capital
$\delta=0.1$
Annual depreciation rate of capital
Climate parameters
$\zeta=0.0021$
Contribution rate of carbon emissions to temperature
$\sigma_{i,t}$
Carbon intensity (GtC per trillion \$)
$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$
Abatement cost parameters
$\pi_{1,i}$, $\pi_{2,i}$
Climate damage parameters
TFP parameters
$g_{i,0},g_{i,t}$
Growth rate of TFP
$d_{i}$
Parameters of TFP growth
A.2 List of Countries
Table A.2: List of countries for regional aggregation.
Region Constituent Countries
Africa
Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
European Union [^13] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience.
Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
Eurasia
Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
Latin America
Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St. Kitts and Nevis, St. Lucia, St. Vincent and the Grenadines, Suriname, Trinidad and Tobago, Uruguay, Bolivarian Republic of Venezuela, and the U.S. Virgin Islands.
Region Constituent Countries
Middle East
Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
Other Non-OECD Asia
Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
Other High-Income
Australia, Canada, Republic of Korea, New Zealand, Singapore.
A.3 Calibration of the TCRE Climate System
Each of the four RCP scenarios (Meinshausen et al., 2011) — RCP 2.6, RCP 4.5, RCP 6, and RCP 8.5 — provide their pathways of emissions, atmospheric carbon concentration, radiative forcing, and atmospheric temperature anomaly. When we calibrate the contribution rate of carbon emissions on temperature, $\zeta$, in a climate system, we use the pathways of emissions and atmospheric temperature anomaly of the four RCP scenarios. Figure A.1 shows that our calibrated TCRE climate system provides a very good projection of the atmospheric temperature anomaly (increase relative to pre-industrial levels) based on cumulative emissions only.
Figure A.1: Calibration of the TCRE climate system.
A.4 Emission Cap Pathways
Figure A.2 shows the regional emission cap pathways generated using the methodology described in Section 5.1. Figure A.3 compares the different emission cap pathways at the global level.
Figure A.2: Regional emission cap pathways under the baseline scenario.
Figure A.3: Global emission cap pathways for the different net zero scenarios.
A.5 GDP Growth Rate beyond this Century
For the GDP growth rate beyond this century, we follow RICE to project $g_{t,i}$ for $t\geq 80$. We begin by assuming the long-run growth rate of TFP in the US is $g_{\mathrm{US},\infty}=0.0033(1-\alpha)=0.00231$ with $\alpha=0.3$. Next, we let $\widetilde{y}_{i,79}=y_{i,79}^{\mathrm{BDD}}y_{i,0}/y_{i,0}^{\mathrm{BDD}}$ be our projected per capita output in 2099, where $y_{i,0}$ is the observed per capita output in 2020. We then assume that the TFP growth in the US is characterized by
$g_{\mathrm{US},t}=g_{\mathrm{US},\infty}+(g_{\mathrm{US},79}-g_{\mathrm{US},\infty})\exp(-0.01(t-79)),$
and let $\widetilde{y}_{\mathrm{US},t+1}=\widetilde{y}_{\mathrm{US},t}\exp(g_{\mathrm{US},t}/(1-\alpha))$ for $t\geq 79$. For the regions other than the US, we assume their TFP growth can be expressed in relation to the TFP growth of the US. Specifically, we assume that, for $t\geq 79$,
$\begin{cases}\widetilde{y}_{i,t+1}=\widetilde{y}_{i,t}\exp(g_{i,t}/(1-\alpha))\\ g_{i,t+1}=g_{\mathrm{US},t+1}+(1-\alpha)\chi\ln(\widetilde{y}_{\mathrm{US},t}/\widetilde{y}_{i,t})\end{cases}$
where $\chi=0.005$ is chosen such that $g_{i,t}$ gradually moves toward $g_{\mathrm{US},t}$ as $t\rightarrow\infty$.  [^14] Assume $\widetilde{y}_{i,t}=A_{i,t}k_{i,t}^{\alpha}$ is GDP per capita where $k_{i,t}$ is capital per capita. We have $\ln\left(\frac{\widetilde{y}_{i,t+1}}{\widetilde{y}_{i,t}}\right)=g_{i,t}+\alpha\ln\left(\frac{k_{i,t+1}}{k_{i,t}}\right).$ If we assume the growth of $k_{t,i}$ is equal to the growth of GDP per capita, then we have $\widetilde{y}_{i,t+1}=\widetilde{y}_{i,t}\exp(g_{i,t}/(1-\alpha)).$
A.6 Calibration of Total Factor Productivity
Figure A.4 shows that with our calibrated $A_{i,t}$ , the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD}}$ from Burke et al. (2018) for all regions.
Figure A.4: Fitting GDP per capita under no climate impact. Lines represent GDP per capita from Burke et al. (2018); marks represent fitted GDP per capita.
A.7 Calibration of Climate Damage
Figure A.5 shows that with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al. (2021), for every region.
Figure A.5: Fitting climate damage parameters. Lines represent the ratios of GDP per capita between RCP 2.6 and RCP8.5 from Kahn et al. (2021); marks represent fitted ratios.
A.8 Additional Simulation Results
A.8.1 Regional Net Emissions under Noncooperation
Figure A.6 displays the regional net emissions under the noncooperative model with the ETS and the baseline emission cap scenario. Latin America is the first to reach net emissions in 2053, followed by Russia in 2055, the US in 2056, OHI in 2057, and Eurasia in 2058. Then, net emissions are achieved by China in 2061, Japan in 2062, the MidWest in 2064, and the EU in 2065. Finally, India achieves net emissions in 2069, followed by Africa and other non-OECD Asian countries in 2071.
Figure A.6: Simulation results of regional net emissions under noncooperation with the ETS and baseline emission caps.
A.8.2 Regional Net Emissions Comparison between Different Cases
Figure A.7 compares regional net emissions between two cases under cooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.7: Simulation results of regional net emissions: comparison between two cases under cooperation.
Figure A.8 compares regional net emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
Figure A.8: Simulation results of regional net emissions: comparison between two cases under noncooperation.
A.8.3 Regional MAC Comparison between Different Cases
Figure A.9 compares the regional MAC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
Figure A.9: Simulation results of regional MAC: comparison between two cases under cooperation.
Figure A.10 compares the regional MAC under noncooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
Figure A.10: Simulation results of regional MAC: comparison between two cases under noncooperation.
A.8.4 Regional SCC Comparison between Different Cases
Figure A.11 compares the regional SCC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
Figure A.11: Simulation results of the regional SCC: comparison between two cases under cooperation.
Figure A.12 compares the regional SCC under noncooperation with the baseline emission caps, comparing two cases: (i) with the ETS, (ii) without the ETS.
Figure A.12: Simulation results of the regional SCC: comparison between two cases under noncooperation.
A.8.5 Regional MAC and SCC under Alternative Emission Caps
In Figure A.13, we compare the MAC for the noncooperative model with the ETS across alternative emission cap scenarios.
Figure A.13: Simulation results of regional MAC: comparison across alternative emission cap scenarios.
Similarly, Figure A.14 displays the SCC of the noncooperative model with the ETS across different emission cap scenarios.
Figure A.14: Simulation results of regional SCC: comparison across alternative emission cap scenarios.

### v2

@@
 We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world.
-We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments outlined in the 2015 Paris Agreement.
-Our model shows that the emission permit price reaches \$749 per ton of carbon by 2050.
+We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and updated in subsequent years.
+Our model shows that the emission permit price reaches \$811 per ton of carbon by 2050.
 We demonstrate that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
@@
 1 Introduction
-Anthropogenic warming of our planet is projected to have disastrous consequences, from rising sea levels to increases in the frequency and intensity of extreme weather events (Arias et al., 2021).
-The first ever international effort to mitigate global emissions was undertaken in 1997 with the signing of the Kyoto protocol.
-More than two decades since the Kyoto protocol, global efforts to tackle climate change have failed, with the world experiencing accelerating rates of greenhouse gas emissions.
-The policy outlined in the 2015 Paris Agreement aimed to codify an objective of limiting climate change to a maximum of 2 degrees Celsius above pre-industrial levels.
-To accomplish this objective, the participating nations committed to Intended Nationally Determined Contributions (INDCs) specifying their emission mitigation goals.
-In November 2021, representatives of almost 200 countries met at Glasgow to renew their efforts in curbing climate change and approved the Glasgow Climate Pact with the overall aim of limiting the rise in global temperatures to 1.5 degrees Celsius through numerous actions.
-Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETS).
+Climate change has been raising concerns of disastrous consequences, from rising sea levels to increases in the frequency and intensity of extreme weather events (Arias et al., 2021).
+International efforts to address climate change have evolved over the past three decades, commencing with the United Nations Framework Convention on Climate Change (UNFCCC) in 1992, followed by the Kyoto Protocol in 1997, the Paris Agreement in 2015, and the Glasgow Pact in 2021.
+The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2 degrees Celsius above pre-industrial levels, with a pursuit of limiting it to the 1.5 degrees Celsius” (Article 2).
+To accomplish this objective, 194 countries (parties) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures.
+Most recently, NDCs have been revisited in the Glasgow Climate Pact in November 2021, reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more progressive efforts to attain the global target of 1.5 degrees Celsius.
+Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETSs).
+[^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for future implementation around the world (Stavins, 2022).
 In a carbon tax regime, a tax is charged on firms for each unit of carbon emissions.
-The optimal global carbon tax rate is often determined at the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus, 2017).
-On the other hand, in an emission trading or a cap-and-trade system, the maximum amount of emissions is fixed in an economy and agents can sell and purchase emission permits at the market price.
-Under the assumptions of complete information and no frictions, properly designed carbon tax regimes and emission trading systems are argued to have equivalent outcomes (Nordhaus and Yang, 1996; Goulder and Schein, 2013).
-The ETS, however, is often considered a more efficient policy tool in practice, since it encourages agents that can reduce emissions at lower costs than others to undertake more emission reduction efforts (Nordhaus, 2017).
-In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, and as of the summer of 2023, the permit price stands at about 90 euros per ton of carbon.
-[^1] Source of EU permit price: https://tradingeconomics.com/commodity/carbon.
-[^2] In addition to the EU ETS, other national or sub-national ETS have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States.
-For a review of emission cap-and-trade systems in practice, see e.g., Schmalensee and Stavins (2017).
-Although a global-level ETS is not currently in place, it has emerged as a prominent potential mechanism to combat climate change, particularly since Article 17 of the Kyoto Protocol proposed emissions trading among countries.
+Previous studies have shown that the optimal global carbon tax rate is determined at the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus, 2017).
+In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price.
+In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery, 1972; Goulder and Schein, 2013; Stavins, 2022).
+In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, [^2] In addition to the EU ETS, other national or sub-national ETSs have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States.
+For the information on most up-to-date policy in practice, see World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/).
+and as of the summer of 2023, the permit price stands at about 90 euros per ton of Carbon Dioxide (CO2).
+[^3] Source of EU permit price: (https://tradingeconomics.com/commodity/carbon).
+Although a global-level ETS is currently not in place, there have been longstanding discussions on its implementation as an international regime to address climate change, as implied by Article 17 of the Kyoto Protocol and Article 6 of the Paris Agreement.
 In this study, we develop a dynamic multi-region model of climate and the economy under a hypothetical global ETS, and use it to project global prices of emission permits and the regional Social Cost of Carbon (SCC).
@@
 Our study makes four contributions.
-The first contribution of this study is a method to solve for the market equilibrium prices of emission permits under a global ETS.
-To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with endogenous prices in a dynamic regional integrated assessment model.
-While prior research on the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus, 2010) highlighted the importance of a multi-region ETS, it did not provide a solution for the market-clearing equilibrium permit price in a noncooperative setting.
-We present an algorithm for finding a Nash equilibrium solution in this context.
-Under the baseline emission cap scenario based on the Paris Agreement and the regional net zero commitments, our simulation results show that the oversupply of global permits in the initial decade results in zero permit prices, but by 2050, the emission permit price can reach up to \$749 per ton of carbon.
-Furthermore, we show that regions such as the United States, India and the EU become major emission permit buyers, and other regions such as Russia and China become permit suppliers.
-The second contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS.
-While carbon taxes and ETS have often been considered as potentially interchangeable policies, we show that within the global ETS framework, the social planner of each region can introduce a carbon tax as a complementary policy tool to achieve its optimal level of net emissions.
-The necessity of a carbon tax arises from the discrepancy in optimal net emissions between the social planner’s choice and the market outcome under the ETS.
-When optimal net emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional MAC and the global permit price under the noncooperative Nash equilibrium with the ETS.
+The first contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS.
+While carbon tax and ETS have often been considered as potentially interchangeable policies, we show that within the global ETS framework, it is optimal for the social planner of each region to introduce a carbon tax as a complementary policy tool to achieve its optimal level of net emissions.
+That is, a hybrid policy with both ETS and carbon tax is optimal.
+The necessity of a carbon tax arises from the discrepancy in optimal net emissions between the social planner’s choice and the market outcome under the ETS only (i.e., firms’ choice under no carbon tax).
+When optimal net emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional MAC and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with the social planner’s.
 We also numerically verify that the regional optimal carbon tax is equal to the regional SCC when net emissions are strictly positive.
-The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Paris Agreement, enhancing the accuracy and relevance of our findings.
+The second contribution of this study is a method to solve for the dynamic Nash equilibrium prices of emission permits under a global ETS.
+To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with fully endogenous prices and emission abatements in a dynamic regional integrated assessment model (IAM) under noncooperation.
+Although some prior research has addressed regional integrated assessment models with a multi-region ETS, their solutions thus far have been obtained in a limited manner.
+One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang, 1996; Nordhaus, 2010), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation.
+Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al., 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni, 2012; Massetti and Tavoni, 2012).
+While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane, 2009; Nordhaus, 2010), our study shows that this does not hold.
+Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between forward-looking regional social planners’ and myopic firms’ choices under the ETS and no carbon tax.
+This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for every region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation with the assumption.
+However, it is well known that even without emission caps, there are still nonzero emission mitigations and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
+We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and, importantly, show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions.
+Under the baseline emission cap scenario based on the Glasgow Pact and the regional net zero commitments, our simulation results show that the oversupply of global permits in the initial decade results in zero permit prices, but by 2050, the emission permit price can reach up to \$811 per ton of carbon.
+Our results further show that regions such as the United States, India and the EU become major emission permit buyers, and other regions such as Russia and China become permit suppliers.
+The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Glasgow Pact, rendering our findings relevant for the real world scenario.
 We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz, 2017), also known as the “Middle-of-the-Road” scenario.
@@
 We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al., 2011).
-Due to its simplicity and effectiveness, the TCRE scheme has found application in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, 2017; Dietz and Venmans, 2019; Mattauch et al., 2020; Barnett et al., 2020.
+Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, 2017; Dietz and Venmans, 2019; Mattauch et al., 2020; Barnett et al., 2020.
 Moreover, Dietz et al.
@@
 Moreover, we show that a stricter emission cap leads to a higher permit price and smaller regional SCCs for every region.
-For example, if all the regions achieve net zero emissions in 2050, the emission permit price can reach up to \$1,538 per ton of carbon by 2049.
-However, even with this strict emission restriction, the global temperature is still projected to rise by 1.64 degrees Celsius by the end of this century.
-These findings suggest that achieving the global target of limiting the temperature rise to 1.5 degrees Celsius requires even stronger emission commitments than those outlined in the Paris Agreement.
+For example, if all the regions achieve net zero emissions in 2050, the emission permit price can reach up to \$1,593 per ton of carbon by 2049.
+However, even with this strict emission restriction, the global temperature is still projected to rise by 1.61 degrees Celsius by the end of this century.
+These findings suggest that achieving the global target of limiting the temperature rise to 1.5 degrees Celsius requires even stronger emission commitments than those outlined in the Glasgow Pact.
+Our results are in qualitative agreement with the analysis of Rogelj et al.
+(2016), which found that the first NDC targets of the countries would result in a median warming of 2.1 to 3.1 degrees Celsius by 2100, and therefore stronger measures are needed to keep the warming under 1.5 degrees Celsius above the pre-industrial levels.
+Riahi et al.
+(2015) and Spencer et al.
+(2015) have similarly concluded that 2030 mitigation targets are insufficient to reach the greenhouse gas stabilization levels of 450 ppm.
 We also find that under cooperation between the regions, the regional SCC is higher for a developed region compared to a developing one.
 That is, there are higher carbon taxes in developed regions under cooperation compared to developing regions because the marginal utility of consumption is higher in developing regions.
-Global aggregate welfare will be higher if developed regions can contribute more to mitigating emissions, thus reducing climate damages in developing regions.
+The global aggregate welfare will be higher if developed regions can contribute more to mitigating emissions, thus reducing climate damages in developing regions.
 In the noncooperative setting, the regional SCCs of the developed and the developing regions are comparable.
-In the literature, the ETS has garnered attention as a promising mechanism for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze the ETS as a potential tool for global climate policy.
-One closely relevant study is Carbone et al.
-(2009), which builds a computable general equilibrium model of the world, incorporating countries’ endogenous participation in an ETS and allocation of emission permits, and solves for the equilibrium permit price.
-In contrast to our work, their study is based on a static setting, so it does not address the critical dynamic aspects of climate change, and lacks a connection with emission abatement decisions.
-Fischer and Springborn (2011) is another relevant work that constructs a dynamic stochastic general equilibrium model to compare the outcomes of the emission cap and tax policies, but their model does not consider emission trading between regions.
-More broadly, a series of recent studies have examined ETS from various perspectives.
-Some studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS or China ETS (Fuss et al., 2018; Hitzemann and Uhrig-Homburg, 2018; Borenstein et al., 2019; Goulder et al., 2022; Perino et al., 2022).
-Others analyze the efficiency gains from integrating regional ETS (Habla and Winkler, 2018; Doda et al., 2019; Holtsmark and Weitzman, 2020; Holtsmark and Midttømme, 2021), or compare the emission permit allocation methods between auctioning and free allocation (Goulder et al., 2010).
-While a global ETS is not yet in place, it is expected to offer multiple benefits compared to a global carbon tax.
-In an international ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al., 2016).
-An international ETS requires little governmental interference, and will help regions to maintain their emission target commitments and thereby contribute towards controlling the global temperature rise.
-On the other hand, a uniform global carbon tax regime is practically challenging.
-This paper therefore contributes to the literature by proposing a global ETS with endogenous permit prices and associated regional carbon taxes under various emission caps.
-Our model framework is closely related to the recent developments in integrated assessment models (IAMs), which aim to capture the interaction between economic growth and climate systems under cooperation and noncooperation.
-A seminal work is the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010), which extends the global DICE model (Nordhaus, 2014) by building a multi-region dynamic framework.
-Nordhaus (2010) computes carbon prices under a cooperative model where global welfare is maximized with weights applied on regional utilities.
-Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region context under cooperation and noncooperation.
-For instance, Hambel et al.
+Our model framework is closely related to the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010), an extension of the global DICE model (Nordhaus, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems.
+In Nordhaus (2010), carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities.
+Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region framework under cooperation and noncooperation.
+For instance, Luderer et al.
+(2012) and Jakob et al.
+(2012) compare the long-term predictions of three regionalized energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm.
+However, these works do not incorporate actual NDC and net-zero year targets of the countries to study the region-wise emission-economy pathways.
+van der Ploeg and de Zeeuw (2016) and Jaakkola and Van der Ploeg (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation.
+Hambel et al.
 (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions.
-van der Ploeg and de Zeeuw (2016) and Jaakkola and Van der Ploeg (2019) focus on stochastic components of the model, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation.
+Iverson and Karp (2021) study a Markov Perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates.
 Cai et al.
@@
 Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw (2016) and Cai et al.
-(2023) and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
+(2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
+In the literature, ETSs have garnered attention as promising mechanisms for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze ETSs as potential instruments for global climate policy.
+Among the most closely related studies are extended versions of the WITCH model (Bosetti et al., 2006): Massetti and Tavoni (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost.
+As noted earlier, our solution mechanism differs from these studies and provides permit prices from a complete noncooperative dynamic optimization problem.
+Carbone et al.
+(2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price.
+However, that study examines the ETS in a static setting, missing the critical dynamic aspects of climate change and lacking a connection with emission abatement decisions.
+Fischer and Springborn (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions.
+Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler, 2018; Doda et al., 2019; Holtsmark and Weitzman, 2020; Holtsmark and Midttømme, 2021).
+More broadly, a substantial body of literature have examined ETSs from various perspectives.
+While in this study, we assume that emission permits are freely allocated following the voluntary commitments from the Paris Agreement, few studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry, 2008; Goulder et al., 2010).
+Some have studied the relative efficiency of carbon policies under the existence of cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions and vice versa (Weitzman, 1974; Stavins, 1996; Karp and Traeger, 2024).
+Some studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg, 2018; Fuss et al., 2018; Perino et al., 2022), China ETS (Goulder et al., 2022), or California ETS (Borenstein et al., 2019).
+An important issue with regional ETSs is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS.
+Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al., 2024; Böhringer et al., 2014; Farrokhi and Lashkaripour, 2021; Fowlie and Reguant, 2022; Levinson, 2023).
+While there is no general agreement among economists on whether carbon tax or ETS is better (Stavins, 2022), the literature have often considered two pricing instruments (trading versus taxes) as policy substitutes.
+For instance, some studies support carbon tax over ETS, as ETS is subject to price volatility (Nordhaus, 2007).
+Conversely, other studies advocate ETS over tax because allocation of permits in ETS allows more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane, 2009).
+ETS could be a preferred instrument over tax in practice, as free allowances could be negotiated among participating agents to redistribute burden and, thus, be used to gain political support (Harstad and Eskeland, 2010; Hahn and Stavins, 2011; Stavins, 2022).
+Newell and Pizer (2003) show that ETS is preferable over carbon tax in the long run, and Knopf et al.
+(2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn, 2008).
+Harstad and Eskeland (2010) argue that an ETS in a perfect market is the first-best system, but frequent government interventions to redistribute allocations among firms results in distortions in the market allocation.
+In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al., 2016).
+For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand (2013); Schmalensee and Stavins (2017); Cai (2021); Stavins (2022).
+This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary.
+A global ETS necessitates regional carbon taxes.
 2 A Dynamic Regional Model of Climate and the Economy with an ETS
@@
 The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption.
-The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement and net zero emissions targets.
+The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets.
+In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
 The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model.
@@
 (2019).
-The amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$.
+We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$.
+[^5] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al., 2006; Bauer et al., 2012; Golosov et al., 2014; Baldwin et al., 2020).
+However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated.
+The emission control rate concept simplifies this.
 That is,
@@
 Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state.
+The concept of the OLNE assumes the regions enter an agreement to commit to the solution path over time at the initial period.
+While this type of equilibrium concept might not be as satisfactory as the feedback Nash equilibrium concept, it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
 4 SCC, MAC, and the Optimal Carbon Tax
@@
 To compute the regional SCC in the cooperative or the noncooperative model, it is equivalent to replace the numerator in equation (15) or (16) with the shadow price of the transition equation of cumulative global emissions (6) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (11) for each region.
-It should be emphasized that our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al.
-(2018).
-Note that, for the global SCC, the marginal rate of substitution between global emissions and global capital (or consumption) is equivalent to the present value of future global climate damages from an additional unit of global emissions released in the current period.
-In this regard, it has been well understood in the literature that the global SCC can be used as the optimal global carbon tax.
-However, in Nordhaus (2017) and Ricke et al.
-(2018), their regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period.
-Such definition implies that their regional SCC is the regional contribution to the global SCC, i.e., the sum of their regional SCCs over all regions is the global SCC.
-This does not make sense in the context of carbon tax.
+Our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al.
+(2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous.
 We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the OLNE.
@@
 (2017) and Cai and Lontzek (2019).
+Moreover, we also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, but our social discount rates are endogenous.
 Similarly, the regional cooperative SCC in our definition is equal to the optimal regional carbon tax under cooperation, when the regional net emissions is not binding at zero and the exogenous emission permit price is zero (or there is no ETS).
@@
 For instance, a developed region can purchase a large amount of permits from a developing region even when its net emissions has been below its emission cap.
+In addition, the regional cooperative SCC is not equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, because a developed region will abate more with a high carbon tax or cooperative SCC to improve the global welfare under cooperation.
 The marginal abatement cost (MAC) is another important concept that captures the additional cost incurred due to an increase in emission abatement.
@@
 $\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$
-When there is no ETS, it is well known that the optimal carbon tax is equal to the MAC.
-This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019).
-However, in the presence of the ETS, the equality does not hold as firms can purchase or sell emission permits instead of abatement.
 4.2 Optimal Carbon Tax under the ETS
@@
 Here we drop the region and time indices for simplicity.
-First of all, the relationship between the optimal net emission ($E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller.
-A region becomes a permit buyer if its optimal net emission is greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement.
-On the other hand, a region becomes a permit seller if its optimal net emission is lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$).
+First of all, the relationship between the regional social planner’s optimal net emissions (denoted $E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller.
+A region becomes a permit buyer if its optimal net emissions are greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement.
+On the other hand, a region becomes a permit seller if its optimal net emissions are lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$).
 Note that this is the optimal solution from the perspective of each region’s social planner.
-One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emission level.
+One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emissions level.
 We assume that the social planner uses carbon taxation as the policy instrument to fix the inconsistency between the optimal net emissions of the regional social planner and the firms.
@@
 Considering that the current level of capital is determined in the previous period, total emissions can be considered as exogenously given from the perspective of the single-period myopic firm.
-Figure 1 depicts the relationship between emissions, permit prices, the carbon tax, and the MAC for buyer and seller regions.
+Figure 1 depicts the relationship between emissions, permit price, carbon tax, and the MAC for buyer and seller regions.
 In Figure 1, the MAC curve is a decreasing function of net emissions ($E$),  [^7] Substituting $E^{A}=E^{\mathrm{Total}}-E$ from equation (3) into the MAC, it follows that $\mathrm{MAC}_{i,t}=1,000\left(\frac{b_{1,i,t}b_{2,i}}{\sigma_{i,t}}\right)\left(\frac{E_{i,t}^{\mathrm{Total}}-E_{i,t}}{\sigma_{i,t}Q_{i,t}}\right)^{b_{2,i}-1}$.
-which reaches zero when net emissions equals total emissions ($E^{\mathrm{Total}}$).
+which reach zero when net emissions equal total emissions ($E^{\mathrm{Total}}$).
 That is, the higher the emission abatement, the smaller the net emissions, resulting in a higher MAC.
@@
 Figure 1: The optimal carbon tax in emission permit buyer and seller regions.
-In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$.
-Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits.
-That is, in the absence of a carbon tax, the social planner’s optimal net emission level, $E^{*}$, may not align with the optimal choice of firms within each region, $E^{m}$, regardless of whether the region is a permit buyer or seller.
-Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at $m$ to increase its profits.
+In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$.
+Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits.
+That is, in the absence of a carbon tax, the social planner’s optimal net emissions, $E^{*}$, may not align with the optimal choice of the firm within a region, $E^{m}$, regardless of whether the region is a permit buyer or seller.
+For example, when $\bar{E}$ is very large for every region, the permit price $m$ is zero, so $E^{m}=E^{\mathrm{Total}}$, but it is well known in the literature that $E^{*}$ is often strictly smaller than $E^{\mathrm{Total}}$ (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
+Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at the price $m$ to increase its profits.
 This is also verified in our simulation results later on.
@@
 For simplicity, we assume that emissions cannot be negative in our model.
-From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
-$\tau=\mathrm{MAC}(E^{*})-m,$ (18)
-when $E^{*}>0$.
-Importantly, it implies that for the representative firm to attain the optimal net emission level $E^{*},$ the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$.
-In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC of a region and the market price of emission permits.
-It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*},$ rather than $E^{m}$.
-When $E^{*}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
-$\tau\geq\mathrm{MAC}(E^{*})-m.$ (19)
+From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm (denoted $E^{*,\mathrm{Firm}}$) leads to the following relationship between the carbon tax, the MAC, and the permit price:
+$\tau=\mathrm{MAC}(E^{*,\mathrm{Firm}})-m,$ (18)
+when $E^{*,\mathrm{Firm}}>0$.
+Importantly, it implies that for the representative firm to attain the regional social planner’s optimal net emissions $E^{*}$ (i.e., $E^{*,\mathrm{Firm}}=E^{*}$), the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$.
+In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal net emissions and the market price of emission permits.
+It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*,\mathrm{Firm}}=E^{*}$, rather than $E^{m}$.
+When $E^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
+$\tau\geq\mathrm{MAC}(E^{*,\mathrm{Firm}})-m.$ (19)
 That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits.
-In particular, the regional $\mathrm{MAC}(E^{*})$ at $E^{*}=0$ could be smaller than the global permit price $m$, in which case the optimal carbon tax should be zero.
-Since the optimal carbon tax can always be chosen to be equal to the gap between the MAC and the permit price when the gap is positive, we have the following general formula for any $E^{*}\geq 0$:
-$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\}.$ (20)
+In particular, the regional $\mathrm{MAC}(E^{*,\mathrm{Firm}})$ at $E^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m$.
+However, when $E^{*}=0$, a nonnegative carbon tax larger than $\mathrm{MAC}(E^{*})-m$ will not make $E^{*,\mathrm{Firm}}$ to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}(E^{*})-m$ if it is positive.
+Therefore, for any regional social planner’s optimal net emissions $E^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
+$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\},$ (20)
+such that the firm’s optimal net emissions are consistent with the social planner’s (i.e., $E^{*,\mathrm{Firm}}=E^{*}$).
 The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model.
-Therefore, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
+Thus, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
+4.3 Optimal Carbon Tax under no ETS
+For the case without the ETS, the cost minimization problem of the representative firm facing the emission cap regulation becomes:
+[…]
+From the Karush-Kuhn-Tucker conditions of (21), when the firm’s optimal solution $E^{*,\mathrm{Firm}}=E^{*}$ is not binding at its lower or upper bound, we have the well known equality:
+$\tau=\mathrm{MAC}(E^{*}).$ (22)
+That is, the optimal carbon tax should be chosen to be equal to the MAC at the social planner’s optimal emissions $E^{*}$ if it is not binding.
+This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019).
+We can also see that when $\overline{E}$ is very large or the cap does not exist, a nonzero $E^{*}$ will not be binding, so the equality holds, then the MAC is nonzero.
+In the case with the ETS and very large caps for every region, which is equivalent to the case without the ETS nor the caps, the permit price should be zero, and so our formula (20) is consistent with (22).
+However, the assumption that the MAC is equal to the permit price, which is often used in the literature (e.g., Keohane, 2009; Nordhaus, 2010), does not hold in this case, unless $E^{*}=E^{\mathrm{Total}}$ (i.e., the social planner’s optimal decision is to have zero abatement), which usually does not appear in those non-business-as-usual solutions in the literature (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
 5 Data and Calibration
@@
 Since there is no global cap-and-trade or ETS currently in place, we use the emission commitments of the Paris Agreement and the regions’ net zero commitments to construct the baseline emission cap constraint, $\overline{E}_{i,t}$.
-We collect reports of the Intended Nationally Determined Contributions (INDCs) of 165 countries that were submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 (INDC, 2023).
+We collect reports of the current Nationally Determined Contributions (NDCs) of 165 countries that were first submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 and later updated (INDC, 2023; NDC, 2024).
 Although there is no international enforcement mechanism, most countries committed to achieving a self-defined mitigation target for 2030 (or 2025).
@@
 Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
-As the first step, we obtain the level of emission targets by 2030 for each country.
-In their INDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level in 2015.
+As the first step, we obtain the level of emission targets for each country.
+In their NDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level at some base year.
 Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in net emissions.
@@
 (2019).
-For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$, Ueckerdt et al.
+For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$ for every region $i$, Ueckerdt et al.
 (2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$.
@@
 Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (26) can hold in an approximate manner for every scenario $j$ and time $t$.
+Appendix A.8.6 also reports a sensitivity analysis on the estimate of the abatement cost.
 5.5 Climate Damage
@@
 Until 2100, we can delineate three periods according to the global emission trading patterns.
-Prior to 2032, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left).
+Prior to 2030, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left).
 This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices.
@@
 This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
-Global emissions are constrained by the global emission caps starting from 2032.
+Global emissions are constrained by the global emission caps starting from 2030.
 To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits.
@@
 The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left).
-The global abatement cost also increases steeply to reach its maximum value of \$5.2 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
+The global abatement cost also increases steeply to reach its maximum value of \$5.18 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
 The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left).
-As the emission cap becomes tighter over the years, the permit price increases to \$749 per ton of carbon in 2050 and \$2,225 in 2069 (Figure 2 bottom-left).
-The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions.
-As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050 for the US, EU, Japan, and OHI regions, and 2060 for Russia and China.
+As the emission cap becomes tighter over the years, the permit price increases to \$811 per ton of carbon in 2050 and \$2,254 in 2069 (Figure 2 bottom-left).
+The kink in the permit price path in 2050 is a result of some regions achieving net zero emissions.
+As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050s for Japan (2050) and EU (2053), 2060 for Russia and China, and 2070 for the rest of regions.
 Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
 [^10] In the bottom-left panel of Figure 2, we plot the emission permit price only when the traded volume is positive.
-By the end of this century, the atmospheric temperature is projected to reach 1.74 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right).
+By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right), restricted by the emission caps.
+[^11] If there is no emission cap, then the temperature anomaly will be much higher.
+This can be seen in a later discussion with alternative emission caps.
 Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right).
-Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
+Overall, our noncooperative model simulation predicts that the global emission caps set by the Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
 It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers.
@@
 In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage.
-After 2032, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Middle East, Africa, and OHI regions, with the US and India being the largest permit buyers.
+After 2029, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Africa, and OHI regions, with the US, EU, and India being the largest permit buyers.
 The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider.
@@
 After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound.
-Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China and Eurasia) and the 2060s (the US, the EU, Japan, MidEast, Latin America, and the OHI).
-India, Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
+Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China, Latin America, and the OHI) and the 2060s (the US, the EU, Japan, Eurasia, India, and MidEast).
+Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
 [^13] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large.
@@
 Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4.
-For example, in 2050, the MAC for the US is \$908 per ton of carbon, the permit price is \$749, and their difference is exactly equal to the regional SCC of the US, \$159.
+For example, in 2050, the MAC for the US is \$971 per ton of carbon, the permit price is \$811, and their difference is exactly equal to the regional SCC of the US, \$160.
 Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative.
-Among all the regions, the US has the highest SCC, starting at \$105 per ton of carbon in 2020 and steadily increasing up to \$210 per ton of carbon by the end of this century.
-The US is followed by the EU (from \$42 to \$138 per ton of carbon) and China (from \$39 to \$115 per ton of carbon) in this century.
+Among all the regions, the US has the highest SCC, starting at \$106 per ton of carbon in 2020 and steadily increasing up to \$213 per ton of carbon by the end of this century.
+The US is followed by the EU (from \$42 to \$139 per ton of carbon) and China (from \$39 to \$117 per ton of carbon) in this century.
 The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon).
@@
 We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation.
-Specifically, cooperation results in approximately 0.19 degrees Celsius lower temperature with the ETS and 0.15 degrees Celsius lower temperature without the ETS.
+Specifically, cooperation results in approximately 0.17 degrees Celsius lower temperature with the ETS and 0.13 degrees Celsius lower temperature without the ETS.
 Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps.
-When comparing the cases with and without the ETS, we find that the ETS may play different roles under noncooperation and cooperation.
+Moreover, the ETS has little impact on the temperature path, by reducing the temperature anomaly in 2010 with only 0.02 degrees Celsius, as the emission caps exist no matter whether the ETS exists or not.
+When comparing the cases with and without the ETS, [^14] Note that the emission caps are still there, no matter whether the ETS is implemented or not.
+we find that the ETS may play different roles under noncooperation and cooperation.
 In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS.
@@
 In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time.
-With the ETS, slightly higher levels of global emissions and temperature rise are observed before 2040 and 2050, respectively, and the opposite trends are observed after that.
+With the ETS, lower levels of global emissions are observed after 2030.
 This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
@@
 and cannot be further mitigated, so its MAC starts to decline over time.
-The effect of the ETS depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU at around 2040); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm).
+The effect of the ETS under cooperation depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm).
+Moreover, the ETS has little impact on the net emissions and the MAC of a developed region as shown in Figures A.7 and A.9 (in Appendices A.8.2 and A.8.3) respectively, because a developed region will abate significantly no matter whether the ETS exists or not under cooperation.
+That is, the ETS plays an important role only in developing regions under cooperation.
 Figure 7: Simulation results of different cases for the US in the cooperative or noncooperative model.
@@
 This is consistent with the high MAC of the US.
-When the US regional net emissions are not binding at zero before 2032 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2032.
+When the US regional net emissions are not binding at zero before 2030 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2030.
 Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions.
@@
 Consequently, the US conducts less emission abatement.
-For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2032 (see Figure A.10 in Appendix A.8.3).
+For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2030 (see Figure A.10 in Appendix A.8.3).
 The bottom-right panel of Figure 7 shows that the noncooperative SCC of the US is much smaller than its cooperative SCC, which is also consistent with the smaller MAC of the US under noncooperation.
@@
 Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS.
-For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2032 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2032 (see the two top panels of Figure 7), which is equal to the optimal regional carbon tax under cooperation.
-For the noncooperative case without the ETS, the US regional net emissions are binding at their caps (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC (see the two bottom panels of Figure 7), which is equal to the optimal regional carbon tax under noncooperation.
-Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Russian under cooperation until 2057 in Figures A.7, A.9, A.11); and Russian under noncooperation until 2042 in Figures A.8, A.10, and A.12).
+For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2033 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2033 (see the two top panels of Figure 7).
+For the noncooperative case without the ETS, the US regional net emissions are binding at their caps after 2022 (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC after 2022 (see the two bottom panels of Figure 7).
+Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Eurasia under cooperation until 2060 in Figures A.7, A.9, A.11; and Eurasia under noncooperation until 2049 in Figures A.8, A.10, and A.12), which is equal to the optimal regional carbon tax.
 This relationship holds for every region as shown in Figures A.7-A.12.
-The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under both cooperation and noncooperation.
+The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under cooperation or noncooperation.
 This pattern holds for every region as shown in Figures A.11 and A.12.
+This happens under noncooperation because the regional noncooperative SCC is the present value of of future climate damages in the region resulting from an additional unit of global emissions released in the current period, and the ETS has little impact on the marginal damages and our endogenous social discount rates.
+Under cooperation, since the ETS has little impact to developed regions and the temperature path, the wealth redistribution among developing regions via the ETS has little impact on the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$,, so the ETS has little impact on the regional cooperative SCC.
+This shows that the regional SCC is complementary to the ETS, as the emission caps exist no matter whether the ETS exists or not.
 6.3 Alternative Net Zero Scenarios
@@
 The top-left and top-right panels in Figure 8 display the emission permit prices and expected temperature increases under different emission cap scenarios.
-Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,538 per ton of carbon in 2049, and the temperature rise is restricted to 1.64 degrees Celsius by the end of this century.
-Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$344 and \$158 per ton of carbon in 2049, respectively.
-In these scenarios, the temperature rise by the end of the century is expected to reach 1.86 degrees Celsius and 2.08 degrees Celsius above the pre-industrial level, respectively.
+Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,593 per ton of carbon in 2049, and the temperature rise is restricted to 1.61 degrees Celsius by the end of this century.
+Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$445 and \$261 per ton of carbon in 2049, respectively.
+In these scenarios, the temperature rise by the end of the century is expected to reach 1.8 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively.
 This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
@@
 Our results show that stricter emission caps lead to higher MACs.
-Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2086 under the net zero 2090 scenario.
+Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2085 under the net zero 2090 scenario.
 This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC.
-We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$216 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$193 in 2100 in the net zero 2090 scenario.
+We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$218 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$198 in 2100 in the net zero 2090 scenario.
 This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive).
@@
 Economic parameters
-$\beta=0.985$
-Discount factor
-$\gamma=1.45$
-Elasticity of marginal utility
-$\alpha=0.3$
-Output elasticity of capital
-$\delta=0.1$
-Annual depreciation rate of capital
+$\beta=0.985$ Discount factor
+$\gamma=1.45$ Elasticity of marginal utility
+$\alpha=0.3$ Output elasticity of capital
+$\delta=0.1$ Annual depreciation rate of capital
 Climate parameters
-$\zeta=0.0021$
-Contribution rate of carbon emissions to temperature
-$\sigma_{i,t}$
-Carbon intensity (GtC per trillion \$)
-$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$
-Abatement cost parameters
-$\pi_{1,i}$, $\pi_{2,i}$
-Climate damage parameters
+$\zeta=0.0021$ Contribution rate of carbon emissions to temperature
+$\sigma_{i,t}$ Carbon intensity (GtC per trillion \$)
+$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$ Abatement cost parameters
+$\pi_{1,i}$, $\pi_{2,i}$ Climate damage parameters
 TFP parameters
-$g_{i,0},g_{i,t}$
-Growth rate of TFP
-$d_{i}$
-Parameters of TFP growth
+$g_{i,0},g_{i,t}$ Growth rate of TFP
+$d_{i}$ Parameters of TFP growth
 A.2 List of Countries
@@
 Region Constituent Countries
-Africa
-Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
+Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
 European Union [^17] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience.
 Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
-Eurasia
-Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
-Latin America
-Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St.
+Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
+Latin America Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St.
 Kitts and Nevis, St.
@@
 Region Constituent Countries
-Middle East
-Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
-Other Non-OECD Asia
-Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
-Other High-Income
-Australia, Canada, Republic of Korea, New Zealand, Singapore.
+Middle East Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
+Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
+Other High-Income Australia, Canada, Republic of Korea, New Zealand, Singapore.
 A.3 Calibration of the TCRE Climate System
@@
 Figure A.14: Simulation results of regional SCC: comparison across alternative emission cap scenarios.
+A.8.6 Sensitivity Analysis on Marginal Abatement Cost
+In Figure A.15, we provide additional sensitivity analysis results on the abatement cost. Recall that, the MAC in the model is
+$\mathrm{MAC}{}_{i,t}=1,000b_{2,i}\mu_{i,t}^{b_{2,i}-1}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t)).$
+We investigate a scenario wherein the MACs across all regions experience a variation of ±20% from the baseline MAC levels at the same emission control rates $\mu_{i,t}$.
+Specifically, we explore the effects of adjusting two key parameters $b_{1,i}$ and $b_{3,i}$, by ±20% from their baseline values.
+The result shows that the equilibrium permit price is closely linked to the level of the MACs.
+When the MACs are elevated across all regions, the permit price also experiences a corresponding increase with almost identical changes of the MACs, and vise versa.
+For example, in 2040, the baseline MAC in the US is \$556, it is changed to \$644 or \$467, and the permit price is changed from \$404 to \$492 or \$315.
+That is, both the MACs and the permit price increase \$88 or decreases \$89 in 2040.
+At the same time, the regional SCC has little changes, as the existence of the emission caps makes the future temperature have little changes and the changes in the MACs have little impact on marginal climate damages in future periods and social discount rates.
+This again verifies our relation between the optimal carbon tax, the MAC, and the permit price (i.e., equation (20)).
+Figure A.15: Sensitivity Analysis on the MAC.

### v3

@@
-Yongyang Cai  Khyati Malik   Hyeseon Shin
+
 Abstract
 We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world.
-We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and updated in subsequent years.
-Our model shows that the emission permit price reaches \$811 per ton of carbon by 2050.
-We demonstrate that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
+We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and later strengthened under the Glasgow Pact.
+Our model shows that the emission permit price reaches \$845 per ton of carbon by 2050, and global average temperature is expected to reach 1.7°C above the pre-industrial level by the end of this century.
+We demonstrate, both theoretically and numerically, that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
+We show that optimal regional carbon tax has significant heterogeneity between regions, and the tax needs to be implemented in both the developed and the developing regions.
 Keywords: Emission trading system, Paris Agreement, dynamic Nash equilibrium, integrated assessment model, carbon tax, social cost of carbon.
@@
 1 Introduction
-Climate change has been raising concerns of disastrous consequences, from rising sea levels to increases in the frequency and intensity of extreme weather events (Arias et al., 2021).
+Climate change has been raising concerns of disastrous consequences, from rising sea levels to increase in the frequency and intensity of extreme weather events (Arias et al.,, 2021).
 International efforts to address climate change have evolved over the past three decades, commencing with the United Nations Framework Convention on Climate Change (UNFCCC) in 1992, followed by the Kyoto Protocol in 1997, the Paris Agreement in 2015, and the Glasgow Pact in 2021.
-The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2 degrees Celsius above pre-industrial levels, with a pursuit of limiting it to the 1.5 degrees Celsius” (Article 2).
-To accomplish this objective, 194 countries (parties) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures.
-Most recently, NDCs have been revisited in the Glasgow Climate Pact in November 2021, reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more progressive efforts to attain the global target of 1.5 degrees Celsius.
+The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2°C above pre-industrial levels, with a pursuit of limiting it to the 1.5°C” (Article 2).
+To accomplish this objective, 194 countries (regions) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures.
+Most recently, the Glasgow Climate Pact, adopted at the UN Climate Change Conference in Glasgow (COP26) in November 2021, revisited Nationally Determined Contributions (NDCs), reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more stringent efforts to attain the global target of 1.5°C.
 Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETSs).
-[^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for future implementation around the world (Stavins, 2022).
-In a carbon tax regime, a tax is charged on firms for each unit of carbon emissions.
-Previous studies have shown that the optimal global carbon tax rate is determined at the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus, 2017).
-In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price.
-In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery, 1972; Goulder and Schein, 2013; Stavins, 2022).
+[^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for implementation globally (Stavins,, 2022).
+In a carbon tax regime, a tax is charged to firms for each unit of carbon emission.
+Previous studies have shown that the optimal global carbon tax rate is equal to the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus,, 2017).
+In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of permissible emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price.
+In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery,, 1972; Goulder and Schein,, 2013; Stavins,, 2022).
 In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, [^2] In addition to the EU ETS, other national or sub-national ETSs have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States.
-For the information on most up-to-date policy in practice, see World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/).
-and as of the summer of 2023, the permit price stands at about 90 euros per ton of Carbon Dioxide (CO2).
+For information on the most up-to-date policy in practice, see World Bank Carbon Pricing Dashboard (https://carbonpricingdashboard.worldbank.org/).
+and as of the winter of 2024, the permit price stands at about 70 euros per ton of carbon dioxide (CO2).
 [^3] Source of EU permit price: (https://tradingeconomics.com/commodity/carbon).
 Although a global-level ETS is currently not in place, there have been longstanding discussions on its implementation as an international regime to address climate change, as implied by Article 17 of the Kyoto Protocol and Article 6 of the Paris Agreement.
-In this study, we develop a dynamic multi-region model of climate and the economy under a hypothetical global ETS, and use it to project global prices of emission permits and the regional Social Cost of Carbon (SCC).
-In this integrated assessment framework, we formed 12 aggregated regions by grouping 198 countries around the world.
-Each region is allocated emission caps in accordance with the emission mitigation targets from the Paris Agreement and regional net zero commitments.
-Our main analysis involves solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions.
-Under the global ETS, regions with higher marginal abatement costs (MAC) are encouraged to purchase permits from the market to meet their emission cap constraints, whereas regions with lower MAC are incentivized to conduct more abatement and make profits by selling permits.
+Prior works have emphasized the need for a global ETS or linking existing regional ETSs (Mehling et al.,, 2018) and for expanding the coverage of carbon trading in specific industries to different regions so as to enhance the global mitigation efforts and reduce carbon leakages (Piris-Cabezas et al.,, 2018).
+In this study, we develop a dynamic multi-region model of climate and the economy under a global ETS with annual time steps, and use it to project global prices of emission permits and the regional SCC.
+In this integrated assessment framework, we form 12 aggregated regions by grouping 152 countries around the world.
+For these regions, we construct emission cap trajectories aligned with the latest NDC mitigation targets from the Paris Agreement and the Glasgow Climate Pact.
+These trajectories are based on near-term targets for 2025/2030 and long-term net-zero commitments for 2050–2070, and our model treats these trajectories as exogenous emission cap constraints.
+Assuming that all regions adhere strictly to their emission cap constraints without future revisions, [^4] In this paper, we assume that regions will not revise their NDC limits in future for the following reasons: (i) relaxing NDC limits could result in penalties from other regions or damage to the region’s reputation; (ii) reducing NDC limits would lead to a loss of benefits from selling emission permits or more cost from purchasing; and (iii) renegotiating NDC limits is often time-consuming and infrequent.
+our main analysis focuses on solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions.
 The emission permit price is endogenously determined by annual supply and demand in the global permit market.
@@
 The first contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS.
-While carbon tax and ETS have often been considered as potentially interchangeable policies, we show that within the global ETS framework, it is optimal for the social planner of each region to introduce a carbon tax as a complementary policy tool to achieve its optimal level of net emissions.
-That is, a hybrid policy with both ETS and carbon tax is optimal.
-The necessity of a carbon tax arises from the discrepancy in optimal net emissions between the social planner’s choice and the market outcome under the ETS only (i.e., firms’ choice under no carbon tax).
-When optimal net emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional MAC and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with the social planner’s.
-We also numerically verify that the regional optimal carbon tax is equal to the regional SCC when net emissions are strictly positive.
+While a carbon tax and ETS have often been considered as potentially interchangeable or substitutable policies, we show that within the global ETS framework, a hybrid policy combining both ETS and carbon tax is optimal: the social planner of each region needs to introduce a carbon tax as a complementary policy tool under the global ETS in order to achieve its optimal level of emissions.
+The necessity of a carbon tax arises from the discrepancy in optimal emissions between the social planner’s choice and the market outcome under an ETS-only regime (i.e., firms’ choice under the ETS with an exogenous emission cap but no carbon tax).
+Specifically, when optimal emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional marginal abatement cost (MAC) and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with that of the social planner’s.
+We also numerically verify that the regional optimal carbon tax is equal to the regional SCC provided emissions are strictly positive.
+Regional SCC is the present value of future regional climate damages resulting from additional global emissions.
 The second contribution of this study is a method to solve for the dynamic Nash equilibrium prices of emission permits under a global ETS.
 To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with fully endogenous prices and emission abatements in a dynamic regional integrated assessment model (IAM) under noncooperation.
-Although some prior research has addressed regional integrated assessment models with a multi-region ETS, their solutions thus far have been obtained in a limited manner.
-One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang, 1996; Nordhaus, 2010), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation.
-Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al., 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni, 2012; Massetti and Tavoni, 2012).
-While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane, 2009; Nordhaus, 2010), our study shows that this does not hold.
-Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between forward-looking regional social planners’ and myopic firms’ choices under the ETS and no carbon tax.
-This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for every region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation with the assumption.
-However, it is well known that even without emission caps, there are still nonzero emission mitigations and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
-We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and, importantly, show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions.
-Under the baseline emission cap scenario based on the Glasgow Pact and the regional net zero commitments, our simulation results show that the oversupply of global permits in the initial decade results in zero permit prices, but by 2050, the emission permit price can reach up to \$811 per ton of carbon.
-Our results further show that regions such as the United States, India and the EU become major emission permit buyers, and other regions such as Russia and China become permit suppliers.
-The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Glasgow Pact, rendering our findings relevant for the real world scenario.
-We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz, 2017), also known as the “Middle-of-the-Road” scenario.
-Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al., 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on cumulative carbon emissions.
-We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al., 2011).
-Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, 2017; Dietz and Venmans, 2019; Mattauch et al., 2020; Barnett et al., 2020.
-Moreover, Dietz et al.
-(2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus, 2017), compared to other more complicated climate systems.
-Our final contribution is a comprehensive comparative analysis of the roles played by cooperation, emission caps, and the global ETS.
-Our simulations reveal that global cooperation leads to significantly lower emissions and lower temperature rises.
-We also find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the first decade.
+Although some prior research has addressed regional IAMs with a multi-region ETS, their solutions thus far have been obtained in a limited manner.
+One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation.
+Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al.,, 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni,, 2012; Massetti and Tavoni,, 2012).
+While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane,, 2009; Nordhaus, 2010b, ), our study shows that this does not hold.
+Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between regional social planners’ choices internalizing the regional climate damages and firms’ choices ignoring the regional climate damages under the ETS regime.
+This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for any region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation.
+However, it is well known that even without emission caps, there are still nonzero emission mitigations from the regional social planners’ perspectives and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang,, 1996; Cai et al.,, 2023).
+That is, regions’ social planners should impose an additional carbon tax on firms, contingent on their NDC commitments, such that the firms’ emission decisions are consistent with theirs.
+We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and importantly, we show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions under the Nash equilibrium.
+Under the baseline emission cap scenario based on the Paris Agreement and Glasgow Pact commitments, our simulation results show that the oversupply of global permits in the initial years results in zero permit prices, but by 2050, the emission permit price can reach up to \$845 per ton of carbon.
+The global average temperature is expected to reach 1.7 degree Celcius above the pre-industrial level by the end of this century.
+The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Paris Agreement and the Glasgow Pact, making our findings directly relevant to real-world scenarios.
+We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz,, 2017), also known as the “Middle-of-the-Road” scenario.
+Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al.,, 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on the cumulative carbon emissions.
+We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al.,, 2011).
+Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas,, 2017; Dietz and Venmans,, 2019; Mattauch et al.,, 2020; Barnett et al.,, 2020.
+Moreover, Dietz et al., (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus,, 2017), compared to other more complicated climate systems.
+Furthermore, our construction of emission cap pathways reflects the latest emission targets for 2030 and net-zero pledges for 2050-2070, contributing to the scant literature on evaluating the economic and environmental implications of these commitments (e.g., van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022).
+Our final contribution is a comprehensive comparative analysis of the roles of emission caps and the global ETS.
+We find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the initial years.
 This implies that it is important to maintain a stringent emission cap at the global level to make the ETS work efficiently.
-Moreover, we show that a stricter emission cap leads to a higher permit price and smaller regional SCCs for every region.
-For example, if all the regions achieve net zero emissions in 2050, the emission permit price can reach up to \$1,593 per ton of carbon by 2049.
-However, even with this strict emission restriction, the global temperature is still projected to rise by 1.61 degrees Celsius by the end of this century.
-These findings suggest that achieving the global target of limiting the temperature rise to 1.5 degrees Celsius requires even stronger emission commitments than those outlined in the Glasgow Pact.
-Our results are in qualitative agreement with the analysis of Rogelj et al.
-(2016), which found that the first NDC targets of the countries would result in a median warming of 2.1 to 3.1 degrees Celsius by 2100, and therefore stronger measures are needed to keep the warming under 1.5 degrees Celsius above the pre-industrial levels.
-Riahi et al.
-(2015) and Spencer et al.
-(2015) have similarly concluded that 2030 mitigation targets are insufficient to reach the greenhouse gas stabilization levels of 450 ppm.
-We also find that under cooperation between the regions, the regional SCC is higher for a developed region compared to a developing one.
-That is, there are higher carbon taxes in developed regions under cooperation compared to developing regions because the marginal utility of consumption is higher in developing regions.
-The global aggregate welfare will be higher if developed regions can contribute more to mitigating emissions, thus reducing climate damages in developing regions.
-In the noncooperative setting, the regional SCCs of the developed and the developing regions are comparable.
-Our model framework is closely related to the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010), an extension of the global DICE model (Nordhaus, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems.
-In Nordhaus (2010), carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities.
+Moreover, we show that a stricter emission cap leads to a higher permit price for every region.
+For example, if all the regions achieve net zero emissions by 2050, the emission permit price can reach up to \$1,539 per ton of carbon by 2049.
+However, even with this strict emission restriction, the global temperature is still projected to rise by 1.62°C by the end of this century.
+These findings suggest that achieving the global target of limiting the temperature rise to 1.5°C requires even stronger emission commitments than those outlined in the Glasgow Pact.
+Our results are in qualitative agreement with the analysis of other IAMs (van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022), which find that the strengthened post-Glasgow NDC targets, encompassing both near-term and long-term goals, are insufficient to limit warming to 1.5°C above pre-industrial levels, but, if fully implemented, could limit warming to below 2°C.
+To ensure that our results are robust, we have performed sensitivity analysis by choosing alternative climate damage and abatement cost parameters obtained by fitting these parameters to other literature data.
+Our model framework is closely related to the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), an extension of the global DICE model (Nordhaus,, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems.
+In Nordhaus, 2010b , carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities.
 Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region framework under cooperation and noncooperation.
-For instance, Luderer et al.
-(2012) and Jakob et al.
-(2012) compare the long-term predictions of three regionalized energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm.
+For instance, Luderer et al., (2012) and Jakob et al., (2012) compare the long-term predictions of three-region energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm.
 However, these works do not incorporate actual NDC and net-zero year targets of the countries to study the region-wise emission-economy pathways.
-van der Ploeg and de Zeeuw (2016) and Jaakkola and Van der Ploeg (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation.
-Hambel et al.
-(2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions.
-Iverson and Karp (2021) study a Markov Perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates.
-Cai et al.
-(2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation.
+van der Ploeg and de Zeeuw, (2016) and Jaakkola and Van der Ploeg, (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation.
+Hambel et al., (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions.
+Iverson and Karp, (2021) study a Markov perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates.
+Cai et al., (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation.
 Nonetheless, none of these works examine the dynamics of an international ETS with an endogenous market price of emission permits under a noncooperative setting.
-Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw (2016) and Cai et al.
-(2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
+Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
 In the literature, ETSs have garnered attention as promising mechanisms for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze ETSs as potential instruments for global climate policy.
-Among the most closely related studies are extended versions of the WITCH model (Bosetti et al., 2006): Massetti and Tavoni (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost.
+Among the most closely related studies are extended versions of the WITCH model (Bosetti et al.,, 2006): Massetti and Tavoni, (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni, (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost.
 As noted earlier, our solution mechanism differs from these studies and provides permit prices from a complete noncooperative dynamic optimization problem.
-Carbone et al.
-(2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price.
+Carbone et al., (2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price.
 However, that study examines the ETS in a static setting, missing the critical dynamic aspects of climate change and lacking a connection with emission abatement decisions.
-Fischer and Springborn (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions.
-Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler, 2018; Doda et al., 2019; Holtsmark and Weitzman, 2020; Holtsmark and Midttømme, 2021).
-More broadly, a substantial body of literature have examined ETSs from various perspectives.
-While in this study, we assume that emission permits are freely allocated following the voluntary commitments from the Paris Agreement, few studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry, 2008; Goulder et al., 2010).
-Some have studied the relative efficiency of carbon policies under the existence of cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions and vice versa (Weitzman, 1974; Stavins, 1996; Karp and Traeger, 2024).
-Some studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg, 2018; Fuss et al., 2018; Perino et al., 2022), China ETS (Goulder et al., 2022), or California ETS (Borenstein et al., 2019).
+Fischer and Springborn, (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions.
+Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler,, 2018; Doda et al.,, 2019; Holtsmark and Weitzman,, 2020; Holtsmark and Midttømme,, 2021; Mehling et al.,, 2018).
+More broadly, a substantial body of literature has examined ETSs from various perspectives.
+While in this study, we assume that emission permits are freely allocated to the regions following the voluntary commitments from the NDCs, several studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins,, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry,, 2008; Goulder et al.,, 2010).
+Other works have studied the relative efficiency of carbon policies under cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions, and vice versa (Weitzman,, 1974; Stavins,, 1996; Karp and Traeger,, 2024).
+In another strand of literature, a number of studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg,, 2018; Fuss et al.,, 2018; Perino et al.,, 2022), China ETS (Goulder et al.,, 2022), or California ETS (Borenstein et al.,, 2019).
 An important issue with regional ETSs is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS.
-Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al., 2024; Böhringer et al., 2014; Farrokhi and Lashkaripour, 2021; Fowlie and Reguant, 2022; Levinson, 2023).
-While there is no general agreement among economists on whether carbon tax or ETS is better (Stavins, 2022), the literature have often considered two pricing instruments (trading versus taxes) as policy substitutes.
-For instance, some studies support carbon tax over ETS, as ETS is subject to price volatility (Nordhaus, 2007).
-Conversely, other studies advocate ETS over tax because allocation of permits in ETS allows more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane, 2009).
-ETS could be a preferred instrument over tax in practice, as free allowances could be negotiated among participating agents to redistribute burden and, thus, be used to gain political support (Harstad and Eskeland, 2010; Hahn and Stavins, 2011; Stavins, 2022).
-Newell and Pizer (2003) show that ETS is preferable over carbon tax in the long run, and Knopf et al.
-(2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn, 2008).
-Harstad and Eskeland (2010) argue that an ETS in a perfect market is the first-best system, but frequent government interventions to redistribute allocations among firms results in distortions in the market allocation.
-In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al., 2016).
-For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand (2013); Schmalensee and Stavins (2017); Cai (2021); Stavins (2022).
+Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al.,, 2024; Böhringer et al.,, 2014; Farrokhi and Lashkaripour,, 2021; Fowlie and Reguant,, 2022; Levinson,, 2023).
+While there is no general agreement among economists on whether a carbon tax or an ETS is better (Stavins,, 2022), the literature have often considered these two pricing instruments (trading versus taxes) as policy substitutes.
+For instance, some studies support carbon tax over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003).
+For instance, Newell and Pizer, (2003) find that carbon taxes can yield higher welfare benefits than an ETS under such uncertainties.
+Conversely, other studies advocate ETS over carbon tax because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009).
+Harstad and Eskeland, (2010), Hahn and Stavins, (2011), and Stavins, (2022) argue that from a practical perspective, ETS may be a preferred instrument over tax, as free allowances could be negotiated among participating agents to redistribute burden and, thus be used to gain political support.
+Knopf et al., (2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn,, 2008).
+In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016).
+However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms results in distortions in the market allocation.
+For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022).
 This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary.
-A global ETS necessitates regional carbon taxes.
+2 Optimal Carbon Tax under the ETS
+While carbon taxes and ETS are often considered as substitutable policies, we demonstrate that each country needs to implement an optimal carbon tax as a complementary policy within a global ETS regime to achieve optimal level of emissions.
+Before introducing the full multi-region dynamic general equilibrium model, it is helpful to first explore the optimal carbon tax within a static framework.
+Under the global ETS framework, we assume that each region $i$ is allocated with an exogenous level of emission cap ($\bar{E}_{i}$), which is the upper limit of emissions allowed for that region and is strictly enforced.
+The region’s gross emissions before abatement is $E_{i}^{\mathrm{Gross}}$, its emissions net of abatement is $E_{i}$, the amount of emission permits purchased (or sold) in the market is $E_{i}^{P}$ (if $E_{i}^{P}<0$, it implies that the region sells emission permits) so we have the emission cap constraint $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$, where the Nash equilibrium permit price is $m$ under the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$.
+Here we assume that each region is a price taker.
+The region’s marginal abatement cost (MAC) function, denoted $\mathrm{MAC}_{i}(E_{i})$, is assumed to be monotonically decreasing over emissions $E_{i}$ with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$.
+A social planner of the region solves the following minimization problem
+[…]
+where the first term of the objective function is the regional damage from the global emissions $\sum_{j\in\mathcal{I}}E_{j}$ (which is the present value of future damage from the global emissions), the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E_{i}^{\mathrm{Gross}}-E_{i})$.
+For simplicity, we assume that emissions cannot be negative in our model.
+In the static Nash equilibrium of all regional social planners, every region solves equation (1) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$.
+Let $E_{j}^{*}$ be the Nash equilibrium emissions for each region $j$ and let $m^{*}$ be the Nash equilibrium price.
+Thus, from the Karush-Kuhn-Tucker conditions of (1), the solution to the region’s optimal emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
+$D_{i}^{\prime}\left(\sum_{j\in\mathcal{I}}E_{j}^{*}\right)=\mathrm{MAC}_{i}\left(E_{i}^{*}\right)-m^{*},$ (2)
+when $E_{i}^{*}>0$.
+The marginal regional damage can also be regarded as the regional social cost of carbon (SCC).
+Note that in the static Nash equilibrium with $E_{i}^{*}>0$, the permit price $m^{*}$ should be equal to the shadow price of the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$.
+That is, if the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ is not binding, or $\sum_{j\in\mathcal{I}}E_{j}^{*}<\sum_{j\in\mathcal{I}}\overline{E}_{j}$ (i.e., there is an over-supply of emission permits), then the permit price is zero.
+Alternatively, if the permit price is nonzero, then the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ must be binding, so the supply of emission permits is equal to the demand (i.e., $\sum_{j\in\mathcal{I}}E_{j}^{*}=\sum_{j\in\mathcal{I}}\overline{E}_{j}$).
+We emphasize that the assumption commonly used in the literature—that the MAC equals the permit price (e.g., Keohane,, 2009; Nordhaus, 2010b, )—however, does not hold in this case, unless the regional marginal damage or SCC is zero, according to equation (2).
+Moreover, we also numerically verify in our dynamic Nash equilibrium that, when $E_{i}^{*}>0$, the regional MAC is also strictly larger than the permit price, and the difference is the regional SCC.
+An important issue arises when a regional social planner cannot directly enforce its optimal emissions level on firms, the actual emission producers in the economy.
+To address this, let us assume that the social planner can employ carbon taxation as a policy instrument to resolve the discrepancy in optimal emissions between the firms and the regional social planner.
+Consider a representative firm in region $i$ that produces emissions and operates under the region’s MAC function and emissions cap constraint in a static framework.
+Since the firm does not internalize the damage from emissions, we assume that the social planner of region $i$ levies a tax on the firm of $\tau_{i}$ per unit of emissions.
+The firm solves the following minimization problem:
+[…]
+where the first term of the objective function is the total tax on emissions.
+In the static Nash equilibrium of all representative firms, every firm solves equation (3) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$.
+Let $E_{j}^{*,\mathrm{Firm}}$ be the Nash equilibrium emissions for each firm $j$ and let $m^{*,\mathrm{Firm}}$ be the Nash equilibrium price.
+Thus, from the Karush-Kuhn-Tucker conditions of (3), the solution to the optimal emissions of the firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
+$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}},$ (4)
+when $E_{i}^{*,\mathrm{Firm}}>0$.
+If the regional social planner imposes the tax to be equal to the regional SCC, i.e., the regional marginal damage, then from equations (2) and (4), we have $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ and $m^{*,\mathrm{Firm}}=m$.
+In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal emissions and the market price of emission permits, capturing the idea that the regional social planner can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its emissions level of $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$.
+When $E_{i}^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (3) shows that
+$\tau_{i}\geq\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}}.$ (5)
+That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits.
+In particular, the regional $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})$ at $E_{i}^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m^{*,\mathrm{Firm}}$, making the right hand side of the inequality (5) to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ if it is positive.
+Therefore, for any regional social planner’s optimal emissions $E_{i}^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
+$\tau_{i}=\max\{0,\ \mathrm{MAC}_{i}(E_{i}^{*})-m^{*}\},$ (6)
+such that the firm’s optimal emissions are consistent with the social planner’s (i.e., $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$).
+The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model.
+Note that this tax is contingent on the imposition of the emission cap on the firm, and the tax level itself cannot make the firm’s emissions to be $E_{i}^{*}$ unless $m^{*}=0$.
+It can be easily shown that our argument is consistent with the existing literature when there is no ETS.
+The cost minimization problem of the representative firm facing the emission cap regulation without the ETS becomes:
+[…]
+From the Karush-Kuhn-Tucker conditions of (7), when the firm’s optimal solution $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ is not binding at its lower or upper bound, we have the well known equality:
+$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*}).$ (8)
+This implies that the optimal carbon tax should be set equal to the MAC at the regional social planner’s optimal emissions $E_{i}^{*}$, provided $E_{i}^{*}$ is not binding.
+This result has been widely used in the literature (e.g., Nordhaus,, 2014; Cai et al.,, 2017; Cai and Lontzek,, 2019).
+Note that our formula (6) under an ETS with zero permit price is consistent with (8) under no ETS.
+A regional social planner in each region $i$ has two targets: one is its exogenous NDC commitment $\overline{E}_{i}$, and the other is the regional social planner’s endogenous optimal emissions $E_{i}^{*}$ after permit trading (note that $\overline{E}_{i}$ and $E_{i}^{*}$ are unequal in general [^5] We believe this is a realistic consideration, as emission cap pathways may not always align precisely with regional optimal emissions over time.
+This discrepancy arises because the negotiation of emission permit allocations is often time-consuming and influenced by political considerations, which may not lead to economically optimal outcomes.
+Our model simulation shows that the emission cap pathways constructed from the Paris Agreement and the Glasgow Climate Pact indeed deviate from the optimal regional emissions.) in the Nash equilibrium of all 12 regional social planners.
+That is, the region $i$’s emissions should be $E_{i}^{*}$, which should also satisfy $E_{i}^{*}\leq\overline{E}_{i}+E_{i}^{P}$ after permit trading in the Nash equilibrium.
+Several policy instruments can be applied to enforce the firms to meet the targets, assuming that the firms can buy or sell emision permits.
+One such instrument is a quantity policy, in which the regional social planner distributes $\overline{E}_{i}$ (e.g., by means of an auction or grandfathering) to its firms, ensuring the firms’ aggregate emissions, $E_{i}^{*,\mathrm{Firm}}$, do not exceed the exogenous allocated cap after permit trading (i.e., $E_{i}^{*,\mathrm{Firm}}\leq\overline{E}_{i}+E_{i}^{P}$) in the Nash equilibrium of firms of all regions.
+[^6] The representative firm framework can be generalized to an environment with multiple firms.
+Assume that there are $n$ homogenous firms in the region $i$, each having $\overline{E}_{i}/n$ as their caps, with a firm’s MAC function at its emissions $e$ is defined as $\mathrm{MAC}_{i}(ne)$, where $\mathrm{MAC}_{i}(\cdot)$ is the aggregate regional MAC function used by the regional social planner and $E=ne$ is the aggregate regional emissions.
+Under this definition, the aggregate total abatement cost of all firms in region $i$ is equal to the integral of $\mathrm{MAC}_{i}(E)$ on $[E_{i},E_{i}^{\mathrm{Gross}}]$, when each firm’s emission is $E_{i}/n$ while $E_{i}^{\mathrm{Gross}}/n$ is each firm’s upper bound under no abatement.
+In this case, the Nash equilibrium of the economy with multiple firms is identical to the equilibrium with a representative firm.
+However, under a quantity-only control policy, $E_{i}^{*,\mathrm{Firm}}$ is not equal to $E_{i}^{*}$, as the zero-tax Nash equilibrium of the firms, $\left\{m^{*,\mathrm{Firm}},E_{i}^{*,\mathrm{Firm}}:j\in\mathcal{I}\right\}$, is not the Nash equilibirum of the regional social planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$.
+This becomes particularly evident in the following extreme case: assume that $\overline{E}_{i}$ is larger than $E_{i}^{\mathrm{Gross}}$ for every region, which leads to $m^{*}=0$ under the ETS.
+The firms’ Nash equilibrium without a tax yields $E_{i}^{*,\mathrm{Firm}}=E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})=m^{*,\mathrm{Firm}}=0$.
+In contrast, the regional social planners’ Nash equilibrium leads to $E_{i}^{*}<E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*})>m^{*}=0$ in general, similar to a model without emission caps.
+Note that this quantity policy cannot directly distribute $E_{i}^{*}$, as doing so would either result in the region violating its commitment if $E_{i}^{*}>\overline{E}_{i}$ or losing the benefit in selling permits if $E_{i}^{*}<\overline{E}_{i}$.
+Another instrument is a carbon tax policy, in which the region $i$ imposes a carbon tax $\tau_{i}$ on the firms, e.g., $\tau_{i}=\mathrm{MAC}_{i}(\overline{E}_{i})$ or $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})$.
+However, a carbon tax policy alone can target only one emission level, either $\overline{E}_{i}$ or $E_{i}^{*}$, in each region, while $\overline{E}_{i}\neq E_{i}^{*}$ in general.
+Therefore, neither instrument can achieve both targets simultaneously.
+We argue that in the presence of a global ETS, it is optimal for each region $i$ to impose both the quantity policy and an additional carbon tax policy simultaneously on firms to enforce both targets.
+Here the additional carbon tax for region $i$ is chosen to be $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ when $E_{i}^{*}>0$, under the Nash equilibirum of the planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$.
+If some regions impose only the quantity policy while the others impose both policies, this leads to a Nash equilibrium of firms, which is not the Nash equilibirum of the planners.
+Note that a tax has no impact on the regional social planners’ decisions; $E_{i}^{*}$ remains the optimal decisions of the regional social planners regardless of the tax levels.
+However, the tax can influence the firms’ decisions.
+Therefore, from the regional social planners’ perspective, any Nash equilibrium of firms deviating from $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$ is worse, and the regions should always impose both the quantity policy and the optimal additional carbon taxes simultaneously on the firms to meet both targets.
+[^7] Our argument is based on the assumption that the emission caps are exogenous in the regional level.
 3 A Dynamic Regional Model of Climate and the Economy with an ETS
-This section introduces a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions.
-Our model framework extends the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps.
+We now introduce a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions.
+Our model framework extends the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps.
 The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption.
@@
 Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions.
-Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial time period is the year 2020.
+Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial year is the year 2020.
 All regions are forward-looking with complete information.
 3.1 Emissions
-The total industrial carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
-$E_{i,t}^{\mathrm{Total}}=\sigma_{i,t}Q_{i,t},$ (1)
-where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al.
-(2019).
-We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$.
-[^5] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al., 2006; Bauer et al., 2012; Golosov et al., 2014; Baldwin et al., 2020).
+The gross carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
+$E_{i}^{\mathrm{Gross}}=\sigma_{i,t}Q_{i,t},$ (9)
+where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al., (2019).
+We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the gross emissions, with an emission control rate $\mu_{i,t}\in[0,1]$.
+[^9] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020).
 However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated.
@@
 $E_{i,t}^{A}=\mu_{i,t}\sigma_{i,t}Q_{i,t}.$ (10)
-Thus the net emissions, $E_{i,t}$, becomes
-$E_{i,t}=E_{i,t}^{\mathrm{Total}}-E_{i,t}^{A}.$ (3)
+Thus the emissions net of abatement, $E_{i,t}$, becomes
+$E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}.$ (11)
 We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions.
-Regions with net emissions exceeding the emission caps can purchase emission permits from other regions.
+Regions with emissions exceeding the emission caps can purchase emission permits from other regions.
 Regions with emissions below their allowance can sell their emission permits to others.
@@
 In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$.
-The Paris Agreement establishes upper limits for net emissions to ensure that the increase in global temperature remains well below 2 degrees Celsius above preindustrial levels by the end of this century.
+The Paris Agreement establishes upper limits for emissions to ensure that the increase in global temperature remains well below 2°C above pre-industrial levels by the end of this century.
 3.2 Climate System
@@
 $T_{t}=\zeta\mathcal{E}_{t},$ (13)
-where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2 degrees Celsius above the pre-industrial level.
-Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al., 2011).
-Figure A.1 in Appendix A.3 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways.
+where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2°C above the pre-industrial level.
+Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011).
+Figure A.1 in Appendix A.3.1 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways.
 Cumulative global emissions evolve according to
@@
 $u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (15)
-where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus (2017) and set $\gamma=1.45$.
+where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus, (2017) and set $\gamma=1.45$.
 The production technology follows a Cobb-Douglas function of capital and labor inputs. The gross output (in trillion USD), or pre-damage output, $Q_{i,t}$, is given by
 $Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (16)
-where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus (2017)) is the elasticity of gross output with respect to capital.
-The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the no-climate-impact GDP projections in Burke et al.
-(2018).
-Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz, 2017).
+where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus, (2017)) is the elasticity of gross output with respect to capital.
+The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the GDP projections in Burke et al., (2018).
+Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz,, 2017).
 We denote $Y_{i,t}$ as the output net of damages induced by climate changes:
 $Y_{i,t}=\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}Q_{i,t},$ (17)
-where $T_{t}$ is the global average temperature increase in degrees Celsius above the preindustrial level.
-We calibrate the climate damage parameters based on the projections of GDP loss in Kahn et al.
-(2021).
-We follow DICE and assume emission abatement cost is
-$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$
+where $T_{t}$ is the global average temperature increase in °C above the pre-industrial level.
+We calibrate the climate damage parameters based on the projections of GDP in Burke et al., (2018).
+We follow DICE and assume that the emission abatement cost is
+$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ (18)
 where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$.
-We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al.
-(2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
+We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al., (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
 The total consumption of each region under the ETS is obtained by
@@
 where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits.
-The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption.
+The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption, but this impact is little as the cost or profit from permit trading is small relative to the consumption or output.
 The capital stock evolves according to the standard neoclassical growth model:
@@
 4 Solving for the Equilibrium
-Based on the model components introduced in Section 2, we now define equilibrium under cooperative and noncooperative models in the ETS.
-These two extreme cases place bounds on what we would observe in a real world scenario.
-3.1 The Cooperative and the Noncooperative Model
-In the cooperative model, all regions work together to maximize their combined welfare.
-More specifically, the objective function is to maximize the overall sum of discounted social utility across all regions as follows:
-${\color[rgb]{0,0,0}\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\sum_{i\in\mathcal{I}}\beta^{t}u(c_{i,t})L_{i,t}\end{split},}$ (13)
+Building on the model components outlined in Section 3, we now define the noncooperative equilibrium of our multi-region dynamic model with the ETS and present an algorithm used to obtain the Nash Equilibrium solution.
+4.1 The Noncooperative Equilibrium
+In the noncooperative model, the regional social planner of each region maximizes the region’s own lifetime social welfare. The maximization problem for each region $i$ is defined as
+$\max_{c_{i,t},E_{i,t}^{P},\mu_{i,t}}\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (22)
 where $\beta$ is the discount factor.
-We follow DICE-2016R (Nordhaus, 2017) and set $\beta=0.985$.
-Note that under cooperation there is no market equilibrium permit price, so we assume permit prices follow an exogenous path.
-Therefore, we define the cooperative model equilibrium given an exogenous path of permit prices.
-DEFINITION:
-Given the initial capital and cumulative global emissions, $\{K_{i,0},\mathcal{E}_{0}:\ i\in\mathcal{I}\}$, and the exogenous paths of emission caps and permit prices, $\{\overline{E}_{i,t},m_{t}:\ i\in\mathcal{I},t\geq 0\}$, the dynamic equilibrium for the cooperative model is a sequence of quantities $\{c_{i,t},E_{i,t}^{P},\mu_{i,t},K_{i,t},\mathcal{E}_{t},T_{t}:\ i\in\mathcal{I},t\geq 0\}$ that solve the maximization problem (13) subject to equations (7)-(12).
-In the noncooperative model, individual regions maximize their own lifetime social welfare, which represents another extreme but relatively more realistic perspective.
-The maximization problem for each region $i$ is defined as
-$\begin{split}\underset{c_{i,t},E_{i,t}^{P},\mu_{i,t}}{\max}&\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t}\end{split}.$ (14)
+We follow DICE-2016R (Nordhaus,, 2017) and set $\beta=0.985$.
+Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simulatenously as a dynamic game.
+Then we define the dynamic Nash equilibrium of the economy as follows.
 DEFINITION:
@@
 The ETS promotes efficient abatement globally by encouraging regions with better abatement technology or capacity (thus, with lower abatement cost) to conduct more abatement, and regions with less efficient abatement technology to purchase permits from other regions.
+Note that the equilibrium concept in our noncooperative model is an Open-Loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state.
+In an OLNE, regions commit to the strategies over time for their decision variables—consumption ($c_{i,t}$), emission purchase ($E_{i,t}^{P}$), and emission control rate ($\mu_{i,t}$)— at the initial period and cannot change their behavior over time.
+This concept contrasts with the Markov Perfect Equilibrium (MPE), where regions may make multiple decisions over time, allowing adaptation in their strategies.
+While the OLNE concept may be less satisfactory than the MPE concept (since OLNE is not subgame perfect), it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
 4.2 The Algorithm for the Noncooperative Model
@@
 Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years.
+Let
+$V_{i,300}(K_{1,300},...,K_{12,300},\mathcal{E}_{300})=u\left(\frac{0.75Y_{i,300}}{L_{i,300}}\right)\frac{L_{i,300}}{1-\beta}$
+be a terminal value function at the terminal year 300, which approximates the present value of utilities after 300 years, assuming that consumption at any $t\geq 300$ is equal to 75 percent of the output $Y_{i,300}$ at $t=300$ and that the exogenous population after 300 years stays at its value at $t=300$.
+Note that $Y_{i,300}$ is computed with a function of the terminal state $K_{i,300}$ and $\mathcal{E}_{300}$.
 Thus, we can transform the infinite horizon models to finite horizon models, where region $i$’s social welfare is rewritten as
-$\sum_{t=0}^{299}\beta^{t}u(c_{i,t})L_{i,t}+\beta^{300}u\left(\frac{0.75Y_{i,300}}{L_{i,300}}\right)\frac{L_{i,300}}{1-\beta}.$
-Here the second term is an approximation of the sum of discounted utility from time $t=300$ to infinity, assuming that consumption at any $t\geq 300$ is identical to 75 percent of the output at $t=300$ and that the exogenous population after 300 years stays at its value at $t=300$.
+$\sum_{t=0}^{299}\beta^{t}u(c_{i,t})L_{i,t}+\beta^{300}V_{i,300}(K_{1,300},...,K_{12,300},\mathcal{E}_{300}),$
+We also numerically verify that this time horizon truncation at 300 years has little impact on the OLNE solution in the first 100 years, by solving the same model but with a time horizon truncation at 400 years.
+This time horizon truncation method is common in solving infinite-horizon non-stationary models.
+For example, DICE-2016 truncates its model’s infinite horizon to 500 years, with a terminal value function being zero everywhere, for obtaining its numerical solution.
 The algorithm to solve the noncooperative model is as follows:
 Step
-1.
-Initialization.
-Give an initial guess of permit prices $\{m_{t}^{0}:\ t\geq 0\}$ and solve the cooperative model with emission cap and trade.
-The optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{i,t}^{0},E_{i,t}^{P,0}:\ t\geq 0\}$, for each region $i$.
-Iterate through steps 2, 3 and 4 for $j=0,1,...$, until convergence.
+1. Initialization. Set an initial guess of permit prices $\{m_{t}^{0}:\ t\geq 0\}$, and emissions $\{E_{i,t}^{0}:\ t\geq 0\}$. Iterate through steps 2, 3 and 4 for $j=1,2,...$, until convergence.
 Step
 2.
-Maximization Step.
-Solve the maximization problem (14) without the market clearing condition (12) for each region.
-Given the permit prices $\{m_{t}^{j}:\ t\geq 0\}$, we solve the optimization problem (14) without the market clearing condition (12) for region 1, where we fix the other regions’ net emissions at $\{E_{i,t}^{j}:\ i=2,...,12,\ t\geq 0\}$.
-We then obtain region 1’s optimal solution, where the optimal net emissions and the amount of emission permits traded with the other regions are denoted as $\{E_{1,t}^{*},E_{1,t}^{P,j+1}:\ t\geq 0\}$ Similarly, we obtain $\{E_{i,t}^{*},E_{i,t}^{P,j+1}:\ t\geq 0\}$ for regions $i=2,...,12$.
+Maximization Step at iteration $j$.
+For each region $i$, solve the maximization problem (22) without the market clearing condition (21), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration $j-1$ when $j>1$.
+The optimal emissions and permits purchased for region $i$ are denoted $\{E_{i,t}^{*,j},E_{i,t}^{P,j}:\ t\geq 0\}$.
 Step
-3. Update the emission permit prices and net emissions. After solving for the optimization problem of all regions, update the permit prices and net emissions as
-$m_{t}^{j+1}$ $=$ $m_{t}^{j}\exp\left(\omega\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}\right),$
-$E_{i,t}^{j+1}$ $=$ $\omega E_{i,t}^{*}+(1-\omega)E_{i,t}^{j},\ \ \forall i\in\mathcal{I},$
-where $\omega=0.1$ is a weight parameter and $\sum_{i\in\mathcal{I}}E_{i,t}^{P,j+1}$ is the net quantity of traded emission permits.
+3. Update Step at iteration $j$. After solving for the optimization problem of all regions respectively in Step 2, update the permit prices and emissions as
+$m_{t}^{j}$ $=$ $m_{t}^{j-1}\exp\left(\omega\sum_{i\in\mathcal{I}}E_{i,t}^{P,j}\right),$
+$E_{i,t}^{j}$ $=$ $\omega E_{i,t}^{*,j}+(1-\omega)E_{i,t}^{j-1},\ \ \forall i\in\mathcal{I},$
+where $\omega=0.1$ is a weight parameter and $\sum_{i\in\mathcal{I}}E_{i,t}^{P,j}$ is the net quantity of traded emission permits.
 Step
@@
 Check the convergence criterion.
-Check if $m_{t}^{j+1}\simeq m_{t}^{j}$, $E_{i,t}^{j+1}\simeq E_{i,t}^{j}$, and $E_{i,t}^{P,j+1}\simeq E_{i,t}^{P,j}$ for every region $i$ and $t\geq 0$.
-If so, stop the iteration.
-Note that $m_{t}^{j+1}=m_{t}^{j}$ implies that the market clearing condition (12) holds at the solution.
+Check if $m_{t}^{j}\simeq m_{t}^{j-1}$, $E_{i,t}^{j}\simeq E_{i,t}^{j-1}$, and $E_{i,t}^{P,j}\simeq E_{i,t}^{P,j-1}$ for every region $i$ and $t\geq 0$.
+If so, stop the iteration, otherwise go to Step 2 by increasing $j$ with 1.
+Note that $m_{t}^{j}=m_{t}^{j-1}$ implies that the market clearing condition (21) holds at the solution.
 This algorithm embodies the concept of market equilibrium.
-When there is a positive net quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices.
-Conversely, when there is a negative net quantity of traded emission permits, indicating excess supply, we lower the permit prices.
+When there is a net positive quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices.
+Conversely, when there is a net negative quantity of traded emission permits, indicating excess supply, we lower the permit prices.
 This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears.
@@
 In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution.
-Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state.
-The concept of the OLNE assumes the regions enter an agreement to commit to the solution path over time at the initial period.
-While this type of equilibrium concept might not be as satisfactory as the feedback Nash equilibrium concept, it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
-4 SCC, MAC, and the Optimal Carbon Tax
-In this section, we first define the regional social cost of carbon (SCC) and marginal abatement cost (MAC) and explain how the optimal carbon tax can be introduced together with the global ETS, in relation to the SCC and MAC.
-4.1 SCC and MAC
-The social cost of carbon (SCC), a central concept in the climate change literature, is widely used to quantify the monetary value of climate damages induced by an additional unit of carbon emissions.
-While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region.
-Similar to van der Ploeg and de Zeeuw (2016) and Cai et al.
-(2023), we define the cooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
-$\mathrm{SCC}_{i,t}^{\mathrm{Coop}}=\frac{-1,000(\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t})}{\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}},$ (15)
-for region $i$, where $V_{t}^{\mathrm{Coop}}$ is the value function of the cooperative model at time $t$, depending on the state variables $\{K_{i,t},\mathcal{E}_{t}:\ i\in\mathcal{I}\}$; that is,
-$V_{t}^{\mathrm{Coop}}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{t}\sum_{i=1}^{12}u(c_{i,s})L_{i,s}.$
-Similarly, the noncooperative SCC of a region is defined as
-$\mathrm{SCC}_{i,t}^{\mathrm{OLNE}}=\frac{-1,000(\partial V_{i,t}^{\mathrm{OLNE}}/\partial\mathcal{E}_{t})}{\partial V_{i,t}^{\mathrm{OLNE}}/\partial K_{i,t}},$ (16)
-where
-$V_{i,t}^{\mathrm{OLNE}}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{t}u(c_{i,s})L_{i,s},$
-for each region $i$ under the OLNE.
-Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon.
-To compute the regional SCC in the cooperative or the noncooperative model, it is equivalent to replace the numerator in equation (15) or (16) with the shadow price of the transition equation of cumulative global emissions (6) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (11) for each region.
-Our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al.
-(2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous.
-We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the OLNE.
-With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional net emissions is not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional net emissions is binding at zero.
-This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al.
-(2017) and Cai and Lontzek (2019).
-Moreover, we also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, but our social discount rates are endogenous.
-Similarly, the regional cooperative SCC in our definition is equal to the optimal regional carbon tax under cooperation, when the regional net emissions is not binding at zero and the exogenous emission permit price is zero (or there is no ETS).
-However, when the exogenous emission permit price is nonzero, the regional cooperative SCC is not equal to the optimal regional carbon tax under cooperation, because the ETS can be used as a means for transferring capital between regions to improve the global welfare under cooperation.
-For instance, a developed region can purchase a large amount of permits from a developing region even when its net emissions has been below its emission cap.
-In addition, the regional cooperative SCC is not equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, because a developed region will abate more with a high carbon tax or cooperative SCC to improve the global welfare under cooperation.
-The marginal abatement cost (MAC) is another important concept that captures the additional cost incurred due to an increase in emission abatement.
-From our model equation (2), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD.
-Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows
-$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$
-4.2 Optimal Carbon Tax under the ETS
-In this subsection, we discuss the relationship between the MAC and carbon taxation under the global ETS.
-While carbon taxes and ETS have often been considered as substitutable policies, we show that each region can internally choose an optimal carbon tax under the global ETS as a complementary policy tool to obtain the optimal level of net emissions.
-Before we discuss the optimal carbon tax, the mechanism behind the emission permit buyer and seller regions needs to be understood.
-Here we drop the region and time indices for simplicity.
-First of all, the relationship between the regional social planner’s optimal net emissions (denoted $E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller.
-A region becomes a permit buyer if its optimal net emissions are greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement.
-On the other hand, a region becomes a permit seller if its optimal net emissions are lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$).
-Note that this is the optimal solution from the perspective of each region’s social planner.
-One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emissions level.
-We assume that the social planner uses carbon taxation as the policy instrument to fix the inconsistency between the optimal net emissions of the regional social planner and the firms.
-Let us consider a representative firm in each region that produces emissions and faces the MAC function and the emission cap of the region.
-We assume that the representative firm is myopic and maximizes its profit in each period, thus it does not consider the dynamic optimization problem of the social planner.
-Given the total emissions, $E^{\mathrm{Total}}$, the representative firm’s profit maximization problem is equivalent to choosing its optimal level of net emissions to minimize its total cost associated with emissions (i.e.
-the abatement cost and revenue or expenditures from emission permit purchases or sales).
-[^6] It is worth mentioning that a region’s total emissions is proportional to its GDP, which, in turn, is determined by capital and the exogenous path of population size.
-Considering that the current level of capital is determined in the previous period, total emissions can be considered as exogenously given from the perspective of the single-period myopic firm.
-Figure 1 depicts the relationship between emissions, permit price, carbon tax, and the MAC for buyer and seller regions.
-In Figure 1, the MAC curve is a decreasing function of net emissions ($E$),  [^7] Substituting $E^{A}=E^{\mathrm{Total}}-E$ from equation (3) into the MAC, it follows that $\mathrm{MAC}_{i,t}=1,000\left(\frac{b_{1,i,t}b_{2,i}}{\sigma_{i,t}}\right)\left(\frac{E_{i,t}^{\mathrm{Total}}-E_{i,t}}{\sigma_{i,t}Q_{i,t}}\right)^{b_{2,i}-1}$.
-which reach zero when net emissions equal total emissions ($E^{\mathrm{Total}}$).
-That is, the higher the emission abatement, the smaller the net emissions, resulting in a higher MAC.
-The market equilibrium price of emission permits is given at $m$.
-Then in the absence of a carbon tax, there is a corresponding cut-off level of net emissions, $E^{m},$ above which conducting the emission abatement is more cost-effective, and below which purchasing emission permits from the permit markets is cheaper.
-Figure 1: The optimal carbon tax in emission permit buyer and seller regions.
-In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$.
-Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits.
-That is, in the absence of a carbon tax, the social planner’s optimal net emissions, $E^{*}$, may not align with the optimal choice of the firm within a region, $E^{m}$, regardless of whether the region is a permit buyer or seller.
-For example, when $\bar{E}$ is very large for every region, the permit price $m$ is zero, so $E^{m}=E^{\mathrm{Total}}$, but it is well known in the literature that $E^{*}$ is often strictly smaller than $E^{\mathrm{Total}}$ (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
-Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at the price $m$ to increase its profits.
-This is also verified in our simulation results later on.
-We now show that a regional carbon tax can serve as a supplementary policy for the representative firm to reach the social planner’s optimal net emissions at $E^{*}$.
-Assume that the social planner levies a tax on the firm of $\tau$ per unit of net emissions.
-Given the total emissions, $E^{\mathrm{Total}}$, the cost minimization problem of the representative firm that faces this carbon tax under the global ETS is as follows:
-[…]
-The first term of the objective function is the total tax on net emissions, the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E^{\mathrm{Total}}-E)$.
-The above cost minimization problem holds for both emission permit buyer and sellers.
-For simplicity, we assume that emissions cannot be negative in our model.
-From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm (denoted $E^{*,\mathrm{Firm}}$) leads to the following relationship between the carbon tax, the MAC, and the permit price:
-$\tau=\mathrm{MAC}(E^{*,\mathrm{Firm}})-m,$ (18)
-when $E^{*,\mathrm{Firm}}>0$.
-Importantly, it implies that for the representative firm to attain the regional social planner’s optimal net emissions $E^{*}$ (i.e., $E^{*,\mathrm{Firm}}=E^{*}$), the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$.
-In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal net emissions and the market price of emission permits.
-It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*,\mathrm{Firm}}=E^{*}$, rather than $E^{m}$.
-When $E^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
-$\tau\geq\mathrm{MAC}(E^{*,\mathrm{Firm}})-m.$ (19)
-That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits.
-In particular, the regional $\mathrm{MAC}(E^{*,\mathrm{Firm}})$ at $E^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m$.
-However, when $E^{*}=0$, a nonnegative carbon tax larger than $\mathrm{MAC}(E^{*})-m$ will not make $E^{*,\mathrm{Firm}}$ to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}(E^{*})-m$ if it is positive.
-Therefore, for any regional social planner’s optimal net emissions $E^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
-$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\},$ (20)
-such that the firm’s optimal net emissions are consistent with the social planner’s (i.e., $E^{*,\mathrm{Firm}}=E^{*}$).
-The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model.
-Thus, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
-4.3 Optimal Carbon Tax under no ETS
-For the case without the ETS, the cost minimization problem of the representative firm facing the emission cap regulation becomes:
-[…]
-From the Karush-Kuhn-Tucker conditions of (21), when the firm’s optimal solution $E^{*,\mathrm{Firm}}=E^{*}$ is not binding at its lower or upper bound, we have the well known equality:
-$\tau=\mathrm{MAC}(E^{*}).$ (22)
-That is, the optimal carbon tax should be chosen to be equal to the MAC at the social planner’s optimal emissions $E^{*}$ if it is not binding.
-This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019).
-We can also see that when $\overline{E}$ is very large or the cap does not exist, a nonzero $E^{*}$ will not be binding, so the equality holds, then the MAC is nonzero.
-In the case with the ETS and very large caps for every region, which is equivalent to the case without the ETS nor the caps, the permit price should be zero, and so our formula (20) is consistent with (22).
-However, the assumption that the MAC is equal to the permit price, which is often used in the literature (e.g., Keohane, 2009; Nordhaus, 2010), does not hold in this case, unless $E^{*}=E^{\mathrm{Total}}$ (i.e., the social planner’s optimal decision is to have zero abatement), which usually does not appear in those non-business-as-usual solutions in the literature (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
 5 Data and Calibration
 To calibrate our model with data, we have two objectives.
-The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their commitments to the Paris Agreement and the net zero emission targets.
-The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al., 2018; Ueckerdt et al., 2019; Kahn et al., 2021).
+The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their emissions commitments under the Glasgow Pact and the long-term net zero emission targets.
+The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$, $b_{4,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al.,, 2018; Ueckerdt et al.,, 2019; Kahn et al.,, 2021).
 We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
-Our model’s 12 regions are formed by aggregating 198 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010).
+Our model’s 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ).
 For the initial year 2020, we obtain country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (GtC) from the World Bank.
-For future projections on population growth, we use the SSP2 scenario (Samir and Lutz, 2017).
+For future projections on population growth, we use the SSP2 scenario (Samir and Lutz,, 2017).
 5.1 Regional Emission Cap Pathways
-Since there is no global cap-and-trade or ETS currently in place, we use the emission commitments of the Paris Agreement and the regions’ net zero commitments to construct the baseline emission cap constraint, $\overline{E}_{i,t}$.
-We collect reports of the current Nationally Determined Contributions (NDCs) of 165 countries that were first submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 and later updated (INDC, 2023; NDC, 2024).
-Although there is no international enforcement mechanism, most countries committed to achieving a self-defined mitigation target for 2030 (or 2025).
-We also obtain the target years to reach net zero emissions of different countries from Climate Action Tracker (2023).
+Since there is no global cap-and-trade system or ETS currently in place, we consider the emission commitments outlined in the Paris Agreement and the COP26 Glasgow Climate Pact.
+Under the Paris Agreement, 195 countries or regions set Nationally Determined Contributions (NDCs) that specify their near-term targets for 2025 or 2030, along with long-term net-zero commitments for 2050 to 2070.
+These near-term targets were further strengthened during the Glasgow Climate Pact in 2021, with most countries and regions submitting updated or new NDCs.
+[^10] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated Nationally Determined Contributions (NDCs) after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker.
+[^11] See https://climateactiontracker.org/.
 Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
-As the first step, we obtain the level of emission targets for each country.
-In their NDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level at some base year.
-Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in net emissions.
+As the first step, we obtain the near-term emission targets for each country.
+In their NDC reports, most countries express targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level at some base year.
+Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in emissions.
 For the countries that did not make a specific emissions reduction pledge, we assume that their carbon intensity reduction and emission reduction percentages are the same as those of the most populous country in that region.
-Next, we generate annual emission cap pathways based on the regions’ historical emission levels (World Bank, 2020), their emission targets for 2030 (or 2025), and net zero emission target years.
+Next, we generate annual emission cap pathways based on the regions’ historical emission levels (World Bank,, 2020), their emission targets for 2030 (or 2025), and net zero emission target years.
 We use five-year emissions data (2014 - 2018) from the World Bank and the INDC emission targets in 2030 (or 2025) to fit a quadratic function and use this fitted function to project the emission pathways for the periods between 2018 and 2030.
@@
 After obtaining the country-level emission pathways, we aggregate them to find the regional emission cap pathways.
-[^8] We also find that our aggregated regional emission cap pathways are close to those used in Nordhaus (2010).
+[^12] We also find that our aggregated regional emission cap pathways are close to those used in Nordhaus, 2010b .
 To assess the implications of varying stringency in the emission caps, we also create additional emission cap pathways by choosing alternative net zero scenarios, wherein we assume that all regions achieve net zero emissions by 2050 (most stringent), 2070, or 2090 (most lax).
-See Figures A.2 and A.3 in Appendix A.4 for the regional emission cap pathways in the baseline cap scenario and the different emission cap pathways at the global level.
-5.2 Total Factor Productivity
-We calibrate total factor productivity (TFP), $A_{i,t}$, for future periods based on the GDP projections in Burke et al.
-(2018), who provide the projected GDP of 165 countries until 2099 assuming no climate-related impacts.
-We aggregate these projections according to our 12 regions and employ the SSP2 population scenario to obtain regional GDP per capita estimates, $y_{i,t}^{\mathrm{BDD}}$, which are used to calibrate the regional TFP under no climate impacts, $A_{i,t}$.
+See Figures A.5 and A.6 in Appendix A.3.4 for the regional emission cap pathways in the baseline cap scenario and the different emission cap pathways at the global level.
+5.2 Total Factor Productivity and Climate Damage
+We follow Cai et al., (2023) to calibrate total factor productivity (TFP), $A_{i,t}$, and the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$, based on Burke et al., (2018), who provide the projected GDP of 165 countries till 2099 assuming no climate-related impacts, and their GDP assuming the contemporaneous climate impact of the RCP4.5 temperature $T_{t}^{\mathrm{RCP4.5}}$ [^13] We choose the RCP4.5 scenario instead of the other RCP scenarios, because the RCP4.5 scenario is the closest which covers the range of temperature in our solution.
+until 2049, under the SSP2 population pathway.
+[^14] It is nontrivial to use historical data to calibrate future TFP, particularly when we need to isolate the climate impacts from the data.
+For simplicity, we use the projected GDP for the calibration in this paper.
+We also use the TFP growth values in RICE to do sensitivity analysis, and find our results are still robust (see Appendix A.4.6).
+We aggregate these projections according to our 12 regions and employ the SSP2 population scenario to obtain regional GDP per capita estimates, $y_{i,t}^{\mathrm{BDD,NoCC}}$ under no climate impact and $y_{i,t}^{\mathrm{BDD}}$ under climate impact, which are used to calibrate the regional TFP, $A_{i,t}$, and the climate damage parameters, $\pi_{1,i}$ and $\pi_{2,i}$.
 For each region $i$, the dynamic path of the TFP is modeled by the relationship $A_{i,t+1}=A_{i,t}\exp\left(g_{i,t}\right)$, where $g_{i,t}$ is the growth rate of $A_{i,t}$ at time $t$.
@@
 Therefore, we follow RICE to generate $g_{t,i}$ for $t\geq 80$.
-See Appendix A.5 for the details.
-In our structural estimation, we obtain $(g_{i,0},d_{i})$ by solving the following minimization problem:
-$\min_{g_{i,0},d_{i}}\ \sum_{t=0}^{79}w_{t}\left(\frac{y_{i,t}^{\mathrm{NoCC}}}{y_{i,0}^{\mathrm{NoCC}}}-\frac{y_{i,t}^{\mathrm{BDD}}}{y_{i,0}^{\mathrm{BDD}}}\right)^{2},$ (24)
-where $w_{t}$ are weights, and $y_{i,t}^{\mathrm{NoCC}}$ is GDP per capita obtained by solving the following optimal growth model with a choice of $(g_{i,0},d_{i})$ and its associated TFP $A_{it}$:
-$\max_{c_{i,t}}$ $\sum_{t=0}^{500}\beta^{t}u(c_{i,t})L_{i,t},$ (25)
+See Appendix A.3.5 for the details.
+In our structural estimation, we obtain $(g_{i,0},d_{i},\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem:
+$\min_{g_{i,0},d_{i},\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{79}\left(\frac{y_{i,t}^{\mathrm{NoCC}}}{y_{i,0}^{\mathrm{NoCC}}}-\frac{y_{i,t}^{\mathrm{BDD,NoCC}}}{y_{i,0}^{\mathrm{BDD,NoCC}}}\right)^{2}+\sum_{t=0}^{29}\left(\frac{y_{i,t}}{y_{i,0}}-\frac{y_{i,t}^{\mathrm{BDD}}}{y_{i,0}^{\mathrm{BDD}}}\right)^{2}.$ (24)
+Here $y_{i,t}^{\mathrm{NoCC}}$ is GDP per capita obtained under no climate impact by solving the following optimal growth model with a choice of $(g_{i,0},d_{i},\pi_{1,i},\pi_{2,i})$ and its associated TFP $A_{it}$:
+$\max_{c_{i,t}}$ $\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (25)
 $\mathrm{s.t.}$ $K_{i,t+1}=(1-\delta)K_{i,t}+\left(y_{i,t}^{\mathrm{NoCC}}-c_{i,t}\right)L_{i,t},$
-where $y_{i,t}^{\mathrm{NoCC}}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{-\alpha}$ is the GDP per capita under no climate impact and $u(c_{i,t})$ is defined as in equation (7).
-The initial TFP $A_{i,0}$ is chosen such that $A_{i,0}K_{i,0}^{\alpha}L_{i,0}^{-\alpha}$ is equal to the observed GDP per capita in 2020.
-Figure A.4 in Appendix A.6 shows that with our calibrated $A_{i,t}$ , the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD}}$ from Burke et al.
-(2018) for all regions.
+where $y_{i,t}^{\mathrm{NoCC}}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{-\alpha}$ and $u(c_{i,t})$ is defined as in equation (15); and $y_{i,t}$ is GDP per capita under climate impact by solving the following optimal growth model:
+$\max_{c_{i,t}}$ $\sum_{t=0}^{\infty}\beta^{t}u(c_{i,t})L_{i,t},$ (26)
+$\mathrm{s.t.}$ $K_{i,t+1}=(1-\delta)K_{i,t}+\left(y_{i,t}-c_{i,t}\right)L_{i,t},$
+where
+$y_{i,t}=\frac{1}{1+\pi_{1,i}T_{t}^{\mathrm{RCP4.5}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP4.5}}\right)^{2}}A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{-\alpha}.$
+The initial TFP $A_{i,0}$ is chosen such that $A_{i,0}K_{i,0}^{\alpha}L_{i,0}^{-\alpha}/\left(1+\pi_{1,i}T_{0}+\pi_{2,i}T_{0}^{2}\right)$ is equal to the observed GDP per capita in 2020.
+Figures A.2 and A.3 in Appendix A.3.2 shows that with our calibrated $(g_{i,0},d_{i},\pi_{1,i},\pi_{2,i})$, the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ or $y_{i,t}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD,NoCC}}$ or $y_{i,t}^{\mathrm{BDD}}$ from Burke et al., (2018), respectively, for all regions.
 5.3 Carbon Intensity
-To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al.
-(2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation.
-As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun211’ in Ueckerdt et al.
-(2019).
-Specifically, this FFrun211 scenario corresponds to climate action from 2020 with full technology portfolio but with zero carbon tax.
-Based on the equation (1), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun211}}^{U}/Q_{i,t,\mathrm{FFrun211}}$, where $E_{i,t,\mathrm{FFrun211}}^{U}$ and $Q_{i,t,\mathrm{FFrun211}}$ are the projected regional emissions and GDP under the FFrun211 scenario for region $i$ at time $t$.
-[^9] Ueckerdt et al.
-(2019) provide data for 11 regions.
+To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al., (2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation.
+As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun451’ in Ueckerdt et al., (2019).
+Specifically, this FFrun451 scenario corresponds to weak policy baseline before Paris Agreement with limited renewable technology portfolio and zero carbon tax.
+Based on the equation (9), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun451}}^{U}/Q_{i,t,\mathrm{FFrun451}}$, where $E_{i,t,\mathrm{FFrun451}}^{U}$ and $Q_{i,t,\mathrm{FFrun451}}$ are the projected regional emissions and GDP under the FFrun451 scenario for region $i$ at time $t$.
+[^15] Ueckerdt et al., (2019) provide data for 11 regions.
 Upon comparison, we find that the countries constituting the ‘Rest of the World (ROW)’ region are the ones that are in the ‘Other High Income (OHI)’ and ‘Eurasia’ region in our study.
-Therefore, the carbon intensity obtained from the ROW in Ueckerdt et al.
-(2019) corresponds to that of OHI and Eurasia regions in our work.
+Therefore, the carbon intensity obtained from the ROW in Ueckerdt et al., (2019) corresponds to that of OHI and Eurasia regions in our work.
 5.4 Abatement Cost
-Our estimation of the abatement cost parameters $b_{1,i}$, $b_{2,i}$, and $b_{3,i}$ relies on the simulation results under ten different levels of carbon taxes in Ueckerdt et al.
-(2019).
-For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$ for every region $i$, Ueckerdt et al.
-(2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$.
-Since Ueckerdt et al.
-(2019) do not consider an ETS, according to the discussion in Section 4.2, we can assume that their carbon taxes $\tau_{t,j}^{U}$ are equal to the marginal abatement costs when net emissions are strictly positive.
+Our estimation of the abatement cost parameters $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$ relies on the simulation results under ten different levels of carbon taxes in Ueckerdt et al., (2019).
+For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$ for every region $i$, Ueckerdt et al., (2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$.
+Since Ueckerdt et al., (2019) do not consider an ETS, according to the discussion in Section 2, we can assume that their carbon taxes $\tau_{t,j}^{U}$ are equal to the marginal abatement costs when emissions are strictly positive.
 That is,
@@
 At first, we estimate the associated emission control rate $\mu_{i,t,j}^{U}$ using the equations (9)-(11) as follows:
-$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun211}}^{U}},$ (27)
-where $E_{i,t,\mathrm{FFrun211}}^{U}$ is the regional emissions from the FFrun211 scenario with zero carbon tax (see Ueckerdt et al.
-(2019) for details).
+$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun451}}^{U}},$ (28)
+where $E_{i,t,\mathrm{FFrun451}}^{U}$ is the regional emission from the FFrun451 scenario with zero carbon tax (see Ueckerdt et al., (2019) for details).
 Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (27) can hold in an approximate manner for every scenario $j$ and time $t$.
-Appendix A.8.6 also reports a sensitivity analysis on the estimate of the abatement cost.
-5.5 Climate Damage
-We calibrate the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$ by considering projections on GDP loss across different climate scenarios in Kahn et al.
-(2021), which shows the percentage loss in GDP per capita by 2030, 2050, and 2100 under the RCP 2.6 and RCP 8.5 scenarios for China, EU, India, Russia, and the US.
-We use their method and data to project the percentage loss in GDP per capita ($\Delta_{i,t}^{\mathrm{RCP26}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}$) every year from 2020 to 2114 under the RCP 2.6 and RCP 8.5 scenarios for each of our 12 regions, employing the baseline setup in Kahn et al.
-(2021).
-Specifically, $\Delta_{i,t}^{\mathrm{RCP26}}=1-y_{i,t}^{\mathrm{RCP26}}/y_{i,t}^{\mathrm{base}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}=1-y_{i,t}^{\mathrm{RCP85}}/y_{i,t}^{\mathrm{base}}$, where $y_{i,t}^{\mathrm{RCP26}}$, $y_{i,t}^{\mathrm{RCP85}}$, and $y_{i,t}^{\mathrm{base}}$ are GDP per capita under RCP 2.6, RCP 8.5, and the baseline scenario, respectively.
-Thus from equation (9) we obtain $(\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem for each region $i$:
-$\min_{\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{94}\left(\frac{1+\pi_{1,i}T_{t}^{\mathrm{RCP85}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP85}}\right)^{2}}{1+\pi_{1,i}T_{t}^{\mathrm{RCP26}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP26}}\right)^{2}}-\frac{1-\Delta_{t,i}^{\mathrm{RCP26}}}{1-\Delta_{t,i}^{\mathrm{RCP85}}}\right)^{2}.$ (28)
-Here $T_{t}^{\mathrm{RCP26}}$ and $T_{t}^{\mathrm{RCP85}}$ are the global average temperature anomalies at time $t$ (deviation from the pre-industrial temperature) under the RCP 2.6 and RCP 8.5 scenarios.
-Figure A.5 in Appendix A.7 shows that with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al.
-(2021), for every region.
 6 Results
 In this section, we discuss our simulation results.
-We first present the results of the noncooperative model under the baseline emission cap scenario for 2020 to 2100.
-Next, we compare the economic outcomes of different models: noncooperation versus cooperation, with or without the emission caps, and with or without the global ETS.
-Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all countries.
-6.1 Noncooperation with ETS under the Baseline Emission Caps
+We first present the results of the noncooperative model under the baseline emission cap scenario for the period from 2020 to 2100.
+Next, we compare the economic outcomes across different models, examining the role of the emission caps and the implementation of a global ETS.
+Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all 12 regions.
+6.1 Key Results under the Baseline Emission Caps
 Figure 1 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario.
 Until 2100, we can delineate three periods according to the global emission trading patterns.
-Prior to 2030, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left).
-This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices.
+Prior to 2025, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5).
+This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al.,, 2018), resulting in zero or very low permit prices.
+This is also consistent with our theoretical analysis for the regional social planners’ Nash equilirbium in Section 2.
 During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 1 top-left and bottom-right).
 This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
-Global emissions are constrained by the global emission caps starting from 2030.
+Global emissions are constrained by the global emission caps starting from 2025.
 To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits.
-This results in a substantial increase in the volumes of both emission abatement and traded permits, along with their associated costs.
-The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left).
-The global abatement cost also increases steeply to reach its maximum value of \$5.18 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
-The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left).
-As the emission cap becomes tighter over the years, the permit price increases to \$811 per ton of carbon in 2050 and \$2,254 in 2069 (Figure 2 bottom-left).
-The kink in the permit price path in 2050 is a result of some regions achieving net zero emissions.
-As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050s for Japan (2050) and EU (2053), 2060 for Russia and China, and 2070 for the rest of regions.
+This results in a substantial increase in the volume of emission abatement, along with the abatement cost and permit price.
+The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2060 (Figure 1 top-left).
+The global abatement cost also increases steeply to reach its maximum value of \$5.45 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
+The emission permit trade gradually decreases to zero by 2070 (Figure 1 top-left).
+As the emission cap becomes tighter over the years, the permit price increases to \$845 per ton of carbon in 2050 and \$2,059 in 2069 (Figure 1 bottom-left).
+The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5.
 Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
-Figure 2: Simulation results at the global scale for the noncooperative model with emission trading under the baseline emission cap scenario.
+Figure 1: Simulation results at the global scale under the baseline emission cap scenario.
 Post-2070 is the period when global emissions are net zero.
@@
 [^16] In the bottom-left panel of Figure 1, we plot the emission permit price only when the traded volume is positive.
-By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right), restricted by the emission caps.
+By the end of this century, the atmospheric temperature is projected to reach 1.7°C above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps.
 [^17] If there is no emission cap, then the temperature anomaly will be much higher.
 This can be seen in a later discussion with alternative emission caps.
-Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right).
-Overall, our noncooperative model simulation predicts that the global emission caps set by the Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
+Lastly, the global climate damage rises almost linearly over the entire period (Figure 1 bottom-right).
+Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5°C.
 It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers.
 Figure 2 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales.
-In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage.
-After 2029, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Africa, and OHI regions, with the US, EU, and India being the largest permit buyers.
-The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider.
-Japan is expected to be involved in a relatively small volume of permit trading.
-Figure 3: Simulation results of regional emission permit purchase for the noncooperative model with emission trading, under the baseline emission cap scenario.
-Based on our simulation results, we now discuss the optimal regional carbon tax under the noncooperative model with the ETS and the baseline emission cap scenario.
-Figure 4 shows the regional MAC along with the market equilibrium price of emission permits.
-We find that the MAC of every region increases drastically and remains strictly greater than the permit price until 2050, and the net emissions are strictly positive for all regions during this time.
-[^12] For regional optimal net emissions, see Figure 3 in Appendix A.8.1.
-As discussed in Section 4.2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ is equivalent to $E_{i,t}^{*}<E_{i,t}^{m}$, implying that when there is no carbon tax, the optimal net emissions for the social planner do not align with the optimal net emissions for the representative firm in each region.
-Therefore, the optimal carbon tax can be determined by $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s net emissions are equal to the social planner’s optimal net emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050.
-Figure 4: Simulation results of regional MAC for the noncooperative model with emission trading under the baseline emission cap scenario.
-In Figure 4, the MAC of each region gradually decreases below the permit price after its net zero emissions (before trading of permits) is achieved.
-Russia is the first region to achieve net zero emissions (before trading of permits) in 2051 (as shown in Figure A.6 in Appendix A.8.1).
-But Russia still sells permits afterwards, as shown in Figure 3, because its after-trade net zero emission target year is 2060 under the baseline emission cap scenario.
-After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound.
-Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China, Latin America, and the OHI) and the 2060s (the US, the EU, Japan, Eurasia, India, and MidEast).
-Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
+Before 2025, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage.
+After 2025, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2036.
+Japan is expected to be involved in a relatively small volume of permit trading, and the other regions (Africa, MidEast, Latin America, and OthAs) change their status of permit suppliers to buyers or vice versa over time.
+Figure 2: Simulation results of regional emission permit purchase under the baseline emission cap scenario.
+6.2 SCC, MAC, and the Optimal Carbon Tax
+Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the MAC, regional social cost of carbon (SCC), and the optimal carbon tax under the global ETS regime.
+The MAC captures the additional cost incurred due to an increase in emission abatement.
+From our model equation (10), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD.
+Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows:
+$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (29)
+With $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^18] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters.
+Figure 3 displays the regional MAC along with the market equilibrium price of emission permits.
+We find that, for every region, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s.
+[^19] For regional optimal emissions, see Figure 11 in Appendix A.4.1.
+As discussed in Section 2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ implies that the optimal emissions for the regional social planner do not align with the optimal emissions for the representative firm under no tax in each region.
+Therefore, the optimal regional carbon tax can be introduced at $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s emissions are equal to the regional social planner’s optimal emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050 or 2070.
+Russia is an exception, where the MAC falls below the permit price starting in 2025, even though it has nonzero emissions until 2050, because global warming leads to increase in productivity rather than causing climate damages in Russia.
+This result is strongly influenced by the climate damage parameters, calibrated using projections from Burke et al., (2018).
+Figure 3: Simulation results of regional MAC under the baseline emission cap scenario.
+In Figure 3, the MAC of each region gradually decreases below the permit price after its net zero emissions is achieved (i.e., $E_{i,t}^{*}=0$).
+Russia is the first region to achieve net zero emissions in 2050 (as shown in Figure A.7 in Appendix A.4.1).
+However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario.
+After 2050, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit.
 [^20] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large.
-Recall that, when the regional net zero emissions (before trading of permits) are achieved, the optimal carbon tax should be larger than the gap between MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$.
-Note that if a region’s after-permit-trade zero emission cap constraint is binding, then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
-Figure 5: Simulation results of regional SCC for the noncooperative model with emission trading under the baseline emission cap scenario.
-We further confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS.
-Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4.
-For example, in 2050, the MAC for the US is \$971 per ton of carbon, the permit price is \$811, and their difference is exactly equal to the regional SCC of the US, \$160.
-Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative.
-Among all the regions, the US has the highest SCC, starting at \$106 per ton of carbon in 2020 and steadily increasing up to \$213 per ton of carbon by the end of this century.
-The US is followed by the EU (from \$42 to \$139 per ton of carbon) and China (from \$39 to \$117 per ton of carbon) in this century.
-The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon).
-Our results show that there is considerable heterogeneity in the SCC across regions and suggest that a significant amount of regional carbon tax implementation is required, particularly for the US, the EU, and China, under the global ETS with the baseline emission cap scenario.
-6.2 Comparative Analysis of Different Cases
-We now modify our baseline model with various cases to understand the implications of cooperation among regions and the ETS.
-We compare outcomes between the noncooperative and the cooperative models, and for each model, we examine the impact of the ETS under the baseline emission caps.
-Figure 6 displays the global emission and temperature time profiles for all the cases.
-Note that for the cooperative model with the ETS, we use the equilibrium solution of permit prices obtained under the noncooperative model with the ETS, so we can compare the difference in the solutions between cooperation and noncooperation, holding permit prices fixed across scenarios.
-Figure 6: Simulation results of different cases under cooperation or noncooperation.
-We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation.
-Specifically, cooperation results in approximately 0.17 degrees Celsius lower temperature with the ETS and 0.13 degrees Celsius lower temperature without the ETS.
-Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps.
-Moreover, the ETS has little impact on the temperature path, by reducing the temperature anomaly in 2010 with only 0.02 degrees Celsius, as the emission caps exist no matter whether the ETS exists or not.
-When comparing the cases with and without the ETS, [^14] Note that the emission caps are still there, no matter whether the ETS is implemented or not.
-we find that the ETS may play different roles under noncooperation and cooperation.
-In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS.
-This is because with the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level.
-This happens until 2050.
-In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time.
-With the ETS, lower levels of global emissions are observed after 2030.
-This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
-We also compare the regional MAC and SCC across the different cases.
-Figure 7 presents the regional MAC and SCC under both the cooperative and the noncooperative models with or without the ETS, using the US as an illustrative example.
-[^15] For comparisons of other regions, refer to Appendices A.8.4 and A.8.3.
-Under cooperation, the MAC of the US (Figure 7 top-left) is very large in the initial years.
-This is because the US, as a developed region, needs to mitigate a significant amount of emissions so that developing regions can mitigate less (as shown in Figure A.9 in Appendix A.8.3), thus improving global social welfare.
-After the 2030s, however, the US net emissions become zero in the cooperative case [^16] For regional net emissions under cooperation and noncooperation, see Appendix A.8.2.
-and cannot be further mitigated, so its MAC starts to decline over time.
-The effect of the ETS under cooperation depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm).
-Moreover, the ETS has little impact on the net emissions and the MAC of a developed region as shown in Figures A.7 and A.9 (in Appendices A.8.2 and A.8.3) respectively, because a developed region will abate significantly no matter whether the ETS exists or not under cooperation.
-That is, the ETS plays an important role only in developing regions under cooperation.
-Figure 7: Simulation results of different cases for the US in the cooperative or noncooperative model.
-The top-right panel of Figure 7 shows that the cooperative SCC of the US is sustained at a high level throughout all periods.
-This is consistent with the high MAC of the US.
-When the US regional net emissions are not binding at zero before 2030 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2030.
-Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions.
-This also implies that a developed region will contribute more to mitigating emissions under cooperation.
-This pattern is consistent across regions: Figure A.11 in Appendix A.8.4 shows that the US, EU, Japan and OHI have a high cooperative SCC, while the developing regions like India and Africa have a small cooperative SCC.
-Under noncooperation, the MAC of the US is substantially lower in the case with the ETS compared to the case without the ETS (Figure 7, bottom-left panel).
-Such discrepancy persists until 2058, the period during which the US is a net permit buyer (shown in Figure 3), reflecting the fact that implementation of the ETS offers the US a more cost-effective means of keeping its net emissions below its cap.
-Consequently, the US conducts less emission abatement.
-For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2030 (see Figure A.10 in Appendix A.8.3).
-The bottom-right panel of Figure 7 shows that the noncooperative SCC of the US is much smaller than its cooperative SCC, which is also consistent with the smaller MAC of the US under noncooperation.
-Moreover, every region has a much smaller SCC under noncooperation than under cooperation, as shown in Figures A.11 and A.12 in Appendix A.8.4.
-This result implies that, in the dynamic Nash equilibrium under noncooperation, no region is willing to impose a high carbon tax thereby allowing other regions to free ride on lower emissions and climate damages.
-Therefore, in the noncooperative model, emissions and temperature rises are much larger than in the cooperative model, as shown in Figure 6.
-Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS.
-For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2033 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2033 (see the two top panels of Figure 7).
-For the noncooperative case without the ETS, the US regional net emissions are binding at their caps after 2022 (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC after 2022 (see the two bottom panels of Figure 7).
-Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Eurasia under cooperation until 2060 in Figures A.7, A.9, A.11; and Eurasia under noncooperation until 2049 in Figures A.8, A.10, and A.12), which is equal to the optimal regional carbon tax.
-This relationship holds for every region as shown in Figures A.7-A.12.
-The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under cooperation or noncooperation.
-This pattern holds for every region as shown in Figures A.11 and A.12.
-This happens under noncooperation because the regional noncooperative SCC is the present value of of future climate damages in the region resulting from an additional unit of global emissions released in the current period, and the ETS has little impact on the marginal damages and our endogenous social discount rates.
-Under cooperation, since the ETS has little impact to developed regions and the temperature path, the wealth redistribution among developing regions via the ETS has little impact on the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$,, so the ETS has little impact on the regional cooperative SCC.
-This shows that the regional SCC is complementary to the ETS, as the emission caps exist no matter whether the ETS exists or not.
-6.3 Alternative Net Zero Scenarios
-Along with the baseline emission cap scenario, we further analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090.
+Similar trends are followed by other regions as they attain net zero emissions in the 2050s ( China, the US, the EU, the OHI, MidEast, Eurasia, and Latin America) and the 2060s (Japan and India).
+Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
+Recall that, when the regional net zero emissions (before trading of permits) are achieved (i.e., $E_{i,t}^{*}=0$), the regional optimal carbon tax should be larger than the gap between the regional MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$.
+Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (12) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
+Given the simulation results for the optimal carbon tax, we now turn to discuss the relationship between the optimal carbon tax and the regional social cost of carbon (SCC).
+The SCC, a central concept in the climate change literature, is widely used to quantify the present value of climate damages induced by an additional unit of carbon emissions.
+While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region.
+Similar to van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), we define the noncooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
+$\mathrm{SCC}_{i,t}=\frac{-1,000(\partial V_{i,t}/\partial\mathcal{E}_{t})}{\partial V_{i,t}/\partial K_{i,t}},$ (30)
+where $V_{i,t}$ is the value function of the noncooperative model at time $t$ for region $i$, depending on the state variables $\{K_{i,t},\mathcal{E}_{t}:\ i\in\mathcal{I}\}$; that is,
+$V_{i,t}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{s-t}u(c_{i,s})L_{i,s},$
+for each region $i$ under the open loop Nash equilibrium.
+Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon.
+[^21] To compute the regional SCC in the noncooperative model, it is equivalent to replace the numerator in equation (30) with the shadow price of the transition equation of cumulative global emissions (14) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (20) for each region.
+Our concept of the regional SCC differs from the definition in Nordhaus, (2017) and Ricke et al., (2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous.
+We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the Nash equilibrium.
+With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional emissions are not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional emissions are binding at zero.
+This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al., (2017) and Cai and Lontzek, (2019).
+[^22] We also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
+Figure 4: Simulation results of regional SCC under the baseline emission cap scenario.
+The simulation results confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS.
+Figure 4 shows that when the emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 3.
+For example, in 2050, the MAC for the US is \$1,081 per ton of carbon, the permit price is \$845, and their difference is exactly equal to the regional SCC of the US, \$236.
+Figure 4 also shows that when the regional emissions are zero, the regional SCC is larger than the gap, which could be negative.
+Among all the regions, the US has the highest SCC in the near term, at around \$198 per ton of carbon in 2030.
+However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$682 for Africa and \$610 for non-OECD Asia by the end of the century.
+Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period.
+Our results show considerable heterogeneity in the SCC across regions and suggest that significant regional carbon tax implementation is required, not only for developed countries such as the US, but also for developing countries like Africa, India, and non-OECD Asia under the global ETS with the baseline emission cap scenario.
+6.3 Effects of ETS Implementation
+Next, we examine the economic and climate implications of the ETS implementation by comparing the global economy with and without the ETS regime, while maintaining the baseline emission caps in both economies.
+The top two panels of Figure 5 show that the ETS implementation results in slightly higher emissions and temperature increases compared to the case without the ETS.
+This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level, but the buyers do not have this option under no ETS.
+That is, the global emissions could be less than the global emission caps under no ETS, while they are equal under the ETS.
+This pattern persists until 2043.
+In the bottom two panels of Figure 5, we compare the MAC and the SCC using the US as an illustrative example.
+The MAC of the US economy is higher without the ETS until 2057, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement.
+The SCC comparison demonstrates that the ETS has little impact on the regional SCC.
+This pattern holds for the other regions as well, as shown in Figure A.10.
+Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results show that the ETS has little impact on the marginal damages and our endogenous social discount rates.
+Figure 5: Comparative analysis: effects of the ETS implementation.
+What are the welfare implications of implementing the global ETS?
+To quantify these effects, we compute a compensating variation (CV) per capita associated with the ETS implementation.
+Specifically, the CV per capita for region $i$ is computed by solving the following equation:
+$W_{i,0}(\mathbf{c}_{i}^{1}-CV)=W_{i,0}(\mathbf{c}_{i}^{0}),$ (31)
+where $\mathbf{c}_{i}^{1}=(c_{i,0}^{1},\cdots,c_{i,t}^{1},\cdots,c_{i,T}^{1})$ represents the vector of optimal consumption per capita under the ETS implementation, $\mathbf{c}_{i}^{0}=(c_{i,0}^{0},\cdots,c_{i,t}^{0},\cdots,c_{i,T}^{0})$ is the vector of optimal consumption per capita without the ETS implementation, and
+$W_{i,0}(\mathbf{c}_{i})=\sum_{t=0}^{T}\beta^{t}u(c_{i,t})L_{i,t}$
+is the social welfare associated with the vector of consumption per capita $\mathbf{c}_{i}=(c_{i,0},\cdots,c_{i,t},\cdots,c_{i,T})$ with the terminal time $T=299$.
+Note that a direct comparison of the noncooperative model with the ETS to the noncooperative model without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global emissions and resulting climate damages differ across the two economies.
+That is, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the emission caps and the ETS implementation.
+Compared with the economy without the ETS, temperature under the ETS implementation is expected to be higher as shown in Figure 5.
+Consequently, some regions may experience additional climate damage from the difference of temperatures between two cases, while it is greater than the reduced abatement cost from purchasing permits or the additional profit from selling permits under the ETS.
+This leads to negative welfare effects of the ETS implementation for some regions.
+To isolate the impact of the ETS, we adjust the emission caps $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which are the optimal emission levels obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS, then we compare economies with and without the ETS under the emission caps $\overline{E}^{\prime}_{i,t}$ as these two economies have the same global emission levels and temperatures.
+Table 1 shows the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$, when the ETS implementation does not impact temperature.
+The findings indicate that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions.
+For example, in the United States, the CV per capita is \$87.22, which is a relatively small fraction of per capita consumption (0.2%).
+Among the 12 regions, Russia experiences the largest welfare improvement, with a CV per capita of \$183.53, equivalent to 2.2% of its per capita consumption.
+In contrast, Africa and non-OECD Asia have the smallest welfare gains from the global ETS.
+These results highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
+Table 1: Welfare effects of the ETS implementation.
+US EU Japan Russia Eurasia China
+CV per capita (\$) 87.22 86.43 35.12 183.53 50.96 24.93
+CV (%) 0.201 0.344 0.124 2.225 1.017 0.250
+India MidEast Africa LatAm OHI OthAs
+CV per capita (\$) 37.09 54.96 13.06 56.93 69.33 8.15
+CV (%) 2.374 0.590 0.857 0.865 0.201 0.348
+7 Sensitivity Analysis
+We perform sensitivity analyses to evaluate alternative emission cap scenarios and to examine the influence of key parameters—climate damages and emission abatement—on our simulation results.
+7.1 Alternative Emission Cap Scenarios
+Along with the baseline emission cap scenario, we analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090, where all regions achieve net zero emissions by the specified year.
 The top-left and top-right panels in Figure 6 display the emission permit prices and expected temperature increases under different emission cap scenarios.
-Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,593 per ton of carbon in 2049, and the temperature rise is restricted to 1.61 degrees Celsius by the end of this century.
-Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$445 and \$261 per ton of carbon in 2049, respectively.
-In these scenarios, the temperature rise by the end of the century is expected to reach 1.8 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively.
-This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
-Figure 8: Simulation results of alternative emission cap scenarios under noncooperation with the ETS.
+Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,539 per ton of carbon in 2049, and the temperature rise is restricted to 1.62°C by the end of this century.
+Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$514 and \$335 per ton of carbon in 2049, respectively.
+In these scenarios, the temperature rise by the end of the century is expected to reach 1.80°C and 1.98°C above the pre-industrial level, respectively.
+This result shows that the global target of restricting the temperature rise to 1.5°C is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
+Figure 6: Sensitivity analysis: alternative emission cap scenarios.
 The bottom-left and bottom-right panels in Figure 6 show the regional MAC and SCC, taking the US as an example.
@@
 Our results show that stricter emission caps lead to higher MACs.
-Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2085 under the net zero 2090 scenario.
+Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,471 per ton of carbon in 2048, compared to the peak of \$1,322 in 2086 under the net zero 2090 scenario.
 This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC.
-We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$218 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$198 in 2100 in the net zero 2090 scenario.
-This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive).
+We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$226 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario.
+This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC (recall that the optimal carbon tax is equal to the gap between the MAC and permit price when emissions are strictly positive).
 Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
+7.2 Alternative Climate Damages and Abatement Costs
+Lastly, recognizing that climate damages and emission abatement costs are key drivers of our model outcomes, including permit prices, MAC, and the SCC, we conduct sensitivity analyses on the parameters for climate damages ($\pi_{1,i}$, $\pi_{2,i}$) and emission abatement ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$).
+While we incorporate climate damage projections from Burke et al., (2018) as our baseline model simulation, we additionally consider projections from Kahn et al., (2021) and Nordhaus, 2010a .
+Specifically, we calibrate the climate damage parameters to match the projections from Kahn et al., (2021) and directly adopt the parameters from Nordhaus, 2010a .
+Figure 7 demonstrates the key economic and climate outcomes under different climate damage assumptions.
+The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,059 by 2069, compared to \$2,282 under the damage parameters estimated from Kahn et al., (2021) and \$2,294 under the damage parameters from Nordhaus, 2010a .
+With the same emission cap constraints imposed, the global temperature outcomes remain unaffected despite variations in the damage parameters.
+For the US, the baseline climate damage estimation from Burke et al., (2018) leads to higher MAC and higher SCC, indicating that the baseline marginal climate damages are relatively higher than those projected by Kahn et al., (2021) and Nordhaus, 2010a .
+However, regional heterogeneity exists; for example, Russia experiences negative SCC under the baseline parameter values, while experiencing small but positive SCC under the parameter values estimated from Kahn et al., (2021) and Nordhaus, 2010a .
+See Appendix A.4.4 for a comparison across all 12 regions.
+Figure 7: Sensitivity analysis: alternative climate damage parameters.
+For our sensitivity analysis of the emission abatement cost parameters, we consider the parameter values from Nordhaus, 2010a in addition to the baseline parameter values calibrated from Ueckerdt et al., (2019), both of which share the same functional form of abatement cost.
+As shown in Figure 8, the permit price rises to \$1,767 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 85% of the permit price projected under the baseline scenario.
+Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70°C by the end of the century, due to the emission cap constraints.
+The lower emission permit price under the abatement cost estimate of Nordhaus, 2010a reflects lower MAC, as illustrated with the US case in Figure 8, with similar patterns observed across most regions (see Appendix A.4.5).
+Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$101 by 2100—just 32% of the baseline scenario.
+Figure 8: Sensitivity analysis: alternative abatement cost parameters
 8 Conclusion
 In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system.
-In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the Paris Agreement and their net zero commitments.
+In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact.
 We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting.
 The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth.
-We further show that carbon taxation can be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal net emissions.
-For strictly positive net emissions, we have shown both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
+We further show that carbon taxation needs to be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal emissions.
+For strictly positive emissions, we show both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
 This work has several policy implications.
-First, our results indicate that the current global target of restricting the global temperature rise to 1.5 degrees Celsius above pre-industrial levels by 2100 is unattainable under noncooperation, even with the current emission commitments outlined in the Paris Agreement.
+First, our results indicate that the current global target of restricting the global temperature rise to 1.5°C above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact.
 Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature.
-Second, our baseline simulation shows that the current emission commitments of the Paris Agreement lead to excess emission permit supply in the initial decade, resulting in permit prices of zero.
+Second, our baseline simulation shows that the current emission commitments lead to excess emission permit supply in the initial years, resulting in permit prices of zero.
 This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits.
@@
 A.1 List of Parameters
+Table A.1 lists the key parameters and their values.
 Table A.1: Key parameters.
-Parameter Description
-Economic parameters
-$\beta=0.985$ Discount factor
-$\gamma=1.45$ Elasticity of marginal utility
-$\alpha=0.3$ Output elasticity of capital
-$\delta=0.1$ Annual depreciation rate of capital
-Climate parameters
-$\zeta=0.0021$ Contribution rate of carbon emissions to temperature
-$\sigma_{i,t}$ Carbon intensity (GtC per trillion \$)
-$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$ Abatement cost parameters
-$\pi_{1,i}$, $\pi_{2,i}$ Climate damage parameters
-TFP parameters
-$g_{i,0},g_{i,t}$ Growth rate of TFP
-$d_{i}$ Parameters of TFP growth
+Parameter Value Description
+1) Economic system parameters (from Nordhaus, (2017))
+$\beta$ $0.985$ Annual discount factor
+$\gamma$ $1.45$ Elasticity of marginal utility
+$\alpha$ $0.3$ Output elasticity of capital
+$\delta$ $0.1$ Annual depreciation rate of capital
+2) Climate system parameters
+$\zeta$ $0.0021$ Contribution rate of carbon emissions to temperature
+Table A.2 lists the values of the baseline abatement cost parameters calibrated from Ueckerdt et al., (2019). The values of carbon intensity at annual time steps will be provided upon request.
+Table A.2: Abatement cost parameters (baseline) calibrated from Ueckerdt et al., (2019)
+US EU Japan Russia Eurasia China
+$b_{1,i}$ 0.462 0.477 0.750 0.292 0.347 0.328
+$b_{2,i}$ 2.859 2.670 2.011 2.499 3.243 2.822
+$b_{3,i}$ 9.920 5.832 2.492 7.625 7.966 7.189
+$b_{4,i}$ 0.182 0.114 0.2 0.2 0.168 0.168
+India MidEast Africa LatAm OHI OthAs
+$b_{1,i}$ 0.594 0.455 0.665 0.286 0.347 0.602
+$b_{2,i}$ 2.802 2.574 3.636 3.828 3.243 3.995
+$b_{3,i}$ 6.336 11.205 6.558 11.496 7.966 6.518
+$b_{4,i}$ 0.2 0.2 0.2 0.2 0.168 0.163
+Table A.3 lists the calibrated values of the climate damage parameters used in the baseline analysis.
+Table A.3: Climate damage parameters (baseline) calibrated from Burke et al., (2018)
+US EU Japan Russia Eurasia China
+$\pi_{1,i}$ 0.0842 0.0489 0.0090 -0.4169 0.2678 0.0003
+$\pi_{2,i}$ 0.0096 0.0011 0.0748 0.3094 0.0002 0.0008
+India MidEast Africa LatAm OHI OthAs
+$\pi_{1,i}$ 0.0017 0.3595 0.1886 0.1801 0.0123 0.2161
+$\pi_{2,i}$ 0.3276 0.0088 0.0764 0.0030 0.0044 0.0224
+Table A.4 lists the values of the TFP parameters calibrated from Burke et al., (2018).
+Table A.4: TFP parameters calibrated from Burke et al., (2018)
+US EU Japan Russia Eurasia China
+$g_{i,0}$ 0.0033 0.0089 0.0085 0.0170 0.0094 0.0345
+$d_{i}$ 0.0011 0.0010 0.0010 0.0154 0.0010 0.0308
+India MidEast Africa LatAm OHI OthAs
+$g_{i,0}$ 0.0332 0.0093 0.0218 0.0134 0.0076 0.0221
+$d_{i}$ 0.0151 0.0010 0.0013 0.0010 0.0010 0.0062
 A.2 List of Countries
-Table A.2: List of countries for regional aggregation.
+Table A.5: List of countries for regional aggregation
 Region Constituent Countries
-Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
-European Union [^17] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience.
-Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
-Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
-Latin America Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St.
-Kitts and Nevis, St.
-Lucia, St.
-Vincent and the Grenadines, Suriname, Trinidad and Tobago, Uruguay, Bolivarian Republic of Venezuela, and the U.S.
-Virgin Islands.
-Region Constituent Countries
-Middle East Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
-Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
-Other High-Income Australia, Canada, Republic of Korea, New Zealand, Singapore.
-A.3 Calibration of the TCRE Climate System
-Each of the four RCP scenarios (Meinshausen et al., 2011) — RCP 2.6, RCP 4.5, RCP 6, and RCP 8.5 — provide their pathways of emissions, atmospheric carbon concentration, radiative forcing, and atmospheric temperature anomaly.
+Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Ethiopia, Gabon, Gambia, The Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Libya, Madagascar, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Rwanda, Senegal, Sierra Leone, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, Zimbabwe.
+European Union [^23] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience..
+Austria, Belgium, Bulgaria, Croatia, Czech Republic, Denmark, Finland, France, Germany, Greece, Hungary, Ireland, Italy, Latvia, Lithuania, Luxembourg, Netherlands, Poland, Portugal, Romania, Slovak Republic, Spain, Sweden, United Kingdom.
+Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, FYR Macedonia, Moldova, Serbia, Slovenia, Tajikistan, Turkey, Ukraine, Uzbekistan.
+Latin America Argentina, Bahamas, The Belize, Bolivia, Brazil, Chile, Colombia, Costa Rica, Cuba, Dominican Republic, Ecuador, El Salvador, Guatemala, Haiti, Honduras, Mexico, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, Trinidad and Tobago, Uruguay.
+Middle East Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Lebanon, Oman, Qatar, Saudi Arabia, United Arab Emirates
+Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, Indonesia, Malaysia, Mongolia, Nepal, Pakistan, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Vanuatu, Vietnam.
+Other High-Income Australia, Canada, Iceland, Republic of Korea, New Zealand, Norway, Switzerland.
+A.3 Details about Calibration and Data
+A.3.1 Calibration of the TCRE Climate System
+Each of the four RCP scenarios (Meinshausen et al.,, 2011) — RCP 2.6, RCP 4.5, RCP 6, and RCP 8.5 — provide their pathways of emissions, atmospheric carbon concentration, radiative forcing, and atmospheric temperature anomaly.
 When we calibrate the contribution rate of carbon emissions on temperature, $\zeta$, in a climate system, we use the pathways of emissions and atmospheric temperature anomaly of the four RCP scenarios.
@@
 Figure A.1: Calibration of the TCRE climate system.
-A.4 Emission Cap Pathways
-Figure A.2 shows the regional emission cap pathways generated using the methodology described in Section 5.1. Figure A.3 compares the different emission cap pathways at the global level.
+A.3.2 Calibration of Total Factor Productivity and Climate Damage
+Figures A.2 and A.3 show that with our calibrated TFP and climate damage coefficients, the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ or $y_{i,t}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD,NoCC}}$ or $y_{i,t}^{\mathrm{BDD}}$ from Burke et al., (2018), respectively, for all regions.
+Figure A.2: Fitting GDP per capita under no climate impact. Lines represent GDP per capita under no climate impact from Burke et al., (2018); marks represent fitted GDP per capita.
+Figure A.3: Fitting GDP under climate impact. Lines represent the GDP per capita under the climate impact of RCP 4.5 from Burke et al., (2018); marks represent fitted values.
+A.3.3 Calibration of Climate Damage from Kahn et al., (2021)
+For sensitivity analysis on climate damage parameters in Section 7.2, we calibrate the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$ by considering projections on GDP loss across different climate scenarios in Kahn et al., (2021), which shows the percentage loss in GDP per capita by 2030, 2050, and 2100 under the RCP 2.6 and RCP 8.5 scenarios for China, EU, India, Russia, and the US.
+We use their method and data to project the percentage loss in GDP per capita ($\Delta_{i,t}^{\mathrm{RCP26}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}$) every year from 2020 to 2114 under the RCP 2.6 and RCP 8.5 scenarios for each of our 12 regions, employing the baseline setup in Kahn et al., (2021).
+Specifically, $\Delta_{i,t}^{\mathrm{RCP26}}=1-y_{i,t}^{\mathrm{RCP26}}/y_{i,t}^{\mathrm{base}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}=1-y_{i,t}^{\mathrm{RCP85}}/y_{i,t}^{\mathrm{base}}$, where $y_{i,t}^{\mathrm{RCP26}}$, $y_{i,t}^{\mathrm{RCP85}}$, and $y_{i,t}^{\mathrm{base}}$ are GDP per capita under RCP 2.6, RCP 8.5, and the baseline scenario, respectively.
+Thus from equation (17) we obtain $(\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem for each region $i$:
+$\min_{\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{94}\left(\frac{1+\pi_{1,i}T_{t}^{\mathrm{RCP85}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP85}}\right)^{2}}{1+\pi_{1,i}T_{t}^{\mathrm{RCP26}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP26}}\right)^{2}}-\frac{1-\Delta_{t,i}^{\mathrm{RCP26}}}{1-\Delta_{t,i}^{\mathrm{RCP85}}}\right)^{2}.$ (A.1)
+Here $T_{t}^{\mathrm{RCP26}}$ and $T_{t}^{\mathrm{RCP85}}$ are the global average temperature anomalies at time $t$ (deviation from the pre-industrial temperature) under the RCP 2.6 and RCP 8.5 scenarios.
+Figure A.3 shows, with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al., (2021).
+Figure A.4: Fitting climate damage parameters. Lines represent the ratios of GDP per capita between RCP 2.6 and RCP8.5 from Kahn et al., (2021); marks represent fitted ratios.
+A.3.4 Regional Emission Cap Pathways
+Figure A.5 displays the regional emission cap pathways, measured in Gigatonne of Carbon (GtC), for the baseline emission cap scenario, generated using the methodology described in Section 5.1.
 Figure A.5: Regional emission cap pathways under the baseline scenario.
+Table A.6 lists the regional emission caps for every region in five-year time steps. The regional emission caps at annual time steps will be provided upon request.
+Table A.6: Regional emission cap pathways under the baseline scenario (unit: GtC)
+2020 2025 2030 2035 2040 2045 2050 2055 2060 2065 2070
+US 1.603 1.330 0.812 0.609 0.406 0.203 0.000 0.000 0.000 0.000 0.000
+EU 1.081 0.923 0.617 0.463 0.309 0.154 0.000 0.000 0.000 0.000 0.000
+Japan 0.308 0.260 0.197 0.148 0.099 0.049 0.000 0.000 0.000 0.000 0.000
+Russia 0.694 0.674 0.621 0.517 0.414 0.310 0.207 0.103 0.000 0.000 0.000
+Eurasia 0.464 0.450 0.390 0.324 0.259 0.193 0.126 0.071 0.030 0.015 0.000
+China 3.433 3.370 3.061 2.551 2.042 1.532 1.023 0.513 0.004 0.002 0.000
+India 0.940 0.773 0.295 0.259 0.222 0.185 0.148 0.111 0.074 0.037 0.000
+MidEast 0.638 0.638 0.530 0.446 0.363 0.280 0.196 0.140 0.083 0.041 0.000
+Africa 0.743 0.712 0.621 0.526 0.434 0.342 0.251 0.188 0.125 0.063 0.000
+LatAm 0.815 0.736 0.598 0.480 0.362 0.244 0.126 0.094 0.063 0.031 0.000
+OHI 0.613 0.538 0.347 0.260 0.173 0.087 0.000 0.000 0.000 0.000 0.000
+OthAs 0.924 0.882 0.630 0.535 0.440 0.345 0.249 0.172 0.094 0.040 0.000
+Figure A.6 presents a comparison of global emission cap pathways under various scenarios: the baseline scenario, net-zero by 2050, net-zero by 2070, and net-zero by 2090.
+In the net-zero scenarios, it is assumed that all countries achieve net-zero emissions by the respective target years.
 Figure A.6: Global emission cap pathways for the different net zero scenarios.
-A.5 GDP Growth Rate beyond this Century
+A.3.5 GDP Growth Rate beyond this Century
 For the GDP growth rate beyond this century, we follow RICE to project $g_{t,i}$ for $t\geq 80$.
@@
 We have $\ln\left(\frac{\widetilde{y}_{i,t+1}}{\widetilde{y}_{i,t}}\right)=g_{i,t}+\alpha\ln\left(\frac{k_{i,t+1}}{k_{i,t}}\right).$ If we assume the growth of $k_{t,i}$ is equal to the growth of GDP per capita, then we have $\widetilde{y}_{i,t+1}=\widetilde{y}_{i,t}\exp(g_{i,t}/(1-\alpha)).$
-A.6 Calibration of Total Factor Productivity
-Figure A.4 shows that with our calibrated $A_{i,t}$ , the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD}}$ from Burke et al.
-(2018) for all regions.
-Figure A.4: Fitting GDP per capita under no climate impact. Lines represent GDP per capita from Burke et al. (2018); marks represent fitted GDP per capita.
-A.7 Calibration of Climate Damage
-Figure A.5 shows that with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al.
-(2021), for every region.
-Figure A.5: Fitting climate damage parameters. Lines represent the ratios of GDP per capita between RCP 2.6 and RCP8.5 from Kahn et al. (2021); marks represent fitted ratios.
-A.8 Additional Simulation Results
-A.8.1 Regional Net Emissions under Noncooperation
-Figure A.6 displays the regional net emissions under the noncooperative model with the ETS and the baseline emission cap scenario.
-Latin America is the first to reach net emissions in 2053, followed by Russia in 2055, the US in 2056, OHI in 2057, and Eurasia in 2058.
-Then, net emissions are achieved by China in 2061, Japan in 2062, the MidWest in 2064, and the EU in 2065.
-Finally, India achieves net emissions in 2069, followed by Africa and other non-OECD Asian countries in 2071.
-Figure A.6: Simulation results of regional net emissions under noncooperation with the ETS and baseline emission caps.
-A.8.2 Regional Net Emissions Comparison between Different Cases
-Figure A.7 compares regional net emissions between two cases under cooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
-Figure A.7: Simulation results of regional net emissions: comparison between two cases under cooperation.
-Figure A.8 compares regional net emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
-Figure A.8: Simulation results of regional net emissions: comparison between two cases under noncooperation.
-A.8.3 Regional MAC Comparison between Different Cases
-Figure A.9 compares the regional MAC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.9: Simulation results of regional MAC: comparison between two cases under cooperation.
+A.4 Additional Simulation Results
+A.4.1 Benchmark Model: Regional Emissions
+Figure A.7 displays the regional emissions under the noncooperative model with the ETS and the baseline emission cap scenario.
+Russia is the first to reach net zero emissions in 2050, followed by China and Latin America in 2056, MidEast in 2057, the US and Eurasia in 2058, and the OHI in 2059.
+Then, net zero emissions are achieved by EU in 2061, Japan and India in 2064.
+Finally, Africa and non-OECD Asia achieve net zero emissions in 2070.
+Figure A.7: Simulation results of regional emissions under the baseline emission caps.
+A.4.2 Model Comparison of ETS Implementation
+Figure A.8 compares regional emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
+Figure A.8: Comparison of the ETS implementation: regional emissions.
 Figure A.9 compares the regional MAC under noncooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.10: Simulation results of regional MAC: comparison between two cases under noncooperation.
-A.8.4 Regional SCC Comparison between Different Cases
-Figure A.11 compares the regional SCC under cooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.11: Simulation results of the regional SCC: comparison between two cases under cooperation.
+Figure A.9: Comparison of the ETS implementation: regional MAC.
 Figure A.10 compares the regional SCC under noncooperation with the baseline emission caps, comparing two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.12: Simulation results of the regional SCC: comparison between two cases under noncooperation.
-A.8.5 Regional MAC and SCC under Alternative Emission Caps
+Figure A.10: Comparison of the ETS implementation: regional SCC.
+A.4.3 Sensitivity Analysis over Emission Caps
 In Figure A.11, we compare the MAC for the noncooperative model with the ETS across alternative emission cap scenarios.
-Figure A.13: Simulation results of regional MAC: comparison across alternative emission cap scenarios.
+Figure A.11: Comparison of regional MAC under different emission caps.
 Similarly, Figure A.12 displays the SCC of the noncooperative model with the ETS across different emission cap scenarios.
-Figure A.14: Simulation results of regional SCC: comparison across alternative emission cap scenarios.
-A.8.6 Sensitivity Analysis on Marginal Abatement Cost
-In Figure A.15, we provide additional sensitivity analysis results on the abatement cost. Recall that, the MAC in the model is
-$\mathrm{MAC}{}_{i,t}=1,000b_{2,i}\mu_{i,t}^{b_{2,i}-1}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t)).$
-We investigate a scenario wherein the MACs across all regions experience a variation of ±20% from the baseline MAC levels at the same emission control rates $\mu_{i,t}$.
-Specifically, we explore the effects of adjusting two key parameters $b_{1,i}$ and $b_{3,i}$, by ±20% from their baseline values.
-The result shows that the equilibrium permit price is closely linked to the level of the MACs.
-When the MACs are elevated across all regions, the permit price also experiences a corresponding increase with almost identical changes of the MACs, and vise versa.
-For example, in 2040, the baseline MAC in the US is \$556, it is changed to \$644 or \$467, and the permit price is changed from \$404 to \$492 or \$315.
-That is, both the MACs and the permit price increase \$88 or decreases \$89 in 2040.
-At the same time, the regional SCC has little changes, as the existence of the emission caps makes the future temperature have little changes and the changes in the MACs have little impact on marginal climate damages in future periods and social discount rates.
-This again verifies our relation between the optimal carbon tax, the MAC, and the permit price (i.e., equation (20)).
-Figure A.15: Sensitivity Analysis on the MAC.
+Figure A.12: Comparison of regional SCC under different.
+A.4.4 Sensitivity Analysis over Climate Damage Parameters
+In Figure A.13, we compare the MAC for the noncooperative model with the ETS under alternative values of the climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on climate damage projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline climate damage projection from Burke et al., (2018).
+Figure A.13: Comparison of regional MAC under different estimates of climate damages.
+Similarly, Figure A.14 displays the SCC of the noncooperative model with the ETS under different values of the climate damage parameters.
+Figure A.14: Comparison of regional SCC under different estimates of climate damages.
+A.4.5 Sensitivity Analysis over Abatement Cost Parameters
+In Figure A.15, we compare the MAC for the noncooperative model with the ETS under alternative estimates of the emissions abatement parameters ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$), calibrated from Ueckerdt et al., (2019) and Nordhaus, 2010a .
+Figure A.15: Comparison of regional MAC under different estimates of emissions abatement cost.
+Similarly, Figure A.16 displays the SCC of the noncooperative model with the ETS under different emissions abatement cost estimates.
+Figure A.16: Comparison of of regional SCC under different estimates of emissions abatement cost.
+A.4.6 Sensitivity over TFP Growth Rates
+Figure A.17 compares key model outcomes under different TFP growth rates: the baseline TFP growth rates derived from Burke et al., (2018) and the alternative rates based on Nordhaus, 2010a .
+Figure A.17: Comparison of simulation results under different TFP growth rates.

### v4

@@
 Abstract
-We build a dynamic multi-region model of climate and economy with emission permit trading among 12 aggregated regions in the world.
-We solve for the dynamic Nash equilibrium under noncooperation, wherein each region adheres to the emission cap constraints following commitments that were first outlined in the 2015 Paris Agreement and later strengthened under the Glasgow Pact.
-Our model shows that the emission permit price reaches \$845 per ton of carbon by 2050, and global average temperature is expected to reach 1.7°C above the pre-industrial level by the end of this century.
-We demonstrate, both theoretically and numerically, that a regional carbon tax is complementary to the global cap-and-trade system, and the optimal regional carbon tax is equal to the difference between the regional marginal abatement cost and the permit price.
-We show that optimal regional carbon tax has significant heterogeneity between regions, and the tax needs to be implemented in both the developed and the developing regions.
+We develop a dynamic multi-region climate–economy model with emissions trading and solve for the dynamic Nash equilibrium under noncooperation, where each region follows Paris Agreement–based emissions caps.
+The permit price reaches \$923 per ton of carbon by 2050, and global temperature rises to 1.7°C above pre-industrial levels by 2100.
+The regional social cost of carbon equals the difference between regional marginal abatement cost and the permit price, highlighting complementarity between carbon taxes and trading.
+We find substantial heterogeneity in regional social costs of carbon, show that lax caps can raise emissions, and demonstrate strong free-rider incentives under partial participation.
 Keywords: Emission trading system, Paris Agreement, dynamic Nash equilibrium, integrated assessment model, carbon tax, social cost of carbon.
@@
 1 Introduction
-Climate change has been raising concerns of disastrous consequences, from rising sea levels to increase in the frequency and intensity of extreme weather events (Arias et al.,, 2021).
+Climate change has raised concerns of disastrous consequences, ranging from rising sea levels to increase in the frequency and intensity of extreme weather events (Arias et al.,, 2021).
 International efforts to address climate change have evolved over the past three decades, commencing with the United Nations Framework Convention on Climate Change (UNFCCC) in 1992, followed by the Kyoto Protocol in 1997, the Paris Agreement in 2015, and the Glasgow Pact in 2021.
-The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2°C above pre-industrial levels, with a pursuit of limiting it to the 1.5°C” (Article 2).
-To accomplish this objective, 194 countries (regions) committed to the Nationally Determined Contributions (NDCs) specifying their emission mitigation goals, which need to be updated every five years by strengthening emission cuts with more intensive measures.
-Most recently, the Glasgow Climate Pact, adopted at the UN Climate Change Conference in Glasgow (COP26) in November 2021, revisited Nationally Determined Contributions (NDCs), reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more stringent efforts to attain the global target of 1.5°C.
-Various market-based approaches to reduce emissions have been proposed, among which the most well-known are carbon tax regimes and emission trading systems (ETSs).
+The policy outlined in the Paris Agreement codified an objective of “limiting global average temperature rise to well below 2 degrees Celsius above pre-industrial levels, with a pursuit of limiting it to 1.5 degrees Celsius” (Article 2).
+To accomplish this objective, 194 countries (regions) committed to the Nationally Determined Contributions (NDCs), which specify their emission mitigation goals.
+The Glasgow Climate Pact, adopted at the UN Climate Change Conference in Glasgow (COP26) in November 2021, revisited the NDCs, reaffirming the previous commitments made in the Paris Agreement and recognizing the need for more stringent efforts to attain the global target of 1.5 degrees Celsius.
+Various market-based approaches have been proposed to reduce emissions, among which carbon tax regimes and emission trading systems (ETSs) are the most prominent.
 [^1] It has been documented that a total of 61 carbon pricing policies, consisting of 30 taxes and 31 ETSs, have been executed or are scheduled for implementation globally (Stavins,, 2022).
 In a carbon tax regime, a tax is charged to firms for each unit of carbon emission.
-Previous studies have shown that the optimal global carbon tax rate is equal to the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of carbon emissions (Nordhaus,, 2017).
-In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of permissible emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at the market price.
-In the absence of transaction costs and uncertainty, properly designed tax and trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery,, 1972; Goulder and Schein,, 2013; Stavins,, 2022).
+Previous studies have argued that the optimal global carbon tax rate is equal to the global social cost of carbon (SCC), which is the present value of global climate damages incurred by an additional unit of atmospheric carbon emissions (Nordhaus,, 2017).
+In an emission trading or a cap-and-trade system, on the other hand, the maximum amount of permissible emissions is fixed in an economy, and agents can sell and purchase emission permits (allowances) at market-determined prices.
+In the absence of transaction costs and uncertainty, properly designed carbon tax and emission trading regimes are argued to yield equivalent outcomes in several key aspects (e.g., incentives for emission reductions, aggregate abatement costs, and carbon leakage) (Montgomery,, 1972; Goulder and Schein,, 2013; Stavins,, 2022).
 In 2005, the European Union (EU) first adopted a legally binding emission trading system among the EU members, [^2] In addition to the EU ETS, other national or sub-national ETSs have been implemented or are in the process of development, including in Canada, China, Japan, New Zealand, South Korea, Switzerland, and the United States.
@@
 [^3] Source of EU permit price: (https://tradingeconomics.com/commodity/carbon).
-Although a global-level ETS is currently not in place, there have been longstanding discussions on its implementation as an international regime to address climate change, as implied by Article 17 of the Kyoto Protocol and Article 6 of the Paris Agreement.
-Prior works have emphasized the need for a global ETS or linking existing regional ETSs (Mehling et al.,, 2018) and for expanding the coverage of carbon trading in specific industries to different regions so as to enhance the global mitigation efforts and reduce carbon leakages (Piris-Cabezas et al.,, 2018).
-In this study, we develop a dynamic multi-region model of climate and the economy under a global ETS with annual time steps, and use it to project global prices of emission permits and the regional SCC.
-In this integrated assessment framework, we form 12 aggregated regions by grouping 152 countries around the world.
-For these regions, we construct emission cap trajectories aligned with the latest NDC mitigation targets from the Paris Agreement and the Glasgow Climate Pact.
-These trajectories are based on near-term targets for 2025/2030 and long-term net-zero commitments for 2050–2070, and our model treats these trajectories as exogenous emission cap constraints.
-Assuming that all regions adhere strictly to their emission cap constraints without future revisions, [^4] In this paper, we assume that regions will not revise their NDC limits in future for the following reasons: (i) relaxing NDC limits could result in penalties from other regions or damage to the region’s reputation; (ii) reducing NDC limits would lead to a loss of benefits from selling emission permits or more cost from purchasing; and (iii) renegotiating NDC limits is often time-consuming and infrequent.
-our main analysis focuses on solving a dynamic Nash equilibrium under noncooperation, where each region makes optimal dynamic decisions on emission abatement, emission permit purchases, and capital investment to maximize its social welfare given the optimal choices of other regions.
+In this study, we develop and quantify a dynamic multi-region model of the climate and economy under an emission trading system (ETS), while the ETS could be global or partial.
+The model framework extends the seminal Regional Integrated Model of Climate and the Economy (RICE) (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), which captures the interactions between economic growth and climate systems in a multi-regional framework.
+As in the standard neoclassical growth model (the so-called Ramsey model), a social planner of each region chooses investment in capital goods to smoothen consumption over time.
+Economic activity generates both output and carbon emissions; the latter induces climate damages that endogenously reduce regional output.
+Emission abatement efforts can reduce carbon emissions but are costly.
+As emissions generate global externalities, each region has less incentive to undertake abatement on its own while benefiting from the abatement efforts of others.
+In a noncooperative environment, a set of the optimal strategies of each regional social planner, involving dynamic choices over consumption and emission abatement, is characterized as a Nash equilibrium concept from game theory (Nordhaus and Yang,, 1996).
+The introduction of the ETS imposes exogenous region-specific emissions caps and allows permit trading across regions.
+Under the ETS, regional social planners jointly determine the optimal paths of consumption, emission abatement, and the amount of permit trading, subject to strictly-enforced emission cap constraints.
+Heterogeneity across regions, in abatement technologies, climate damages, productivity, carbon intensity, and emissions cap constraints, causes divergence in regional marginal abatement costs (MAC) and the regional social cost of carbon across regions, shaping the path of the market equilibrium permit price over time.
+In an ETS scheme, the primary forces that shape the equilibrium outcome include regional climate damages, revenues or costs from permit trading, and regional abatement costs.
+Using a simple static framework where a regional social planner $i$ internalizes climate damages and chooses optimal net emissions, $E_{i}$, (i.e., emissions after abatement), we show that the equilibrium condition for optimal abatement can be expressed as: $\mathrm{SCC}=\mathrm{MAC}(E_{i})-m$, where $m$ is the market price of emission permits.
+This relationship between regional SCC, MAC, and emission price has an important implication: a positive regional SCC implies $\mathrm{MAC}(E_{i})>m$, suggesting that a region experiencing climate damages has an incentive to abate emissions up to a level where the MAC exceeds the market price of emission permits.
+We emphasize that this equilibrium condition represents the necessary condition for achieving the optimal level of net emissions from the perspective of each regional social planner.
+In the literature, the equality between $\mathrm{MAC}$ and permit price is used as a competitive equilibrium condition under an ETS regime in a decentralized economy, where individual firms internalize the emission price.
+This condition is derived from the perspective of individual firms in the ETS when they face no additional carbon tax.
+If a carbon tax $\tau_{i}$ is levied on firms on top of the permit price, the competitive equilibrium condition becomes $\tau_{i}=\mathrm{MAC}(E_{i}^{\prime})-m^{\prime}$ in the decentralized economy, where $E_{i}^{\prime}$ and $m^{\prime}$ are the corresponding net emissions and permit price under the decentralized equilibrium with tax $\tau_{i}$.
+If the carbon tax is set equal to the regional SCC in each region, then $E_{i}^{\prime}$ and $m^{\prime}$ are equal to the regional social planner’s optimal net emissions and permit price.
+Moreover, because climate damages—and hence the regional SCC—differ across regions, a uniform global carbon price alone may not achieve regionally optimal level of emission pricing.
+The appropriate equilibrium condition in this context, $\mathrm{SCC}=\mathrm{MAC}(E_{i})-m$, further implies that carbon taxation and cap-and-trade are not competing instruments but rather complementary.
+We bring the model to data by aggregating 152 countries into 12 world regions and simulating the model at annual time steps.
+The model is calibrated by fitting it to historical data and the recent predicted trends of regional climate damage (Burke et al.,, 2018), regional total factor productivity (Burke et al.,, 2018), regional abatement costs (Ueckerdt et al.,, 2019), and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz,, 2017), also known as the “Middle-of-the-Road” scenario.
+Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al.,, 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on the cumulative carbon emissions.
+We demonstrate that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al.,, 2011).
+Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas, (2017), Dietz and Venmans, (2019), Mattauch et al., (2020), and Barnett et al., (2020).
+Additionally, Dietz et al., (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus,, 2017), compared to other more complicated climate systems.
+For regional emission cap scenarios, we construct emission cap pathways reflecting the latest emission targets for 2030 and net-zero pledges for 2050-2070, contributing to the scant literature on evaluating the economic and environmental implications of these commitments (e.g., van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022).
+For each region, emissions cap trajectories are constructed to align with the latest nationally determined contributions (NDCs) under the Paris Agreement and the Glasgow Climate Pact.
+These trajectories reflect near-term targets for 2025 and 2030, as well as long-term net-zero commitments spanning 2050 to 2070, and are treated as exogenous constraints.
+Our numerical simulations, based on newly calibrated parameters, provide comprehensive results on the potential economic and environmental outcomes under the ETS.
+Assuming full compliance with these emission caps and no future revisions, [^4] This paper assumes that regions do not revise their NDC limits (emission cap constraints) in the future.
+While primarily introduced to address computational challenges, this assumption is also supported by economic and institutional considerations: (i) relaxing NDC limits could result in penalties from other regions or damage to the region’s reputation; (ii) reducing NDC limits would lead to a loss of benefits from selling emission permits or more cost from purchasing; and (iii) renegotiating NDC limits is often time-consuming and infrequent.
 the emission permit price is endogenously determined by annual supply and demand in the global permit market.
-Our study makes four contributions.
-The first contribution is demonstrating, both theoretically and numerically, that regional optimal taxes can be internally implemented under the global ETS.
-While a carbon tax and ETS have often been considered as potentially interchangeable or substitutable policies, we show that within the global ETS framework, a hybrid policy combining both ETS and carbon tax is optimal: the social planner of each region needs to introduce a carbon tax as a complementary policy tool under the global ETS in order to achieve its optimal level of emissions.
-The necessity of a carbon tax arises from the discrepancy in optimal emissions between the social planner’s choice and the market outcome under an ETS-only regime (i.e., firms’ choice under the ETS with an exogenous emission cap but no carbon tax).
-Specifically, when optimal emissions from the social planner’s perspective are strictly positive, the regional optimal carbon tax is equal to the difference between the regional marginal abatement cost (MAC) and the global permit price under the noncooperative Nash equilibrium with the ETS, thereby ensuring firms’ choice under the tax is consistent with that of the social planner’s.
-We also numerically verify that the regional optimal carbon tax is equal to the regional SCC provided emissions are strictly positive.
-Regional SCC is the present value of future regional climate damages resulting from additional global emissions.
-The second contribution of this study is a method to solve for the dynamic Nash equilibrium prices of emission permits under a global ETS.
-To the best of our knowledge, this is the first attempt to integrate an emission permit trading system with fully endogenous prices and emission abatements in a dynamic regional integrated assessment model (IAM) under noncooperation.
-Although some prior research has addressed regional IAMs with a multi-region ETS, their solutions thus far have been obtained in a limited manner.
-One seminal work in IAMs is the Regional Integrated Model of Climate and the Economy (RICE; Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), but it did not provide a solution for the market-clearing equilibrium permit prices under noncooperation.
-Other studies have explored an ETS by extending the World Induced Technical Change Hybrid Model framework (WITCH; Bosetti et al.,, 2006), yet their permit prices and emission abatements are not fully endogenized as their prices are assumed to be equal to their regional MACs (De Cian and Tavoni,, 2012; Massetti and Tavoni,, 2012).
-While it has been long considered in the literature that the market equilibrium permit price is equalized to the regional MACs across all countries in an ETS regime (e.g., Keohane,, 2009; Nordhaus, 2010b, ), our study shows that this does not hold.
-Specifically, the assumption that the permit price equals the regional MACs ignores the inconsistency between regional social planners’ choices internalizing the regional climate damages and firms’ choices ignoring the regional climate damages under the ETS regime.
-This can also be understood with an extreme case: if the emission caps are very large such that they are not binding for any region (which is equivalent to a case without emission caps), then both permit prices and trading quantities are zero, implying zero emission mitigation.
-However, it is well known that even without emission caps, there are still nonzero emission mitigations from the regional social planners’ perspectives and the optimal regional taxes are equal to the nonzero regional MACs under the Nash equilibrium (see, e.g., Nordhaus and Yang,, 1996; Cai et al.,, 2023).
-That is, regions’ social planners should impose an additional carbon tax on firms, contingent on their NDC commitments, such that the firms’ emission decisions are consistent with theirs.
-We present a numerical algorithm for finding a Nash equilibrium solution where forward-looking regional social planners are making dynamic optimal decisions, and importantly, we show that the market clearing permit price is not equal to the regional MACs and the regional MACs are not equalized across regions under the Nash equilibrium.
-Under the baseline emission cap scenario based on the Paris Agreement and Glasgow Pact commitments, our simulation results show that the oversupply of global permits in the initial years results in zero permit prices, but by 2050, the emission permit price can reach up to \$845 per ton of carbon.
-The global average temperature is expected to reach 1.7 degree Celcius above the pre-industrial level by the end of this century.
-The third contribution is that our model’s results are based on newly calibrated parameters and regional emission cap pathways under the Paris Agreement and the Glasgow Pact, making our findings directly relevant to real-world scenarios.
-We calibrate our model by fitting it to historical data and the recent predicted trends of regional climate damage, regional GDP, regional abatement costs, and population projections based on the Shared Socioeconomic Pathway 2 (SSP2; Samir and Lutz,, 2017), also known as the “Middle-of-the-Road” scenario.
-Besides this calibration, our model uses a stylized but stable climate system, called the Transient Climate Response to Emissions (TCRE; Matthews et al.,, 2009), which assumes that increases in the global averaged atmospheric temperature have a nearly linear functional dependence on the cumulative carbon emissions.
-We show that this temperature system can be calibrated to match closely with the various Representative Concentration Pathways (RCPs; Meinshausen et al.,, 2011).
-Due to its simplicity and effectiveness, the TCRE scheme has found applications in recent economic analyses, as evidenced by studies such as Brock and Xepapadeas,, 2017; Dietz and Venmans,, 2019; Mattauch et al.,, 2020; Barnett et al.,, 2020.
-Moreover, Dietz et al., (2021) show that the TCRE scheme does not lead to a large difference in economic analysis with the seminal DICE framework (Nordhaus,, 2017), compared to other more complicated climate systems.
-Furthermore, our construction of emission cap pathways reflects the latest emission targets for 2030 and net-zero pledges for 2050-2070, contributing to the scant literature on evaluating the economic and environmental implications of these commitments (e.g., van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022).
-Our final contribution is a comprehensive comparative analysis of the roles of emission caps and the global ETS.
-We find that the global ETS leads to higher emissions under noncooperation, because regions with binding emission caps can purchase permits from regions with less restrictive caps, exploiting the excessively abundant emission permits in the initial years.
-This implies that it is important to maintain a stringent emission cap at the global level to make the ETS work efficiently.
-Moreover, we show that a stricter emission cap leads to a higher permit price for every region.
-For example, if all the regions achieve net zero emissions by 2050, the emission permit price can reach up to \$1,539 per ton of carbon by 2049.
-However, even with this strict emission restriction, the global temperature is still projected to rise by 1.62°C by the end of this century.
-These findings suggest that achieving the global target of limiting the temperature rise to 1.5°C requires even stronger emission commitments than those outlined in the Glasgow Pact.
-Our results are in qualitative agreement with the analysis of other IAMs (van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022), which find that the strengthened post-Glasgow NDC targets, encompassing both near-term and long-term goals, are insufficient to limit warming to 1.5°C above pre-industrial levels, but, if fully implemented, could limit warming to below 2°C.
-To ensure that our results are robust, we have performed sensitivity analysis by choosing alternative climate damage and abatement cost parameters obtained by fitting these parameters to other literature data.
-Our model framework is closely related to the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ), an extension of the global DICE model (Nordhaus,, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems.
-In Nordhaus, 2010b , carbon prices are obtained under a cooperative model where the global welfare is maximized with weights applied on regional utilities.
-Following the RICE model, a group of studies have explored climate policy in a dynamic multi-region framework under cooperation and noncooperation.
-For instance, Luderer et al., (2012) and Jakob et al., (2012) compare the long-term predictions of three-region energy-economy models, when certain environmental goals are set, like stabilizing the CO2 levels to 450 ppm.
-However, these works do not incorporate actual NDC and net-zero year targets of the countries to study the region-wise emission-economy pathways.
-van der Ploeg and de Zeeuw, (2016) and Jaakkola and Van der Ploeg, (2019) focus on stochastic components, such as climate tipping points or technological breakthroughs, comparing the results under different levels of cooperation.
-Hambel et al., (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs, under certain model assumptions.
+Under the baseline emission cap scenario based on the Paris Agreement and Glasgow Pact commitments, our simulation results show that the oversupply of global permits in the initial years results in zero permit prices, but by 2050, the emission permit price can reach up to \$923 per ton of carbon.
+The corresponding global average temperature is expected to reach 1.7 degree Celsius above the pre-industrial level by the end of this century.
+Furthermore, our simulation results numerically confirm that the regional SCC exactly equals the difference between the regional MAC and the permit price, as explored in the predictions of the static framework.
+For instance, in 2050, the MAC for the United States is estimated at \$1,159 per ton of carbon, the permit price at \$923, and the resulting difference of \$236 matches the regional SCC calculated independently from the model.
+A comparative analysis of the roles of emission caps and the ETS further illustrates several key findings.
+First, the findings indicate that the ETS with the baseline emission caps leads to higher emissions under noncooperation, as regions with binding emission caps can purchase permits from less constrained regions, exploiting the surplus permits in the initial years.
+This highlights the necessity of maintaining stringent global emission caps to ensure the efficient functioning of the ETS.
+Second, our welfare analysis reveals that when the global emission cap is sufficiently tight, the ETS can lead to welfare improvements for all participating regions.
+These welfare improvements reflect efficiency gains achieved by reallocating abatement efforts across regions through market mechanisms.
+Additionally, leveraging the flexibility of our model, we evaluate two additional set of policy scenarios, including a partial ETS participation and alternative emission caps consistent with different net-zero targets.
+Recognizing the challenges of achieving international cooperation on climate policies, we consider a scenario in which one major region (the United States) opts out of the global cap-and-trade system.
+The results show that the United States experiences a notable welfare gain.
+This outcome underscores strong free-rider incentives, as the opt-out region faces a lighter abatement burden while benefiting from the emission reductions undertaken by participating regions.
+In a second policy experiment, we find that tightening the global emission cap leads to higher permit prices across future periods.
+For instance, if all regions achieve net-zero emissions by 2050, the permit price could reach \$1,621 per ton of carbon by 2049.
+However, even with such strict emissions restrictions, the global temperature is projected to rise by 1.62°C by the end of the century.
+These results suggest that meeting the global target of limiting warming to 1.5°C will require even more stringent emission reduction commitments than those set out in the Glasgow Pact, albeit at the cost of significantly higher permit prices.
+Our results are qualitatively consistent with the findings of other IAMs (van de Ven et al.,, 2023; Meinshausen et al.,, 2022; den Elzen et al.,, 2022), which indicate that the strengthened post-Glasgow NDC and net-zero targets, covering both near-term and long-term goals, are insufficient to limit warming to 1.5°C above the pre-industrial level, though full implementation could restrict warming to below 2°C.
+Finally, to ensure the robustness of our results, we conduct sensitivity analyses using alternative parameter values for climate damages and abatement costs, calibrated to empirical estimates reported in the key existing literature.
+2 Related Literature
+This paper contributes to the literature on macroeconomic modeling of climate change.
+Our model framework is closely related to the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ; Yang,, 2023), an extension of the global DICE model (Nordhaus,, 2014) in a multi-region framework, capturing interactions between economic growth and climate systems.
+However, Nordhaus and Yang, (1996) and Yang, (2023) do not investigate an ETS, and Nordhaus, 2010b does not solve a problem with an ETS under noncooperation.
+Following the RICE model, a group of studies have explored climate policy in a dynamic multi-regional framework under cooperation and noncooperation.
+For instance, Luderer et al., (2012) and Jakob et al., (2012) compare the long-term predictions of three-region energy-economy models under specific environmental targets, such as stabilizing the atmospheric CO2 concentrations at 450 ppm.
+However, these studies do not incorporate actual NDCs or regional net-zero targets to assess the region-specific emissions and economic pathways.
+Other studies, such as van der Ploeg and de Zeeuw, (2016) and Jaakkola and Van der Ploeg, (2019), focus on stochastic events, including climate tipping points and technological breakthroughs, comparing the results under different levels of cooperation.
+Cai et al., (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation.
+Hambel et al., (2021) extend the RICE framework by integrating endogenous international trade under noncooperation and provide a closed-form analytical solution for the regional SCCs under certain model assumptions.
 Iverson and Karp, (2021) study a Markov perfect equilibrium in a dynamic game with a social planner deciding climate policies and non-constant discount rates.
-Cai et al., (2023) build a dynamic IAM for two economic regions (North and Tropic/South) to compute regional SCCs under cooperation and noncooperation.
-Nonetheless, none of these works examine the dynamics of an international ETS with an endogenous market price of emission permits under a noncooperative setting.
-Our study builds on the concept of regional SCCs defined in van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), and provides a comprehensive analysis of regional SCCs and their relation with the global ETS in a multi-region dynamic framework under noncooperation.
-In the literature, ETSs have garnered attention as promising mechanisms for mitigating global emissions, along with carbon taxes, but there have been relatively limited attempts to comprehensively analyze ETSs as potential instruments for global climate policy.
-Among the most closely related studies are extended versions of the WITCH model (Bosetti et al.,, 2006): Massetti and Tavoni, (2012) analyze an ETS for Asia in particular, where emission quotas are defined according to specific climate objectives, while De Cian and Tavoni, (2012) study an ETS where countries can choose abatement and investment in innovation that lowers marginal abatement cost.
-As noted earlier, our solution mechanism differs from these studies and provides permit prices from a complete noncooperative dynamic optimization problem.
-Carbone et al., (2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, solving for the equilibrium permit price.
-However, that study examines the ETS in a static setting, missing the critical dynamic aspects of climate change and lacking a connection with emission abatement decisions.
-Fischer and Springborn, (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission cap and tax policies, albeit without consideration of emission trading between regions.
-Other strand of studies have analyzed the potential efficiency gains from integrating regional ETSs into a global ETS (Habla and Winkler,, 2018; Doda et al.,, 2019; Holtsmark and Weitzman,, 2020; Holtsmark and Midttømme,, 2021; Mehling et al.,, 2018).
+For a comprehensive review of macroeconomic models of climate change, see Fernández-Villaverde et al., (2025).
+Nonetheless, the literature has yet to examine the dynamics of an ETS with an endogenous market price of emission permits and regional SCCs under a noncooperative setting.
+This paper addresses this gap by providing a comprehensive analysis of an ETS as well as identifying the relationship between permit prices and regional SCCs under a multi-region dynamic noncooperative framework.
+This paper also adds to the extensive literature on environmental economic policy, with a particular focus on carbon pricing.
+ETSs, alongside carbon taxes, have garnered attention as promising mechanisms for reducing global emissions.
+However, comprehensive analyses of ETSs as instruments for global climate policy remain relatively limited.
+Among the most closely related studies are extensions of the WITCH model (Bosetti et al.,, 2006) that incorporate an ETS, including the analyses of regional ETSs for Asian countries (Massetti and Tavoni,, 2012), endogenous technological change (De Cian and Tavoni,, 2012), and banking of emission permits (Bosetti et al.,, 2009).
+The WITCH model considers alternative emission permit allocation schemes in which permits are distributed according to population or current emission shares, whereas our study applies emission caps consistent with the Paris Agreement and the Glasgow Climate Pact.
+In addition, our study departs from previous WITCH analyses by examining regional SCCs and their economic policy implications under an ETS, which were not explored in earlier studies.
+Carbone et al., (2009) present another relevant study, constructing a computable general equilibrium model incorporating countries’ endogenous participation in an ETS and allocation of emission permits, and solving for the equilibrium permit price.
+Yet, Carbone et al., (2009) examine the ETS in a static setting, missing the crucial dynamic aspects of climate change and its connection with emission abatement decisions.
+Fischer and Springborn, (2011) develop a dynamic stochastic general equilibrium model to compare outcomes of emission caps and tax policies, but they do not allow for emission trading between regions.
+Another group of studies examines the potential efficiency gains from integrating regional ETSs into a global system (Habla and Winkler,, 2018; Doda et al.,, 2019; Holtsmark and Weitzman,, 2020; Holtsmark and Midttømme,, 2021; Mehling et al.,, 2018).
+Some studies highlight that differentiated emission pricing under ETSs based on location-specific damages can be welfare improving (Muller and Mendelsohn,, 2009; Holland and Yates,, 2015; Fowlie and Muller,, 2019).
 More broadly, a substantial body of literature has examined ETSs from various perspectives.
-While in this study, we assume that emission permits are freely allocated to the regions following the voluntary commitments from the NDCs, several studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins,, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry,, 2008; Goulder et al.,, 2010).
-Other works have studied the relative efficiency of carbon policies under cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions, and vice versa (Weitzman,, 1974; Stavins,, 1996; Karp and Traeger,, 2024).
+Several studies have analyzed how the initial allocation of permits affects the equilibrium outcomes (Hahn and Stavins,, 2011), or have compared allocation methods between auctioning and free allocation (Goulder and Parry,, 2008; Goulder et al.,, 2010).
+Other studies have compared the relative efficiency of carbon policies under cost uncertainty, showing that carbon tax can be more efficient than ETS under certain conditions, and vice versa (Weitzman,, 1974; Stavins,, 1996; Karp and Traeger,, 2024).
 In another strand of literature, a number of studies have focused on empirically analyzing the regional emission trading markets currently in practice, such as the EU ETS (Hitzemann and Uhrig-Homburg,, 2018; Fuss et al.,, 2018; Perino et al.,, 2022), China ETS (Goulder et al.,, 2022), or California ETS (Borenstein et al.,, 2019).
-An important issue with regional ETSs is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS.
+An important issue with a regional ETS is carbon leakage, which refers to the shift of emission intensive production to regions outside the jurisdiction of the ETS.
 Previous studies have compared different policy instruments aimed at mitigating this problem and providing a level-playing field to the firms operating within the ETS (Ambec et al.,, 2024; Böhringer et al.,, 2014; Farrokhi and Lashkaripour,, 2021; Fowlie and Reguant,, 2022; Levinson,, 2023).
-While there is no general agreement among economists on whether a carbon tax or an ETS is better (Stavins,, 2022), the literature have often considered these two pricing instruments (trading versus taxes) as policy substitutes.
-For instance, some studies support carbon tax over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003).
-For instance, Newell and Pizer, (2003) find that carbon taxes can yield higher welfare benefits than an ETS under such uncertainties.
-Conversely, other studies advocate ETS over carbon tax because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009).
+While there is no general agreement among economists on whether carbon taxation or ETS is better (Stavins,, 2022), previous works have often considered these two pricing instruments (trading versus taxes) as policy substitutes.
+For instance, some studies support carbon taxation over ETS, highlighting concerns about price volatility in ETS (Nordhaus,, 2007), or the presence of uncertainty regarding emission abatement costs (Newell and Pizer,, 2003).
+For instance, Newell and Pizer, (2003) find that carbon taxation can yield higher welfare benefits than ETS under such uncertainties.
+Conversely, other studies advocate ETS over carbon taxation because allocation of permits in ETS allows for more flexibility, and ETS faces less uncertainty in controlling the cumulative amount of carbon emissions compared to taxes (Keohane,, 2009).
 Harstad and Eskeland, (2010), Hahn and Stavins, (2011), and Stavins, (2022) argue that from a practical perspective, ETS may be a preferred instrument over tax, as free allowances could be negotiated among participating agents to redistribute burden and, thus be used to gain political support.
-Knopf et al., (2012) favor a global ETS because carbon taxes that increase over time may accelerate extraction of fossil fuels due to profit maximization of fossil fuel firms, an argument that has been made by other researchers as well (Sinn,, 2008).
-In a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016).
-However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms results in distortions in the market allocation.
-For a comprehensive comparison on carbon tax and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022).
-This study contributes to the literature by demonstrating that the two pricing instruments are not competing policies but rather complementary.
-2 Optimal Carbon Tax under the ETS
-While carbon taxes and ETS are often considered as substitutable policies, we demonstrate that each country needs to implement an optimal carbon tax as a complementary policy within a global ETS regime to achieve optimal level of emissions.
-Before introducing the full multi-region dynamic general equilibrium model, it is helpful to first explore the optimal carbon tax within a static framework.
-Under the global ETS framework, we assume that each region $i$ is allocated with an exogenous level of emission cap ($\bar{E}_{i}$), which is the upper limit of emissions allowed for that region and is strictly enforced.
-The region’s gross emissions before abatement is $E_{i}^{\mathrm{Gross}}$, its emissions net of abatement is $E_{i}$, the amount of emission permits purchased (or sold) in the market is $E_{i}^{P}$ (if $E_{i}^{P}<0$, it implies that the region sells emission permits) so we have the emission cap constraint $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$, where the Nash equilibrium permit price is $m$ under the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$.
-Here we assume that each region is a price taker.
-The region’s marginal abatement cost (MAC) function, denoted $\mathrm{MAC}_{i}(E_{i})$, is assumed to be monotonically decreasing over emissions $E_{i}$ with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$.
-A social planner of the region solves the following minimization problem
+Moreover, in a global ETS, emission permits will be traded globally with one price for all nations based on market forces, which may alleviate the problem of carbon leakage (Fowlie et al.,, 2016).
+However, Harstad and Eskeland, (2010) argue that although an ETS in a perfect market is the first-best system, frequent government interventions to redistribute allocations among firms may result in distortions in the market allocation.
+[^5] Environmental pollution has also been studied from the context of fiscal federalism, which considers what level of government should regulate pollution (Oates,, 1999; Ogawa and Wildasin,, 2009; Banzhaf and Chupp,, 2012; Williams III,, 2012).
+For a comprehensive comparison on carbon taxation and trading regimes, see e.g., Strand, (2013); Schmalensee and Stavins, (2017); Cai, (2021); Stavins, (2022).
+This study contributes to the literature by showing that regional carbon tax is complementary to ETS under noncooperation.
+3 A Static Framework of Climate and the Economy with an ETS
+This section presents a simplified static model framework with a global ETS to provide intuition on the relationship between the regional SCC and regional MAC.
+In a global ETS regime, the main forces determining the equilibrium are climate damages, costs (or revenue) from emissions trading, and costs from regional emission abatement efforts.
+In this simplified static framework, it can be easily shown that the regional SCC equals the regional MAC minus the market price of emission permits.
+Building on these insights, the next section presents the multi-region dynamic general equilibrium model, followed by the quantification strategies and numerical analysis.
+Consider a world economy with multiple regions under a global ETS regime, and let $\mathcal{I}$ denote the set of regions.
+Each region $i\in\mathcal{I}$ is allocated an exogenous emission cap $\bar{E}_{i}$, which represents the region’s maximum allowable emissions and is strictly enforced.
+Economic activity generates emissions, with $E_{i}^{\mathrm{Gross}}$ denoting a region’s gross emissions, which are treated as exogenous in this simplified framework.
+[^6] It is innocuous to assume that gross emissions are exogenous in this static framework.
+In the full dynamic model introduced later, regional gross emissions are proportional to GDP, which is determined by capital and the exogenous population growth.
+Given that the current level of capital is determined in the previous period, gross emissions can be considered exogenous for each period in a static equilibrium.
+To comply with the emission cap constraint, each region’s social planner can either (i) purchase or sell emission permits in the global emission market, or (ii) undertake costly abatement efforts.
+Let $E_{i}$ denote its emissions net of abatement (henceforth referred to as net emissions), and $E_{i}^{P}$ the quantity of emission permits purchased in the market.
+A positive $E_{i}^{P}$ indicates that the region purchases permits, while a negative $E_{i}^{P}$ indicates that the region sells permits.
+The emission cap constraint for each region is then given by $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$.
+In this economy, the regional social planner aims to minimize the total economic costs associated with its emissions, including (i) climate damages due to global emissions, (ii) costs (or revenue) from emission trading, (iii) and costs from own regional emission abatement.
+Formally, a social planner of each region solves the following minimization problem:
 […]
-where the first term of the objective function is the regional damage from the global emissions $\sum_{j\in\mathcal{I}}E_{j}$ (which is the present value of future damage from the global emissions), the second term is the cost from purchasing permits at the price $m$ (or the profit from selling permits), and the third term is the total abatement cost for reducing emissions by the amount $(E_{i}^{\mathrm{Gross}}-E_{i})$.
-For simplicity, we assume that emissions cannot be negative in our model.
-In the static Nash equilibrium of all regional social planners, every region solves equation (1) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$.
-Let $E_{j}^{*}$ be the Nash equilibrium emissions for each region $j$ and let $m^{*}$ be the Nash equilibrium price.
-Thus, from the Karush-Kuhn-Tucker conditions of (1), the solution to the region’s optimal emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
-$D_{i}^{\prime}\left(\sum_{j\in\mathcal{I}}E_{j}^{*}\right)=\mathrm{MAC}_{i}\left(E_{i}^{*}\right)-m^{*},$ (2)
-when $E_{i}^{*}>0$.
-The marginal regional damage can also be regarded as the regional social cost of carbon (SCC).
-Note that in the static Nash equilibrium with $E_{i}^{*}>0$, the permit price $m^{*}$ should be equal to the shadow price of the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$.
-That is, if the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ is not binding, or $\sum_{j\in\mathcal{I}}E_{j}^{*}<\sum_{j\in\mathcal{I}}\overline{E}_{j}$ (i.e., there is an over-supply of emission permits), then the permit price is zero.
-Alternatively, if the permit price is nonzero, then the inequality $E_{i}\leq\overline{E}_{i}+E_{i}^{P}$ must be binding, so the supply of emission permits is equal to the demand (i.e., $\sum_{j\in\mathcal{I}}E_{j}^{*}=\sum_{j\in\mathcal{I}}\overline{E}_{j}$).
-We emphasize that the assumption commonly used in the literature—that the MAC equals the permit price (e.g., Keohane,, 2009; Nordhaus, 2010b, )—however, does not hold in this case, unless the regional marginal damage or SCC is zero, according to equation (2).
-Moreover, we also numerically verify in our dynamic Nash equilibrium that, when $E_{i}^{*}>0$, the regional MAC is also strictly larger than the permit price, and the difference is the regional SCC.
-An important issue arises when a regional social planner cannot directly enforce its optimal emissions level on firms, the actual emission producers in the economy.
-To address this, let us assume that the social planner can employ carbon taxation as a policy instrument to resolve the discrepancy in optimal emissions between the firms and the regional social planner.
-Consider a representative firm in region $i$ that produces emissions and operates under the region’s MAC function and emissions cap constraint in a static framework.
-Since the firm does not internalize the damage from emissions, we assume that the social planner of region $i$ levies a tax on the firm of $\tau_{i}$ per unit of emissions.
-The firm solves the following minimization problem:
+The first term of the objective function represents regional climate damages, where $D_{i}(\cdot)$ is a function that captures the regional climate damages from global net emissions, $\sum_{j\in\mathcal{I}}E_{j}$.
+Here $D_{i}(\cdot)$ can also be considered as the present value of future climate damages from the current global net emissions in the corresponding dynamic framework.
+The second term accounts for the cost (or revenue) from purchasing (or selling) permits at price $m$, which is the Nash equilibrium price satisfying the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$.
+The quantity of emission permits purchased may influence the global permit price, making the market price depend on their emission permit purchase choices, which in turn are contingent on the trade decisions of other regions.
+Lastly, the third term reflects the total abatement cost incurred to reduce emissions by $(E_{i}^{\mathrm{Gross}}-E_{i})$, where $\mathrm{MAC}_{i}(\cdot)$ denotes the region’s marginal abatement cost and is assumed to be monotonically increasing over $E_{i}$, with $\mathrm{MAC}_{i}(E_{i}^{\mathrm{Gross}})=0$.
+For simplicity, it is assumed that net emissions are nonnegative throughout the paper.
+[^7] This assumption ensures that no region can abate more than its gross emissions solely to sell permits to other regions.
+In the static Nash equilibrium of the world economy, every regional social planner simultaneously solves equation (1) while satisfying the market-clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$, which implies that $\sum_{j\in\mathcal{I}}E_{j}\leq\sum_{j\in\mathcal{I}}\overline{E}_{j}$.
+We assume that each region is a price taker for permits, as a large number of firms in each region participate in the permit trading market in practice.
+Thus, the Karush-Kuhn-Tucker conditions of equation (1) imply that the solution to the region’s optimal net emissions leads to the following relationship between the marginal regional damage, the MAC, and the permit price:
 […]
-where the first term of the objective function is the total tax on emissions.
-In the static Nash equilibrium of all representative firms, every firm solves equation (3) simultaneously with the market clearing condition $\sum_{j\in\mathcal{I}}E_{j}^{P}=0$.
-Let $E_{j}^{*,\mathrm{Firm}}$ be the Nash equilibrium emissions for each firm $j$ and let $m^{*,\mathrm{Firm}}$ be the Nash equilibrium price.
-Thus, from the Karush-Kuhn-Tucker conditions of (3), the solution to the optimal emissions of the firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
-$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}},$ (4)
-when $E_{i}^{*,\mathrm{Firm}}>0$.
-If the regional social planner imposes the tax to be equal to the regional SCC, i.e., the regional marginal damage, then from equations (2) and (4), we have $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ and $m^{*,\mathrm{Firm}}=m$.
-In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal emissions and the market price of emission permits, capturing the idea that the regional social planner can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its emissions level of $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$.
-When $E_{i}^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (3) shows that
-$\tau_{i}\geq\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})-m^{*,\mathrm{Firm}}.$ (5)
-That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits.
-In particular, the regional $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})$ at $E_{i}^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m^{*,\mathrm{Firm}}$, making the right hand side of the inequality (5) to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ if it is positive.
-Therefore, for any regional social planner’s optimal emissions $E_{i}^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
-$\tau_{i}=\max\{0,\ \mathrm{MAC}_{i}(E_{i}^{*})-m^{*}\},$ (6)
-such that the firm’s optimal emissions are consistent with the social planner’s (i.e., $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$).
-The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model.
-Note that this tax is contingent on the imposition of the emission cap on the firm, and the tax level itself cannot make the firm’s emissions to be $E_{i}^{*}$ unless $m^{*}=0$.
-It can be easily shown that our argument is consistent with the existing literature when there is no ETS.
-The cost minimization problem of the representative firm facing the emission cap regulation without the ETS becomes:
-[…]
-From the Karush-Kuhn-Tucker conditions of (7), when the firm’s optimal solution $E_{i}^{*,\mathrm{Firm}}=E_{i}^{*}$ is not binding at its lower or upper bound, we have the well known equality:
-$\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*}).$ (8)
-This implies that the optimal carbon tax should be set equal to the MAC at the regional social planner’s optimal emissions $E_{i}^{*}$, provided $E_{i}^{*}$ is not binding.
-This result has been widely used in the literature (e.g., Nordhaus,, 2014; Cai et al.,, 2017; Cai and Lontzek,, 2019).
-Note that our formula (6) under an ETS with zero permit price is consistent with (8) under no ETS.
-A regional social planner in each region $i$ has two targets: one is its exogenous NDC commitment $\overline{E}_{i}$, and the other is the regional social planner’s endogenous optimal emissions $E_{i}^{*}$ after permit trading (note that $\overline{E}_{i}$ and $E_{i}^{*}$ are unequal in general [^5] We believe this is a realistic consideration, as emission cap pathways may not always align precisely with regional optimal emissions over time.
-This discrepancy arises because the negotiation of emission permit allocations is often time-consuming and influenced by political considerations, which may not lead to economically optimal outcomes.
-Our model simulation shows that the emission cap pathways constructed from the Paris Agreement and the Glasgow Climate Pact indeed deviate from the optimal regional emissions.) in the Nash equilibrium of all 12 regional social planners.
-That is, the region $i$’s emissions should be $E_{i}^{*}$, which should also satisfy $E_{i}^{*}\leq\overline{E}_{i}+E_{i}^{P}$ after permit trading in the Nash equilibrium.
-Several policy instruments can be applied to enforce the firms to meet the targets, assuming that the firms can buy or sell emision permits.
-One such instrument is a quantity policy, in which the regional social planner distributes $\overline{E}_{i}$ (e.g., by means of an auction or grandfathering) to its firms, ensuring the firms’ aggregate emissions, $E_{i}^{*,\mathrm{Firm}}$, do not exceed the exogenous allocated cap after permit trading (i.e., $E_{i}^{*,\mathrm{Firm}}\leq\overline{E}_{i}+E_{i}^{P}$) in the Nash equilibrium of firms of all regions.
-[^6] The representative firm framework can be generalized to an environment with multiple firms.
-Assume that there are $n$ homogenous firms in the region $i$, each having $\overline{E}_{i}/n$ as their caps, with a firm’s MAC function at its emissions $e$ is defined as $\mathrm{MAC}_{i}(ne)$, where $\mathrm{MAC}_{i}(\cdot)$ is the aggregate regional MAC function used by the regional social planner and $E=ne$ is the aggregate regional emissions.
-Under this definition, the aggregate total abatement cost of all firms in region $i$ is equal to the integral of $\mathrm{MAC}_{i}(E)$ on $[E_{i},E_{i}^{\mathrm{Gross}}]$, when each firm’s emission is $E_{i}/n$ while $E_{i}^{\mathrm{Gross}}/n$ is each firm’s upper bound under no abatement.
-In this case, the Nash equilibrium of the economy with multiple firms is identical to the equilibrium with a representative firm.
-However, under a quantity-only control policy, $E_{i}^{*,\mathrm{Firm}}$ is not equal to $E_{i}^{*}$, as the zero-tax Nash equilibrium of the firms, $\left\{m^{*,\mathrm{Firm}},E_{i}^{*,\mathrm{Firm}}:j\in\mathcal{I}\right\}$, is not the Nash equilibirum of the regional social planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$.
-This becomes particularly evident in the following extreme case: assume that $\overline{E}_{i}$ is larger than $E_{i}^{\mathrm{Gross}}$ for every region, which leads to $m^{*}=0$ under the ETS.
-The firms’ Nash equilibrium without a tax yields $E_{i}^{*,\mathrm{Firm}}=E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*,\mathrm{Firm}})=m^{*,\mathrm{Firm}}=0$.
-In contrast, the regional social planners’ Nash equilibrium leads to $E_{i}^{*}<E_{i}^{\mathrm{Gross}}$, with $\mathrm{MAC}_{i}(E_{i}^{*})>m^{*}=0$ in general, similar to a model without emission caps.
-Note that this quantity policy cannot directly distribute $E_{i}^{*}$, as doing so would either result in the region violating its commitment if $E_{i}^{*}>\overline{E}_{i}$ or losing the benefit in selling permits if $E_{i}^{*}<\overline{E}_{i}$.
-Another instrument is a carbon tax policy, in which the region $i$ imposes a carbon tax $\tau_{i}$ on the firms, e.g., $\tau_{i}=\mathrm{MAC}_{i}(\overline{E}_{i})$ or $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})$.
-However, a carbon tax policy alone can target only one emission level, either $\overline{E}_{i}$ or $E_{i}^{*}$, in each region, while $\overline{E}_{i}\neq E_{i}^{*}$ in general.
-Therefore, neither instrument can achieve both targets simultaneously.
-We argue that in the presence of a global ETS, it is optimal for each region $i$ to impose both the quantity policy and an additional carbon tax policy simultaneously on firms to enforce both targets.
-Here the additional carbon tax for region $i$ is chosen to be $\tau_{i}=\mathrm{MAC}_{i}(E_{i}^{*})-m^{*}$ when $E_{i}^{*}>0$, under the Nash equilibirum of the planners, $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$.
-If some regions impose only the quantity policy while the others impose both policies, this leads to a Nash equilibrium of firms, which is not the Nash equilibirum of the planners.
-Note that a tax has no impact on the regional social planners’ decisions; $E_{i}^{*}$ remains the optimal decisions of the regional social planners regardless of the tax levels.
-However, the tax can influence the firms’ decisions.
-Therefore, from the regional social planners’ perspective, any Nash equilibrium of firms deviating from $\left\{m^{*},E_{j}^{*}:j\in\mathcal{I}\right\}$ is worse, and the regions should always impose both the quantity policy and the optimal additional carbon taxes simultaneously on the firms to meet both targets.
-[^7] Our argument is based on the assumption that the emission caps are exogenous in the regional level.
+when $E_{i}>0$.
+The left hand side of this equation represents the marginal regional damage, which corresponds to the regional SCC.
+This term captures the additional regional economic cost imposed by a unit increase in global emissions, reflecting the region’s contribution to global climate change.
+The right hand side captures the difference between $\mathrm{MAC}_{i}(E_{i})$ and $m$, reflecting the deviation of regional abatement efforts from the market equilibrium.
+The equation (2) simplifies to $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$, where $\mathrm{SCC}_{i}$ denotes the regional SCC.
+A positive regional SCC implies $\mathrm{MAC}_{i}(E_{i})>m$, suggesting that a region experiencing climate damages has an incentive to abate emissions to a level where the marginal abatement cost exceeds the market price of emission permits.
+In the environmental economics literature, it is well established that when production generates emissions as a negative externality in an economy, a regulator can achieve the socially efficient level of abatement by setting an emission cap at that level and issuing tradable permits to firms.
+The resulting market equilibrium under cap-and-trade is equivalent to the outcome of an optimal emission tax equal to the MAC at the efficient abatement level, in the absence of cost uncertainty or transaction costs.
+However, in a multi-region framework where climate damages are region-specific, the decentralized equilibrium outcome under the global ETS does not necessarily coincide with the regional social planner’s optimal outcome.
+The underlying reason is that, because climate damages are heterogeneous across regions, a uniform global carbon price alone may not yield regionally optimal levels of emission abatement or the corresponding emission prices.
+The intuition behind $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ can be further illustrated through an extreme case.
+If emission caps are sufficiently large such that they are not binding for any region—effectively a scenario without emission constraints—permit prices and trading volumes would be zero.
+Imposing the equilibrium condition $\mathrm{MAC}=m$ in this situation would imply no abatement (i.e., $\mathrm{MAC}=0$ when $m=0$).
+However, it is well established that even in the absence of binding caps, the optimal level of emission abatement from the perspective of a regional social planner remains strictly positive (i.e., $\mathrm{SCC}=\mathrm{MAC}>0=m$) under noncooperative Nash equilibrium (see, e.g., Nordhaus and Yang,, 1996; Cai et al.,, 2023).
+This example further highlights the gap between the regional social planner’s optimal abatement level and the decentralized market outcome, as firms undertake no abatement in the absence of an emission price.
+The basic intuition explored here carries over to the dynamic general equilibrium model introduced in the next section, where we further confirm numerically that $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ continues to hold.
+An additional important implication of this relationship is that the two pricing instruments—emission taxes and cap-and-trade—are not competing policies but rather complementary under the global ETS framework.
+Although our model does not explicitly introduce firms, the regional MAC is effectively obtained as an aggregation of individual firms’ marginal abatement quantities at a given price level within a region (Keohane and Olmstead,, 2016).
+To achieve the optimal level of net emissions, $E_{i}$, from each regional social planners perspective under the global ETS, the optimal policy for each regional planner is to impose, within its own economy, a carbon tax equal to its regional $\mathrm{SCC}_{i}$ on top of the permit price from the global ETS.
+This ensures that individual firms internalize both the regional SCC and the global permit price.
+With the regional tax incorporated, each individual firm chooses net emissions such that its own MAC equals the sum of the permit price and the regional carbon tax, i.e., $\mathrm{MAC_{i}}(E_{i}^{\prime})=\mathrm{\tau}_{i}+m^{\prime}$, where the optimal carbon tax is $\mathrm{\tau}_{i}=\mathrm{SCC}_{i}$ from the perspective of the regional social planner.
+For a partial ETS, if region $i$ participates in the ETS, then the relation $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})-m$ still holds, where $m$ is the equilibrium permit price between the participating regions; if region $i$ does not participate, then the relation is changed to $\mathrm{SCC}_{i}=\mathrm{MAC}_{i}(E_{i})$, which is a standard relation under no ETS.
 4 A Dynamic Regional Model of Climate and the Economy with an ETS
-We now introduce a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions.
+We now introduce a dynamic regional model of climate and economy that integrates a global ETS across multiple regions.
 Our model framework extends the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps.
-The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption.
-The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets.
-In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
+One of the main focuses of the model is to characterize the equilibrium path of carbon prices in a noncooperative environment, where a social planner of each region maximizes its own social welfare by optimally choosing emissions abatement, permit purchases in the global carbon market, and consumption over time.
+Future regional emissions are constrained by emission caps based on commitments established under the Paris Agreement, later updated by the Glasgow Pact, and by net-zero targets.
 The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model.
-We follow the RICE model and assume that the world consists of 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs).
-[^8] See Appendix 1 A.2 for the full list of countries and regional aggregation.
 Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions.
-Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial year is the year 2020.
-All regions are forward-looking with complete information.
+Time is discrete and infinite, with annual time steps indexed by $t=0,1,2,\cdots$.
+All agents are forward-looking with complete information.
+The model presented here abstracts from uncertainty, excludes international trade of goods other than emission permits, and assumes frictionless trading in the ETS.
+4.1 The Economic System
+Each region consists of a representative household with a population size of $L_{i,t}$, and utility of the representative consumer is given by
+$u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (3)
+where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. Following DICE-2016R (Nordhaus,, 2017), $\gamma$ is set to 1.45.
+There is a representative firm in each region that employs a Cobb-Douglas production technology using capital and labor as inputs, and produces a numeraire good whose price is normalized to 1.
+The representative household owns all input factors and the firm of the regional economy.
+The gross output, or pre-damage output, $Q_{i,t}$, is given by
+$Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (4)
+where $K_{i,t}$ is capital stock, and $\alpha=0.3$ is the elasticity of gross output with respect to capital, as in Nordhaus, (2017).
+Consistent with the standard neoclassical growth model, capital depreciates over time, and the firm invests to replenish and accumulate capital stock.
+The evolution of capital follows
+$K_{i,t+1}=(1-\delta)K_{i,t}+I_{i,t},$ (5)
+where $I_{i,t}$ is investment and $\delta=0.1$ is the rate of depreciation of capital stock.
+Regions experience climate damages resulting from the externalities of global emissions, with the extent of these damages varying across regions.
+The output net of climate-induced damages, denoted as $Y_{i,t}$, is given by:
+$Y_{i,t}=\left(\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}\right)Q_{i,t},$ (6)
+where $\pi_{1,i}$ and $\pi_{2,i}$ are region-specific climate damage parameters, and $T_{t}$ is the global average temperature increase in degrees Celsius above the pre-industrial level.
+The specification captures the adverse (or potentially beneficial for some regions) effects of rising global average temperature, where local damages increase nonlinearly with temperature, following the quadratic form commonly used in the climate economics literature (Nordhaus,, 2014, 2017).
 4.2 Emissions
-The gross carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
+Each region’s economic activity produces carbon emissions, and the representative firm faces emission constraints of the region.
+The gross emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
 $E_{i}^{\mathrm{Gross}}=\sigma_{i,t}Q_{i,t},$ (7)
-where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al., (2019).
-We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the gross emissions, with an emission control rate $\mu_{i,t}\in[0,1]$.
-[^9] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020).
-However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated.
-The emission control rate concept simplifies this.
-That is,
+where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity.
+The firm may choose to reduce a fraction of gross emissions, with its efforts represented by emission control rate $\mu_{i,t}\in[0,1]$.
+[^8] Alternatively, abatement can also be modeled through the reduced use of fossil fuel energy inputs in the production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020).
+However, such models often require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producers, and other sectors, increasing the complexity of the model.
+The emission control rate approach simplifies this by offering a more streamlined representation.
+The amount of emissions abated by each region, $E_{i,t}^{A}$, is then expressed as follows.
 $E_{i,t}^{A}=\mu_{i,t}\sigma_{i,t}Q_{i,t}.$ (8)
-Thus the emissions net of abatement, $E_{i,t}$, becomes
-$E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}.$ (11)
-We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions.
-Regions with emissions exceeding the emission caps can purchase emission permits from other regions.
-Regions with emissions below their allowance can sell their emission permits to others.
-The emission cap constraint is represented by
-$E_{i,t}-E_{i,t}^{P}\leq\overline{E}_{i,t},$ (12)
-where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions traded with other regions.
-Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$.
-In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$.
-The Paris Agreement establishes upper limits for emissions to ensure that the increase in global temperature remains well below 2°C above pre-industrial levels by the end of this century.
-3.2 Climate System
-The global average temperature rises as carbon emissions accumulate in the atmosphere.
-We follow the TCRE climate system and assume that the global average temperature increase above the pre-industrial level is approximately proportional to cumulative global emissions $\mathcal{E}_{t}$, i.e.,
-$T_{t}=\zeta\mathcal{E}_{t},$ (13)
-where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2°C above the pre-industrial level.
-Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011).
-Figure A.1 in Appendix A.3.1 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways.
-Cumulative global emissions evolve according to
-$\mathcal{E}_{t+1}=\mathcal{E}_{t}+\sum_{i\in\mathcal{I}}E_{i,t}.$ (14)
-3.3 Economic System
-A representative consumer in region $i$ has a power utility per period as follows:
-$u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (15)
-where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus, (2017) and set $\gamma=1.45$.
-The production technology follows a Cobb-Douglas function of capital and labor inputs. The gross output (in trillion USD), or pre-damage output, $Q_{i,t}$, is given by
-$Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (16)
-where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus, (2017)) is the elasticity of gross output with respect to capital.
-The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the GDP projections in Burke et al., (2018).
-Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz,, 2017).
-We denote $Y_{i,t}$ as the output net of damages induced by climate changes:
-$Y_{i,t}=\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}Q_{i,t},$ (17)
-where $T_{t}$ is the global average temperature increase in °C above the pre-industrial level.
-We calibrate the climate damage parameters based on the projections of GDP in Burke et al., (2018).
-We follow DICE and assume that the emission abatement cost is
+The emission abatement efforts incur costs to the firm and are heterogeneous across regions due to technological differences. The emission abatement cost, $\Phi_{i,t}$, is specified as
 $\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ (9)
 where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$.
-We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al., (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
-The total consumption of each region under the ETS is obtained by
+The parameters $b_{1,i}$, $b_{2,i}$, $b_{3,i},$ and $b_{4,i}$ govern the cost structure of emission abatement cost, $\Phi_{i,t}$, which depends on both the gross output $Q_{i,t}$ and the exponential function of the emission control rate, $\mu_{i,t}$.
+This specification captures the dynamic nature of abatement costs, reflecting potential technological advancements that reduce the cost of emissions abatement over time.
+In a global ETS, or a cap-and-trade system, each region is provided with an emission allowance and can trade emission permits with other regions.
+The representative firm in each region that emits beyond its cap can purchase permits, while those emitting below their allowance can sell excess permits.
+Denoting the emissions net of abatement (net emissions) as $E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}$, the emission cap constraint is represented by
+$E_{i,t}-E_{i,t}^{P}\leq\overline{E}_{i,t},$ (10)
+where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions purchased from other regions.
+Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$.
+In a model that does not consider emission permit trade, the emissions purchase is simply set at $E_{i,t}^{P}=0$ for all $i$ and $t$.
+4.3 The Climate System
+The global average temperature rises as carbon emissions accumulate in the atmosphere.
+Adopting the TCRE climate system representation (Matthews et al.,, 2009), it is assumed that the global average temperature increase above the pre-industrial level is approximately linear to cumulative global emissions $\mathcal{E}_{t}$, i.e.,
+$T_{t}=\zeta\mathcal{E}_{t},$ (11)
+where $\zeta$ represents the contribution rate of cumulative global emissions to temperature. The cumulative global emissions evolve according to
+$\mathcal{E}_{t+1}=\mathcal{E}_{t}+\sum_{i\in\mathcal{I}}E_{i,t}.$ (12)
+This dynamic process captures the accumulation of emissions in the atmosphere over time, with each region contributing to the global emissions stock and, consequently to the increase in global temperature.
+4.4 Market Clearing
+The goods market clearing implies that the total consumption of each region under the ETS is constrained by
 $c_{i,t}L_{i,t}=Y_{i,t}-I_{i,t}-\Phi_{i,t}-m_{t}E_{i,t}^{P},$ (13)
-where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits.
-The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption, but this impact is little as the cost or profit from permit trading is small relative to the consumption or output.
-The capital stock evolves according to the standard neoclassical growth model:
-$K_{i,t+1}=(1-\delta)K_{i,t}+I_{i,t},$ (20)
-where $\delta=0.1$ is the rate of depreciation of capital stock. We assume that the emission trading market clears in each period. That is,
-$\sum_{i\in\mathcal{I}}E_{i,t}^{P}=0,$ (21)
-for each time period $t$.
+where $m_{t}$ is the market equilibrium price of emission permits.
+As in the static framework, the last term on the right-hand side reflects the cost or revenue generated from the emission permit trade.
+Finally, the emission trading market clears each period, given by: [^9] The current model assumes that intertemporal lending or borrowing of emission permits is not allowed.
+$\sum_{i\in\mathcal{I}}E_{i,t}^{P}=0.$ (14)
+If it is a partial ETS and region $i$ does not participate in the partial ETS, then we set $E_{i,t}^{P}$ to be fixed at zero.
 5 Solving for the Equilibrium
@@
 We follow DICE-2016R (Nordhaus,, 2017) and set $\beta=0.985$.
-Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simulatenously as a dynamic game.
+Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simultaneously as a dynamic game.
 Then we define the dynamic Nash equilibrium of the economy as follows.
@@
 In particular, finding an equilibrium solution for the emission permit prices that satisfies the optimality conditions for each region as well as the market clearing condition poses significant computational challenges.
-While several studies have incorporated emission prices in their models, no study has successfully solved for the emission permit price in a noncooperative context.
 Here we outline the algorithm we develop to obtain the optimal solution for our model.
-Since $\gamma=1.45$ implies utilities of per capita consumption have negative values and that long-run per capita consumption will be large, the long-run per-capita utility has a small magnitude.
-Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years.
+With the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years.
 Let
@@
 Maximization Step at iteration $j$.
-For each region $i$, solve the maximization problem (22) without the market clearing condition (21), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration $j-1$ when $j>1$.
+For each region $i$, solve the maximization problem (15) without the market clearing condition (14), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration j1j-1 when $j>1$.
 The optimal emissions and permits purchased for region $i$ are denoted $\{E_{i,t}^{*,j},E_{i,t}^{P,j}:\ t\geq 0\}$.
@@
 6 Data and Calibration
-To calibrate our model with data, we have two objectives.
+When taking our model to data, two key objectives are pursued.
 The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their emissions commitments under the Glasgow Pact and the long-term net zero emission targets.
@@
 We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
-Our model’s 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ).
-For the initial year 2020, we obtain country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (GtC) from the World Bank.
-For future projections on population growth, we use the SSP2 scenario (Samir and Lutz,, 2017).
+We assume that the world is divided into 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs).
+These 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ).
+[^10] See Appendix 1 A.2 for the full list of countries and regional aggregation.
+Note that, while we present our simulation results at a regionally aggregated level for computational tractability, our methods are applicable to a larger number of regions and can be readily extended to obtain country-level outcomes.
+For example, the regional SCC for Africa can be disaggregated to approximate country-level SCCs, since the regional SCC is equal to the sum of the SCCs of the individual countries within the region.
+[^11] Regional SCC is the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period.
+The initial year is set to 2020, with country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (CO2 equivalent, GtC) sourced from the World Bank.
+[^12] To obtain initial capital stock, we used capital formation data from the World Bank.
+Specifically, we computed $K_{i,2020}=\sum_{g=0}^{20}(0.9)^{g}\text{CF}_{i.2020-g}$, where $\text{CF}_{i,t}$denotes capital formation adjusted to 2020 values in trillions of USD.
+For future population pathways, projections from the SSP2 scenario (Samir and Lutz,, 2017) are employed.
 6.1 Regional Emission Cap Pathways
@@
 These near-term targets were further strengthened during the Glasgow Climate Pact in 2021, with most countries and regions submitting updated or new NDCs.
-[^10] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated Nationally Determined Contributions (NDCs) after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker.
+[^13] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated NDCs after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker.
 [^14] See https://climateactiontracker.org/.
@@
 To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al., (2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation.
-As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun451’ in Ueckerdt et al., (2019).
-Specifically, this FFrun451 scenario corresponds to weak policy baseline before Paris Agreement with limited renewable technology portfolio and zero carbon tax.
-Based on the equation (9), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun451}}^{U}/Q_{i,t,\mathrm{FFrun451}}$, where $E_{i,t,\mathrm{FFrun451}}^{U}$ and $Q_{i,t,\mathrm{FFrun451}}$ are the projected regional emissions and GDP under the FFrun451 scenario for region $i$ at time $t$.
+As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun111’ in Ueckerdt et al., (2019).
+Specifically, this FFrun111 scenario corresponds to climate action from 2010 with full technology portfolio and no carbon tax.
+.
+Based on the equation (7), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun111}}^{U}/Q_{i,t,\mathrm{FFrun111}}$, where $E_{i,t,\mathrm{FFrun111}}^{U}$ and $Q_{i,t,\mathrm{FFrun111}}$ are the projected regional emissions and GDP under the FFrun111 scenario for region $i$ at time $t$.
 [^18] Ueckerdt et al., (2019) provide data for 11 regions.
@@
 At first, we estimate the associated emission control rate $\mu_{i,t,j}^{U}$ using the equations (7)-(8) as follows:
-$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun451}}^{U}},$ (28)
-where $E_{i,t,\mathrm{FFrun451}}^{U}$ is the regional emission from the FFrun451 scenario with zero carbon tax (see Ueckerdt et al., (2019) for details).
-Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (27) can hold in an approximate manner for every scenario $j$ and time $t$.
-6 Results
-In this section, we discuss our simulation results.
-We first present the results of the noncooperative model under the baseline emission cap scenario for the period from 2020 to 2100.
-Next, we compare the economic outcomes across different models, examining the role of the emission caps and the implementation of a global ETS.
-Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all 12 regions.
-6.1 Key Results under the Baseline Emission Caps
-Figure 1 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario.
+$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun111}}^{U}},$ (21)
+[…]
+6.5 Climate System: Transient Climate Response to Emissions
+In the TCRE climate system, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ in year 2020 is 1.2 degrees Celsius above the pre-industrial level.
+The contribution rate of cumulative global emissions to temperature is calibrated at $\zeta=0.0021$ by using the projections of emissions and temperatures across the four Representative Carbon Concentration Pathways (RCPs) scenarios: RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011).
+That is, we solve the following minimization problem:
+$\min_{\zeta}$ $\sum_{t}\sum_{j}\left\|T_{t}^{j}-\zeta\mathcal{E}_{t}^{j}\right\|$
+$s.t.$ $\mathcal{E}_{t+1}^{j}=\mathcal{E}_{t}^{j}+E_{t}^{j},\ \forall t,j,$
+where $j$ represents one of four RCP scenarios, $E_{t}^{j}$ and $T_{t}^{j}$ are the exogenous projections of emissions and temperatures, respectively, at time $t$ for RCP scenario $j$.
+Figure A.1 in Appendix A.3.1 demonstrates that the calibrated TCRE climate system matches well all four RCP temperature pathways using their associated RCP emission pathways.
+7 Baseline Model Results
+This section presents the numerical simulation results under the assumption that all regions participate in the global ETS.
+The baseline emission cap scenario is defined in accordance with the NDCs in the near term and the net-zero targets in the mid- to long term.
+We report the results through the end of the century (2020–2100), focusing on both economic and environmental outcomes across key variables.
+Additionally, we provide an in-depth analysis of the numerical simulation results to evaluate the effects of implementing a global cap-and-trade system.
+7.1 Emissions, Temperature, and Permit Prices
+Figure 1 displays key simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario.
 Until 2100, we can delineate three periods according to the global emission trading patterns.
-Prior to 2025, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5).
+Prior to 2024, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5).
 This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al.,, 2018), resulting in zero or very low permit prices.
-This is also consistent with our theoretical analysis for the regional social planners’ Nash equilirbium in Section 2.
 During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 1 top-left and bottom-right).
 This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
-Global emissions are constrained by the global emission caps starting from 2025.
+Global emissions are constrained by the global emission caps starting from 2024.
 To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits.
 This results in a substantial increase in the volume of emission abatement, along with the abatement cost and permit price.
-The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2060 (Figure 1 top-left).
-The global abatement cost also increases steeply to reach its maximum value of \$5.45 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
-The emission permit trade gradually decreases to zero by 2070 (Figure 1 top-left).
-As the emission cap becomes tighter over the years, the permit price increases to \$845 per ton of carbon in 2050 and \$2,059 in 2069 (Figure 1 bottom-left).
-The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5.
+The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2070 (Figure 1 top-left).
+The global abatement cost also increases steeply to reach its maximum value of \$7.82 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070.
+The emission permit trade also gradually decreases to zero by 2070 (Figure 1 top-left).
+As the emission cap becomes tighter over the years, the permit price increases to \$923 per ton of carbon in 2050 and \$2,105 in 2069 (Figure 1 bottom-left).
+The kinks in the permit price path in 2030 result from a kink in emission cap pathways (see Figure A.6), while those in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5.
 Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
 Figure 1: Simulation results at the global scale under the baseline emission cap scenario.
-Post-2070 is the period when global emissions are net zero.
+Post-2070 is the period when global emissions are at net zero.
 For this period, emission permits are no longer traded and all emissions are abated in each region.
 [^19] In the bottom-left panel of Figure 1, we plot the emission permit price only when the traded volume is positive.
-By the end of this century, the atmospheric temperature is projected to reach 1.7°C above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps.
+By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps.
 [^20] If there is no emission cap, then the temperature anomaly will be much higher.
@@
 Lastly, the global climate damage rises almost linearly over the entire period (Figure 1 bottom-right).
-Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5°C.
+Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
 It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers.
 Figure 2 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales.
-Before 2025, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage.
-After 2025, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2036.
+Before 2024, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage.
+After 2024, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2032.
 Japan is expected to be involved in a relatively small volume of permit trading, and the other regions (Africa, MidEast, Latin America, and OthAs) change their status of permit suppliers to buyers or vice versa over time.
 Figure 2: Simulation results of regional emission permit purchase under the baseline emission cap scenario.
-6.2 SCC, MAC, and the Optimal Carbon Tax
-Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the MAC, regional social cost of carbon (SCC), and the optimal carbon tax under the global ETS regime.
+7.2 Regional MAC and SCC
+Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the regional MAC, SCC, and the market equilibrium emissions permit price under the global ETS regime.
 The MAC captures the additional cost incurred due to an increase in emission abatement.
-From our model equation (10), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD.
+From our model equation (8), the total abatement cost is $\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD.
 Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows:
-$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (29)
-With $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^18] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters.
+$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\Phi_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (22)
+Note that, with $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^21] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters.
 Figure 3 displays the regional MAC along with the market equilibrium price of emission permits.
-We find that, for every region, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s.
-[^19] For regional optimal emissions, see Figure 11 in Appendix A.4.1.
-As discussed in Section 2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ implies that the optimal emissions for the regional social planner do not align with the optimal emissions for the representative firm under no tax in each region.
-Therefore, the optimal regional carbon tax can be introduced at $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s emissions are equal to the regional social planner’s optimal emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050 or 2070.
-Russia is an exception, where the MAC falls below the permit price starting in 2025, even though it has nonzero emissions until 2050, because global warming leads to increase in productivity rather than causing climate damages in Russia.
-This result is strongly influenced by the climate damage parameters, calibrated using projections from Burke et al., (2018).
+The simulation result shows that, for most regions, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s.
+[^22] For regional optimal emissions, see Figure A.7 in Appendix A.4.1.
+Russia is an exception, where the MAC falls below the permit price starting in 2021, which we will elaborate shortly in the next paragraph.
+The MAC of each region gradually decreases below the permit price after its net zero emission level is achieved (i.e., $E_{i,t}^{*}=0$).
+Russia is the first region to achieve net zero emissions in 2048 (as shown in Figure A.7 in Appendix A.4.1).
+However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario.
+After 2048, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit.
+[^23] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large.
+Similar trends are observed in other regions as they attain net zero emissions in the 2050s (China, Latin America, MidEast, the US, the OHI, Eurasia) and the 2060s (the EU, Japan and India).
+Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
+Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (10) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
 Figure 3: Simulation results of regional MAC under the baseline emission cap scenario.
-In Figure 3, the MAC of each region gradually decreases below the permit price after its net zero emissions is achieved (i.e., $E_{i,t}^{*}=0$).
-Russia is the first region to achieve net zero emissions in 2050 (as shown in Figure A.7 in Appendix A.4.1).
-However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario.
-After 2050, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit.
-[^20] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large.
-Similar trends are followed by other regions as they attain net zero emissions in the 2050s ( China, the US, the EU, the OHI, MidEast, Eurasia, and Latin America) and the 2060s (Japan and India).
-Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time.
-Recall that, when the regional net zero emissions (before trading of permits) are achieved (i.e., $E_{i,t}^{*}=0$), the regional optimal carbon tax should be larger than the gap between the regional MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$.
-Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (12) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
-Given the simulation results for the optimal carbon tax, we now turn to discuss the relationship between the optimal carbon tax and the regional social cost of carbon (SCC).
 The SCC, a central concept in the climate change literature, is widely used to quantify the present value of climate damages induced by an additional unit of carbon emissions.
@@
 for each region $i$ under the open loop Nash equilibrium.
-Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon.
 [^24] To compute the regional SCC in the noncooperative model, it is equivalent to replace the numerator in equation (23) with the shadow price of the transition equation of cumulative global emissions (12) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (5) for each region.
-Our concept of the regional SCC differs from the definition in Nordhaus, (2017) and Ricke et al., (2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous.
-We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the Nash equilibrium.
-With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional emissions are not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional emissions are binding at zero.
-This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al., (2017) and Cai and Lontzek, (2019).
-[^22] We also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
+Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in monetary unit of 2020 USD per ton of carbon.
+[^25] The choice of discount rates has been a critical factor contributing to the gap in social cost of carbon estimates across different studies (Guo et al.,, 2006; Weitzman,, 2013).
+Our concept of regional SCC differs from the definition in Nordhaus, (2017), where the regional SCC is obtained by assuming that it is a fraction of the global SCC, with the share determined by the discounted value of regional output using an exogenous, constant discount rate.
+We verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
+The simulation results confirm the relationship between the regional SCC, MAC, and permit prices in a multi-region economy with a global ETS, as explored in Section 3.
+Figure 4 demonstrates that, when a region’s net emissions are strictly positive, the regional SCC is exactly equal to the gap between the regional MAC and the permit price shown in Figure 3.262626When regional net emissions are at zero, the regional SCC can be larger than the gap, $\left(\mathrm{MAC}_{i}(E_{i})-m^{*}\right)$, which could be negative, as shown in Figure 4.
+For example, in 2050, the MAC for the US is \$1,159 per ton of carbon, the permit price is \$923, and their difference is exactly equal to the regional SCC of the US, \$236.
+Russia is an exceptional case with a negative SCC in our simulation, suggesting that global warming creates benefits rather than causing climate damages in Russia, a result heavily influenced by the climate damage parameters calibrated using projections from Burke et al., (2018).
+Nevertheless, the relationship between the SCC, MAC, and permit price remains intact; for example, in 2040, the SCC for Russia is $\mathdollar 41$, which corresponds to the difference between the MAC (\$602) and the permit price (\$643).
+This explains why the MAC falls below the permit price starting in 2021, despite Russia having nonzero emissions until 2050.
+Among all the regions, the US has the highest SCC in the near term, at around \$199 per ton of carbon in 2030.
+However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$675 for Africa and \$599 for non-OECD Asia by the end of the century.
+Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period.
+Our results show considerable heterogeneity in the SCC across regions.
 Figure 4: Simulation results of regional SCC under the baseline emission cap scenario.
-The simulation results confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS.
-Figure 4 shows that when the emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 3.
-For example, in 2050, the MAC for the US is \$1,081 per ton of carbon, the permit price is \$845, and their difference is exactly equal to the regional SCC of the US, \$236.
-Figure 4 also shows that when the regional emissions are zero, the regional SCC is larger than the gap, which could be negative.
-Among all the regions, the US has the highest SCC in the near term, at around \$198 per ton of carbon in 2030.
-However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$682 for Africa and \$610 for non-OECD Asia by the end of the century.
-Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period.
-Our results show considerable heterogeneity in the SCC across regions and suggest that significant regional carbon tax implementation is required, not only for developed countries such as the US, but also for developing countries like Africa, India, and non-OECD Asia under the global ETS with the baseline emission cap scenario.
 7.3 Effects of ETS Implementation
@@
 The top two panels of Figure 5 show that the ETS implementation results in slightly higher emissions and temperature increases compared to the case without the ETS.
-This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level, but the buyers do not have this option under no ETS.
-That is, the global emissions could be less than the global emission caps under no ETS, while they are equal under the ETS.
 This pattern persists until 2043.
+This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level.
+In contrast, without the ETS, regions with binding regional emission caps cannot utilize surplus emission permits from other regions, resulting in global net emissions that are lower than or equal to those in the ETS scenario over time.
 In the bottom two panels of Figure 5, we compare the MAC and the SCC using the US as an illustrative example.
-The MAC of the US economy is higher without the ETS until 2057, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement.
+The MAC of the US economy without the ETS is higher until 2056, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement.
+[^27] Before 2030, the emission control rate grows fast so the MAC of the US increases rapidly along time, but after 2030, the emission control rate grows slowly such that the improvement of emission abatement technology makes the MAC decline along time until 2042.
 The SCC comparison demonstrates that the ETS has little impact on the regional SCC.
 This pattern holds for the other regions as well, as shown in Figure A.10.
-Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results show that the ETS has little impact on the marginal damages and our endogenous social discount rates.
+Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results indicate that the ETS has little impact on either the marginal climate damages or our endogenous social discount rates.
 Figure 5: Comparative analysis: effects of the ETS implementation.
@@
 To quantify these effects, we compute a compensating variation (CV) per capita associated with the ETS implementation.
-Specifically, the CV per capita for region $i$ is computed by solving the following equation:
+Specifically, the CV per capita for region $i$ is computed by numerically solving the following equation:
 $W_{i,0}(\mathbf{c}_{i}^{1}-CV)=W_{i,0}(\mathbf{c}_{i}^{0}),$ (24)
@@
 is the social welfare associated with the vector of consumption per capita $\mathbf{c}_{i}=(c_{i,0},\cdots,c_{i,t},\cdots,c_{i,T})$ with the terminal time $T=299$.
-Note that a direct comparison of the noncooperative model with the ETS to the noncooperative model without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global emissions and resulting climate damages differ across the two economies.
-That is, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the emission caps and the ETS implementation.
-Compared with the economy without the ETS, temperature under the ETS implementation is expected to be higher as shown in Figure 5.
-Consequently, some regions may experience additional climate damage from the difference of temperatures between two cases, while it is greater than the reduced abatement cost from purchasing permits or the additional profit from selling permits under the ETS.
-This leads to negative welfare effects of the ETS implementation for some regions.
-To isolate the impact of the ETS, we adjust the emission caps $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which are the optimal emission levels obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS, then we compare economies with and without the ETS under the emission caps $\overline{E}^{\prime}_{i,t}$ as these two economies have the same global emission levels and temperatures.
-Table 1 shows the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$, when the ETS implementation does not impact temperature.
-The findings indicate that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions.
-For example, in the United States, the CV per capita is \$87.22, which is a relatively small fraction of per capita consumption (0.2%).
-Among the 12 regions, Russia experiences the largest welfare improvement, with a CV per capita of \$183.53, equivalent to 2.2% of its per capita consumption.
-In contrast, Africa and non-OECD Asia have the smallest welfare gains from the global ETS.
-These results highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
-Table 1: Welfare effects of the ETS implementation.
+Table 1 shows welfare effects of the global ETS, measured by the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$.
+The results reveal significant heterogeneity in welfare effects across regions: some regions experience welfare gains, with Russia benefiting the most, showing a CV per capita of \$552.75, equivalent to 6.699% of its per capita consumption.
+In contrast, other regions, including Africa, the Middle East, and non-OECD Asia, experience negative welfare effects.
+This outcome may seem counterintuitive, as one might expect an additional market mechanism (ETS) to enhance welfare by providing greater flexibility in managing emissions.
+It is important to note that a direct comparison of the noncooperative model with and without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global net emissions and resulting climate damages differ across the two economies.
+Recall that, global average temperature is higher under ETS implementation in the baseline emission cap scenario, as shown in Figure 5.
+Therefore, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the global net emissions (and resulting climate damages) and the ETS implementation.
+As a result, some regions may experience additional climate damages due to higher temperatures, outweighing the benefits of lower abatement costs from permit purchases or additional profits from permit sales.
+This explains why some regions experience negative welfare effects from ETS implementation.
+Table 1: Welfare effects of the ETS implementation: pre-emission cap adjustment.
 US EU Japan Russia Eurasia China
-CV per capita (\$) 87.22 86.43 35.12 183.53 50.96 24.93
-CV (%) 0.201 0.344 0.124 2.225 1.017 0.250
+CV per capita (\$) 38.94 83.04 -0.03 552.75 53.18 42.90
+CV per capita (%) 0.089 0.330 0.000 6.699 1.062 0.429
 India MidEast Africa LatAm OHI OthAs
-CV per capita (\$) 37.09 54.96 13.06 56.93 69.33 8.15
-CV (%) 2.374 0.590 0.857 0.865 0.201 0.348
-7 Sensitivity Analysis
-We perform sensitivity analyses to evaluate alternative emission cap scenarios and to examine the influence of key parameters—climate damages and emission abatement—on our simulation results.
-7.1 Alternative Emission Cap Scenarios
+CV per capita (\$) 22.94 -36.03 -5.40 66.68 109.69 -18.93
+CV per capita (%) 1.469 -0.387 -0.354 1.001 0.317 -0.806
+To isolate the impact of the ETS, we adjust the emission cap $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which is the optimal level of net emissions obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS.
+We then compare economies with and without the ETS under the emission cap $\overline{E}^{\prime}_{i,t},$ ensuring two economies have the same pathways of global net emissions and temperature.
+Table 2 shows that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions.
+For example, in the United States, the CV per capita is \$116.18, which is a relatively small fraction of per capita consumption (0.267%).
+Among all regions, Russia still experiences the largest welfare improvement, with a CV per capita of \$235.17, equivalent to 2.850% of its per capita consumption, though its welfare gains are smaller than in the pre-adjustment analysis.
+This is because Russia, a country with a negative SCC, benefits from higher temperatures, meaning that its welfare gains in the pre-adjustment analysis reflect both climate-induced benefits and ETS implementation.
+Conversely, Africa and non-OECD Asia experience the smallest welfare gains from the global ETS, indicating that additional climate damages from higher temperatures contributed to the negative welfare effects observed in the pre-adjustment analysis.
+Overall, the results in the post-adjustment analysis highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
+Table 2: Welfare effects of the ETS implementation: post-emission cap adjustment.
+US EU Japan Russia Eurasia China
+CV per capita (\$) 116.18 100.26 38.60 235.17 57.49 34.62
+CV per capita (%) 0.267 0.399 0.136 2.850 1.148 0.346
+India MidEast Africa LatAm OHI OthAs
+CV per capita (\$) 33.34 42.73 16.82 75.30 100.63 12.40
+CV per capita (%) 2.135 0.459 1.103 1.130 0.291 0.529
+8 Alternative Policy Simulations
+Our model framework allows flexibility to explore alternative policy simulations.
+Recognizing that the actual implementation of a global ETS requires substantial international commitment, we consider two alternative policy scenarios to gain additional insights.
+First, we examine a partial ETS in which a major player in the world economy—the United States—does not participate, and the ETS operates among the remaining regions.
+Second, we analyze alternative net-zero pathways that vary in the stringency of global emission reduction commitments.
+8.1 Partial ETS: US Non-participation
+Recent political developments under the Trump administration have indicated that the United States is unlikely to participate in international cooperation aimed at mitigating global emissions in the near term.
+In light of this potential fragmentation of global climate cooperation, it is useful to numerically examine the economic and environmental consequences of a partial ETS.
+To this end, we simulate a scenario in which the United States refrains from any form of climate action—neither enforcing its emission cap nor participating in the ETS—while the remaining 11 regions continue to operate a cap-and-trade system.
+Figure 6 compares the key outcomes under the baseline (“full ETS”) and the partial ETS scenarios.
+As intuitively expected, US non-participation leads to a substantial increase in global emissions and a corresponding rise in global temperature, reaching 1.83 degrees Celsius—about 0.12 degrees Celsius higher than in the baseline.
+Under the partial ETS scenario, the US MAC is equal to its SCC which rises to \$336 by 2100 due to higher global average temperatures.
+Figure 6: Comparative analysis: effects of US non-participation in the global ETS
+Table 3 shows welfare effects of the US non-participation by comparing the partial ETS to the full ETS, further illustrating the distributional effects of fragmented climate cooperation.
+The United States experiences a significant welfare gain, with CV per capita of \$565—approximately 1.3% of per capita consumption—reflecting its reduced abatement burden and the low level of MACs as shown in the figure above.
+On the other hand, the other regions (except for Russia) all experience significant welfare losses, ranging from -3.8% (MidEast) to -0.01% (OHI) of per capita consumption, as they bear the costs of both higher climate damages and continued emission constraints under the ETS.
+Russia is the only exception among the participating regions, experiencing welfare gains because higher global average temperatures are expected to benefit its economy, as noted earlier.
+Overall, this exercise underscores the critical importance of sustained and collective international commitment to climate mitigation, given the strong incentives for free-riding under noncooperative settings.
+Table 3: Welfare effects of US non-participation in the global ETS
+US EU Japan Russia Eurasia China
+CV per capita (\$) 564.96 -54.00 -115.07 105.06 -129.73 -10.15
+CV per capita (%) 1.309 -0.215 -0.404 1.226 -2.561 -0.103
+India MidEast Africa LatAm OHI OthAs
+CV per capita (\$) -28.94 -327.27 -52.56 -121.99 -3.23 -79.91
+CV per capita (%) -1.867 -3.793 -3.409 -1.966 -0.009 -3.378
+8.2 Net Zero Emission Scenarios
 Along with the baseline emission cap scenario, we analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090, where all regions achieve net zero emissions by the specified year.
 The top-left and top-right panels in Figure 7 display the emission permit prices and expected temperature increases under different emission cap scenarios.
-Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,539 per ton of carbon in 2049, and the temperature rise is restricted to 1.62°C by the end of this century.
-Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$514 and \$335 per ton of carbon in 2049, respectively.
-In these scenarios, the temperature rise by the end of the century is expected to reach 1.80°C and 1.98°C above the pre-industrial level, respectively.
-This result shows that the global target of restricting the temperature rise to 1.5°C is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
+Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,621 per ton of carbon in 2049, and the temperature rise is restricted to 1.62 degrees Celsius by the end of this century.
+Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$616 and \$446 per ton of carbon in 2049, respectively.
+In these scenarios, the temperature rise by the end of the century is expected to reach 1.80 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively.
+This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
 Figure 7: Sensitivity analysis: alternative emission cap scenarios.
@@
 Our results show that stricter emission caps lead to higher MACs.
-Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,471 per ton of carbon in 2048, compared to the peak of \$1,322 in 2086 under the net zero 2090 scenario.
+Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,493 per ton of carbon in 2048, compared to the peak of \$1,322 in 2082 under the net zero 2090 scenario.
 This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC.
-We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$226 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario.
-This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC (recall that the optimal carbon tax is equal to the gap between the MAC and permit price when emissions are strictly positive).
-Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
-7.2 Alternative Climate Damages and Abatement Costs
+We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$225 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario.
+This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC.
+9 Sensitivity Analysis
 Lastly, recognizing that climate damages and emission abatement costs are key drivers of our model outcomes, including permit prices, MAC, and the SCC, we conduct sensitivity analyses on the parameters for climate damages ($\pi_{1,i}$, $\pi_{2,i}$) and emission abatement ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$).
+9.1 Climate Damage Parameters
+.
 While we incorporate climate damage projections from Burke et al., (2018) as our baseline model simulation, we additionally consider projections from Kahn et al., (2021) and Nordhaus, 2010a .
@@
 Figure 8 demonstrates the key economic and climate outcomes under different climate damage assumptions.
-The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,059 by 2069, compared to \$2,282 under the damage parameters estimated from Kahn et al., (2021) and \$2,294 under the damage parameters from Nordhaus, 2010a .
+The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,105 by 2069, compared to \$2,322 under the damage parameters estimated from Kahn et al., (2021) and \$2,334 under the damage parameters from Nordhaus, 2010a .
 With the same emission cap constraints imposed, the global temperature outcomes remain unaffected despite variations in the damage parameters.
@@
 Figure 8: Sensitivity analysis: alternative climate damage parameters.
+9.2 Abatement Cost Parameters
 For our sensitivity analysis of the emission abatement cost parameters, we consider the parameter values from Nordhaus, 2010a in addition to the baseline parameter values calibrated from Ueckerdt et al., (2019), both of which share the same functional form of abatement cost.
-As shown in Figure 8, the permit price rises to \$1,767 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 85% of the permit price projected under the baseline scenario.
-Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70°C by the end of the century, due to the emission cap constraints.
+As shown in Figure 9, the permit price rises to \$1,717 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 81% of the permit price projected under the baseline scenario.
+Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70 degrees Celsius by the end of the century, due to the emission cap constraints.
 The lower emission permit price under the abatement cost estimate of Nordhaus, 2010a reflects lower MAC, as illustrated with the US case in Figure 9, with similar patterns observed across most regions (see Appendix A.4.6).
-Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$101 by 2100—just 32% of the baseline scenario.
+Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$102 by 2100—just 33% of the baseline scenario.
 Figure 9: Sensitivity analysis: alternative abatement cost parameters
@@
 In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system.
-In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact.
-We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting.
-The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth.
-We further show that carbon taxation needs to be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal emissions.
-For strictly positive emissions, we show both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
+In our integrated assessment framework, regions participating in a global or partial ETS are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact.
+We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting among the participating regions.
+The permit prices are endogenously determined by demand and supply of emission permits in the permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth.
+For strictly positive emissions, we show both theoretically and numerically that the regional SCC is equal to the difference between the regional MAC and the market price of permits for regions participating in the ETS.
 This work has several policy implications.
-First, our results indicate that the current global target of restricting the global temperature rise to 1.5°C above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact.
+First, our results indicate that the current global target of restricting the global temperature rise to 1.5 degrees Celsius above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact.
 Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature.
@@
 This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits.
-Third, we demonstrate that the global ETS is not a pure substitute for the carbon tax, but instead that both can be implemented together to improve the efficiency of global climate policies.
+Third, we demonstrate that an ETS is not a perfect substitute for a carbon tax; rather, the two instruments are complementary and can be jointly employed to enhance policy efficiency.
+Finally, our numerical analysis highlights the strong free-rider incentives in a partial ETS, illustrating why achieving sustained global cooperation in carbon mitigation remains a major challenge.
+11 Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work, the authors used generative AI (ChatGPT) in order to improve the manuscript’s readability during the revision process.
+After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
 Appendix for Online Publication
@@
 A.4.2 Model Comparison of ETS Implementation
-Figure A.8 compares regional emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
-Figure A.8: Comparison of the ETS implementation: regional emissions.
+Figure A.8 compares regional net emissions between two cases under noncooperation with the baseline emission caps: (i) with the ETS, (ii) without the ETS.
+Figure A.8: Comparison of regional net emissions with and without the ETS.
 Figure A.9 compares the regional MAC under noncooperation with the baseline emission caps. We compare two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.9: Comparison of the ETS implementation: regional MAC.
+Figure A.9: Comparison of regional MAC with and without the ETS.
 Figure A.10 compares the regional SCC under noncooperation with the baseline emission caps, comparing two cases: (i) with the ETS, (ii) without the ETS.
-Figure A.10: Comparison of the ETS implementation: regional SCC.
-A.4.3 Sensitivity Analysis over Emission Caps
+Figure A.10: Comparison of regional SCC with and without ETS.
+A.4.3 Alternative Policy Simulations: Partial ETS
+Figure A.11 compares the regional MAC under noncooperation, comparing two cases: (i) with the full ETS, (ii) with a partial ETS.
+Figure A.11: Comparison of regional MAC under full and partial ETS scenarios.
+Figure A.12 compares the regional SCC under noncooperation, comparing two cases: (i) with the full ETS, (ii) with the partial ETS without participation of the United States.
+Figure A.12: Comparison of regional SCC under the full and the partial ETS scenarios.
+A.4.4 Alternative Policy Simulations: Net Zero Scenarios
 In Figure A.13, we compare the MAC for the noncooperative model with the ETS across alternative emission cap scenarios.
@@
 Similarly, Figure A.14 displays the SCC of the noncooperative model with the ETS across different emission cap scenarios.
-Figure A.12: Comparison of regional SCC under different.
-A.4.4 Sensitivity Analysis over Climate Damage Parameters
-In Figure A.13, we compare the MAC for the noncooperative model with the ETS under alternative values of the climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on climate damage projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline climate damage projection from Burke et al., (2018).
+Figure A.14: Comparison of regional SCC under different emission caps.
+A.4.5 Sensitivity Analysis over Climate Damage Parameters
+In Figure A.15, we compare the MAC for the noncooperative model with the ETS under alternative climate damage parameters ($\pi_{1,i}$, $\pi_{2,i}$), based on projections from Kahn et al., (2021) and Nordhaus, 2010a alongside the baseline projection from Burke et al., (2018).
 Figure A.15: Comparison of regional MAC under different estimates of climate damages.
@@
 Figure A.16: Comparison of regional SCC under different estimates of climate damages.
-A.4.5 Sensitivity Analysis over Abatement Cost Parameters
+A.4.6 Sensitivity Analysis over Abatement Cost Parameters
 In Figure A.17, we compare the MAC for the noncooperative model with the ETS under alternative estimates of the emissions abatement parameters ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$), calibrated from Ueckerdt et al., (2019) and Nordhaus, 2010a .
@@
 Figure A.18: Comparison of of regional SCC under different estimates of emissions abatement cost.
-A.4.6 Sensitivity over TFP Growth Rates
+A.4.7 Sensitivity over TFP Growth Rates
 Figure A.19 compares key model outcomes under different TFP growth rates: the baseline TFP growth rates derived from Burke et al., (2018) and the alternative rates based on Nordhaus, 2010a .